### v1

# Computing Wasserstein Barycenters through Gradient Flows

Abstract
Wasserstein barycenters provide a powerful tool for aggregating probability measures, while leveraging the geometry of their ambient space. Existing discrete methods suffer from poor scalability, as they require access to the complete set of samples from input measures. We address this issue by recasting the original barycenter problem as a gradient flow in the Wasserstein space. Our approach offers two advantages. First, we achieve scalability by sampling mini-batches from the input measures. Second, we incorporate functionals over probability measures, which regularize the barycenter problem through internal, potential, and interaction energies. We present two algorithms for empirical and Gaussian mixture measures, providing convergence guarantees under the Polyak-Łojasiewicz inequality. Experimental validation on toy datasets and domain adaptation benchmarks show that our methods outperform previous discrete and neural net-based methods for computing Wasserstein barycenters.
1 Introduction
(a) Static barycenter.
(b) Barycenter flow.
Figure 1: In (a), we show the usual static notion of the Wasserstein barycenter (in red), which minimizes the sum of distances to the input measures (in blue). In (b), we show our notion of barycenter as a gradient flow, flowing an initial measure $P_{0}$ (purple) to the barycenter $P^{\star}$ (yellow) of the input measures.
Wasserstein barycenters [1] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [2]. This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures. Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [3]. As such, they contributed to several areas of machine learning, including model averaging [4], ensembling [5], data augmentation [6], distillation [7, 8], dictionary learning [9], domain adaptation [10, 11, 12], and Bayesian learning [13].
Existing methods for computing Wasserstein barycenters can be divided into three categories. The first category relies on discretization of the input measures and the barycentric measure through empirical measures. These algorithms may be subdivided into fixed-support [14, Algorithm 1][15, 16], and free-support [14, Algorithm 2][11].
The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models. In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [17, 18]. For Gaussian mixtures, [19] shows connections with multi-marginal OT [20] between the components of the input mixtures, while [12] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [14, Algorithm 2].
The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples. In this sense, [21] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [22]. Furthermore, [23] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [24]. More recently, [25] proposes computing barycenters with normalizing flows.
Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$. In this sense, neural networks have a clear advantage over discrete methods, since they can operate at the level of mini-batches. There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$. So far, only [25] has considered experiments in this context. In this regard, methods such as [21, 26] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [27, 28], which is an appealing framework for optimization in the space of probability measures. This new understanding of Wasserstein barycenters has several advantages. First, we can scale the discrete barycenter problem, by drawing mini-batches from the input measures. Second, we can scale with respect to the number of input measures, as the representation of the barycentric measure is independent of $K$. Third, based on the established literature on gradient flows, we can define regularized Wasserstein barycenters with suitable functionals on the space of probability measures, which were not possible with previous methods. Lastly, we can modify the underlying metric to integrate label information between samples, leading to barycenters that respect class structure and provide more accurate and semantically meaningful estimates. We show an overview of our approach in Figure 1. Our contributions are:
  • We introduce a unified framework for computing Wasserstein barycenters based on gradient flows, which provides a principled and flexible approach to optimizing in the space of probability measures.
  • We generalize algorithms from the state-of-the-art [14, 11, 12] under a common framework.
  • We offer theoretical guarantees for the convergence of our algorithms, which also apply to previous works.
  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [19, 12].
  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [21, 23, 24, 25], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling. We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
Paper organization. Section 2 presents the notation and preliminaries of our paper. Section 3 covers our proposed algorithms, including the theoretical results. Section 4 describes our experiments. Finally, section 5 concludes this paper.
2 Background
Throughout this paper, $(\Omega,d)$ is a metric space. We denote by $P$ and $Q$, probability measures on $\mathcal{P}(\Omega)$, which is the space of measures over $\Omega$. For $p\in[1,+\infty)$ we further denote by $\mathcal{P}_{p,ac}(\Omega)$, the sub-set of absolutely continuous measures w.r.t. the Lebesgue measure such that $\int_{\Omega}d(z,z_{0})^{p}dP(z)<+\infty$. In our case, we have access to these measures through their samples, denoted $z^{(P)}\sim P$. The set $\Delta_{K}$ denotes the $K$-simplex, i.e., $y\in\Delta_{n}=\{y\in\mathbb{R}^{n}:\sum_{i=1}^{n}y_{i}=1\text{ and }y_{i}\geq 0\text{ }\forall i\}$. We denote functionals over $\mathcal{P}_{2,\text{ac}}$ with blackboard bold letters (e.g., $\mathbb{F}$), and their (Wasserstein) gradients by $\nabla\mkern-10.0mu\nabla$. We offer a short introduction in Section 3 in the Appendix.
2.1 Empirical Optimal Transport
In this section, given $n$ i.i.d. samples from $P$, we can approximate a probability $P$ empirically through,
$\hat{P}(z)=\dfrac{1}{n}\sum_{i=1}^{n}\delta(z-z_{i}^{(P)}),$ (1)
which gives a finite parametrization for $P$. Henceforth we use a hat to indicate empirical measures (e.g., $\hat{P}$).
OT [2] is a field of mathematics concerned with the transportation of mass at least effort. We refer readers to [29] for a computational treatment of the subject, and [30] for applications in machine learning. Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost. OT was originally founded by Monge [31]. In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\int_{\Omega}c(z,T(z))dP(z).$ (2)
Alternatively, Kantorovich [32] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}\int_{\Omega}c(z,z^{\prime})\gamma(z,z)$ (3)
Given $p\in[1,+\infty)$, when $c(z,z^{\prime})=d(z,z^{\prime})^{p}$, the previous problems define the $p$-Wasserstein distance,
$\mathbb{W}_{p}(P,Q)^{p}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\int_{\Omega}\int_{\Omega}d(z,z^{\prime})^{p}d\gamma(z,z^{\prime}).$ (4)
Equation 4 defines a metric in $\mathcal{P}_{p,ac}(\Omega)$. Conceptually, the Wasserstein distance lifts the metric $d$ on $\Omega$ to $\mathbb{W}_{p}$ on $\mathcal{P}_{p,ac}(\Omega)$. Based on this idea, we can define barycenters and gradient flows in the space of probability measures.
2.2 Gaussian Mixture Optimal Transport
In comparison with equation 1, we can approximate $P$ through a Gaussian Mixture Model (GMM),
$P(z)=\sum_{i=1}^{n}\pi_{i}^{(P)}\mathcal{N}(z|\mu_{i}^{(P)},\Sigma_{i}^{(P)}),$ (5)
where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM. The general theory of OT between GMMs was first presented in [19]. The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
$\omega^{\star}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$
$\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan. The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [33] in terms of their parameters,
$\mathbb{W}_{2}(P,P^{\prime})^{2}=\lVert\mu-\mu^{\prime}\rVert_{2}^{2}+\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{\frac{1}{2}}\Sigma^{\prime}\Sigma^{\frac{1}{2}})^{\frac{1}{2}}).$
Based on these ideas, one may define a Wasserstein-type distance between GMMs,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{inf}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$ (6)
2.3 Wasserstein Barycenter and its Variants
In the metric setting, the barycenter problem is known as Fréchet [34] or Karcher [35] means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
$P^{\star}=\underset{P\in\mathcal{P}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{Q}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}\biggr\}.$ (7)
In general, equation 7 does not have a closed-form solution. However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [14], Gaussian measures [1, 36] or Gaussian mixtures [19, 12], and a further link can be made to multi-marginal OT [20].
The work of [14] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures. The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
$z_{\tau+1,i}^{(P)}=(1-\alpha)z_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(z_{\tau,i}^{(P)}),$ (8)
where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$. This strategy is reminiscent of the fixed-point approach of [36]. Furthermore, [11] shows that, for feature-label joint measures, the labels can be propagated through [37],
$y_{\tau+1,i}^{(P)}=(1-\alpha)y_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(y_{\tau,i}^{(P)}),$ (9)
where $y_{\tau,i}^{(P)}$ are one-hot encoded vectors.
Previous work has extended equation 8 to the GMM setting, when the components are axis-aligned Gaussians, i.e., $\Sigma_{i}^{(P)}=\text{diag}(\sigma_{i}^{(P)})$. In this case,
$\mu_{i,\tau+1}^{(P)}=(1-\alpha)\mu_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\mu_{i,\tau}^{(P)}),$ (10)
$\sigma_{i,\tau+1}^{(P)}=(1-\alpha)\sigma_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\sigma_{i,\tau}^{(P)}),$
There are a few limitations with equations 8 and 10. First, the empirical iterations do not scale well to large datasets. Indeed, they assume the availability of all samples of $Q_{k}$ per iteration. Second, the GMM iterations in equation 10 are restricted to axis-aligned GMM s. This paper addresses these gaps using gradient flows.
2.4 Gradient Flows
In the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional. A gradient flow is the solution to the differential equation,
$\begin{cases}\dot{x}(t)=-\nabla F(x(t))&\text{for }t>0,\\ x(0)=x_{0},\end{cases}$ (11)
where $x_{0}$ is the initial condition. In $\mathcal{P}_{p,\text{ac}}(\Omega)$, gradient flows correspond to the continuity equation,
$\partial_{t}P_{t}=-\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})),$ (12)
where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures. Usually, these functionals take the following form [38],
$\mathbb{F}(P)=$ $\underbrace{\int G(P(z))dz}_{\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\mathbb{V}(P)}+$ (13)
$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\mathbb{U}(P)},$
where $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ are the internal, potential, and interaction energies, respectively. Here, $G:\mathbb{R}\rightarrow\mathbb{R}$ is convex and superlinear, and $V:\Omega\rightarrow\mathbb{R}$ and $U:\Omega^{2}\rightarrow\mathbb{R}$ are convex and sufficiently smooth. Henceforth, we denote $\mathbb{F}^{\star}=\text{inf}_{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}\mathbb{F}(P)$.
3 Wasserstein Barycenters as Gradient Flows
In this section, we describe a new method for computing empirical and Gaussian mixture Wasserstein barycenters. Our main idea is using the functional:
$\mathbb{F}(P)=\mathbb{B}_{\mathcal{Q}}(P)+\mathbb{G}(P)+\mathbb{V}(P)+\mathbb{U}(P),$ (14)
where $\mathbb{B}_{\mathcal{Q}}(P)$ is the barycenter objective defined in equation 7, and $\mathbb{G},\mathbb{V}$ and $\mathbb{U}$ are the energies defined in equation 13.
3.1 Empirical Flow
In the empirical case, the barycenter is approximated through a finite set of particles (see equation 1) $\{z_{i}^{(P)}\}_{i=1}^{n}$. Hence, we may rewrite equation 14 as,
$\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{min}}\,\underbrace{\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}d(z_{i}^{(P)},z_{j}^{(Q_{k})})^{p}}_{\mathbb{B}_{\mathcal{Q}}(\hat{P})}$+
$\underbrace{\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)})}_{\mathbb{V}(\hat{P})}+\underbrace{\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)})}_{\mathbb{U}(\hat{P})}$ (15)
Here, we omit the internal energy term, $\mathbb{G}(P)$, as it is not defined for empirical measures. We solve the previous optimization problem through a block-coordinate descent strategy, that is,
$\gamma_{k}$ $=\underset{\gamma\in\Gamma(\hat{P}_{\tau},\hat{Q}_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{i,j}d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{p},$
$z_{\tau+1,i}^{(P)}$ $=z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}).$ (16)
In equation 16, $\gamma_{k}$ is used to compute $\mathbb{B}_{\mathcal{Q}}$ and is updated for each $\tau$. We provide in Algorithm 1 the pseudo-code for this strategy. Our algorithm assumes access to $Q_{k}$ through sampling, that is, we do not necessarily have access to all samples of $Q_{k}$ at once.
Algorithm 1 Empirical barycenter Wasserstein flow using Gradient Descent.
[…]
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
5:  for $k\leftarrow 1$ to $K$ do
6:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i.i.d. from $Q_{k}$.
7:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{p}(\hat{P}_{\tau},\hat{Q}_{k})^{p}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\tau})$
10:end for
Remark 3.1.
In line 5 of Algorithm 1, $\hat{Q}_{k}$ denotes the empirical measure (c.f., equation 1) supported on the mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ sampled from $Q_{k}$.
3.2 Gaussian Mixture Flow
In the case of GMMs, we flow the parameters $\theta=\{(\pi_{k},\mu_{k},\Sigma_{k})\}_{k=1}^{K}$ of $P_{\theta}=\sum\pi_{i}\mathcal{N}(\cdot|\mu_{i},\Sigma_{i})$. Similarly to equation 16, we have:
$\omega_{k}$ $=\underset{\omega\in\Gamma(\pi,\pi_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\omega_{i,j}\mathbb{W}_{2}(P_{\theta_{\tau},i},Q_{k,j})^{2},$
$\theta_{\tau+1,i}$ $=\theta_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}}).$ (17)
Here, $\omega_{k}$ is a transport plan between GMM components. In general, the energy functionals do not have closed form with respect to the GMM parameters. However, we can estimate them using Monte-Carlo and the reparametrization trick, for instance,
$z_{i}$ $=L_{i}\varepsilon+\mu_{i},\quad i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id}),$ (18)
This strategy allows us to propagate the gradients to the parameters of $P_{\tau}$. We show the overall approach in Algorithm 2.
Algorithm 2 GMM barycenter flow using Gradient Descent.
[…]
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  Sample $\{z_{i}=L_{i}\varepsilon+\mu_{i}:i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id})\}$,
5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{G}(\hat{P}_{\tau})+\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
6:  for $k\leftarrow 1$ to $K$ do
7:   $\mathbb{F}(P_{\theta_{\tau}})\leftarrow\mathbb{F}(P_{\theta_{\tau}})+\lambda_{k}\mathbb{MW}_{2}(P_{\theta_{\tau}},Q_{k})^{2}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}})$
10:end for
Covariance matrix parametrization. Optimizing $\mathbb{F}$ with respect to $\Sigma_{k}$ is challenging, due to the symmetric positive definite constraint. We enforce this constraint by writing $\Sigma_{k}=L_{k}L_{k}^{T}$, where $L_{k}$ is the lower-triangular Cholesky factor. There are two advantages to this strategy. First, optimizing with respect to $L_{k}$ is empirically more stable than $\Sigma_{k}$. Second, we can use $L_{k}$ directly in the reparametrization trick in equation 18.
3.3 Flows over Joint Measures
In [14], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work. As we show in our experiments, in machine learning applications one encounters $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Whenever we deal with labeled measures (i.e., $z=(x,y)$), we adopt,
$d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (19)
where $\beta\geq 0$ is a parameter that balances the feature distance terms and the label distance terms.
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in [10] and [39], is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
$y_{i,c}^{(P)}=\text{softmax}(\ell_{i,1}^{(P)},\cdots,\ell_{i,n_{c}}^{(P)})=\dfrac{\exp(\ell_{i,c}^{(P)})}{\sum_{c=1}^{n_{c}}\exp(\ell_{i,c}^{(P)})},$
thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$. From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
For GMMs, we equip their components with labels $\nu_{k}\in\Delta_{n_{\text{classes}}}$. Using the ground-cost in equation 19, we are able to write the $\mathbb{MW}_{2}(P,Q)^{2}$ as,
$\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}^{\star}(\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\beta\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}^{2}),$
where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal. This approach was first proposed in [12], but the authors relied on heuristic arguments to justify this modeling choice. We present the following proposition justifying the label term,
Proposition 3.1.
Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be two GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$. Let the ground cost $c$ be,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
Functionals for joint measures. One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by [39] for transfer learning problems. Here we propose the following functionals,
$V_{E}(z)$ $=-\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (20)
$U_{R}(z,z^{\prime})$ $=\begin{cases}h(d(x,x^{\prime}))&\text{if }y\neq y^{\prime},\\ 0&\text{otherwise}\end{cases},$ (21)
where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below. For instance, in our experiments we use the hinge loss, $h(u)=\text{max}(0,\text{margin}-d)$, where $\text{margin}\geq 0$ is a margin parameter. Equations 20 and 21 correspond to entropy and repulsion terms. The first functional penalizes barycenters that have fuzzy labels. The second functional encourages classes to be well separated.
3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [28, Section 4.4]. As a result, we are minimizing a non-convex functional. Henceforth, we focus on $p=2$, as this is the setting where most results are available. Explicit constants and proofs are available in Section 4 in the Appendix.
Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [40, 41],
$\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (22)
which is covered in [42]. Under this assumption, we have the following convergence result,
Theorem 3.1.
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball on $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Assume the inequality 22. Then, the following holds,
$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,m}}{n}}\end{aligned},$ (23)
where the expectation on the l.h.s. is taken with respect to samples from $Q_{k}$. The constants $C_{R}$ and $C_{d,m}$ depend on the radius $R$, number of barycenter samples $m$, and dimensions $d$.
The right hand side of inequality 23 is composed of two terms. The first term comes from the PL inequality, and covers the convergence towards the minimizer of $\mathbb{B}$. The second term covers the error of empirical approximations. In this sense, letting $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximation.
Next, we derive a new result for the convergence of the GMM gradient flow (cf. Algorithm 2). Here, we assume $P$ and $Q_{k}\in\mathcal{Q}$ are GMMs. Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm. Our result relies on [43, Theorem 3.1]. Therefore, we assume,
  1. 1.
    bounded GMM parameters, i.e., $\lVert\mu_{i}^{(Q_{k})}\lVert_{2}\leq R_{\mu}$, $\sqrt{\text{Tr}(\Sigma_{i}^{(Q_{k})})}\leq R_{\Sigma}$ $\forall i,\forall k$ (resp. $\hat{\mu}_{i}^{(Q_{k})}$ and $\hat{\Sigma}_{i}^{(Q_{k})}$, the estimated parameters); furthermore,
  2. 2.
    Empirical convergence rates, i.e., $\mathbb{E}[\lVert\pi^{(Q_{k})}-\hat{\pi}^{(Q_{k})}\rVert_{1}]\leq\rho_{\pi}$, $\mathbb{E}[\lVert\mu_{i}^{(Q_{k})}-\hat{\mu}_{i}^{(Q_{k})}\rVert_{2}]\leq\rho_{\mu}$, and $\mathbb{E}[d_{\text{Bures}}(\Sigma_{i}^{(Q_{k})},\hat{\Sigma}_{i}^{(Q_{k})})]\leq\rho_{\Sigma}$, $\forall i$ and $\forall k$.
Here the expectations are taken with respect to samples drawn from $Q_{k}$. Under these conditions we have,
Theorem 3.2.
Let $P$ and $Q$ be two labeled GMMs with bounded parameters and satisfying the empirical convergence rates. Then, the following holds,
$\mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}\tau}}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{\lambda,\mathcal{Q}},$
where $C_{\lambda,\mathcal{Q}}$ is a constant that depends on the coordinates $\lambda$ and GMMs in $\mathcal{Q}$.
Here $C_{\lambda,\mathcal{Q}}$ depends on the $(\pi^{(Q_{k})},\mu^{(Q_{k})},\Sigma^{(Q_{k})})$ of each GMM, but not on $\nu^{(Q_{k})}$. Indeed, since we fit a GMM per class, the labels are estimated exactly (i.e., $P_{i,x}$ and $\hat{P}_{i,x}$ belong to the same class).
4 Experiments
We divide our experiments into a toy example (Section 4.1) to illustrate our methods, and multi-source domain adaptation experiments (Section 4.2). Due to length constraints, we include additional experiments in Sections 5 and 6 in the Appendix. In particular, in Table 1 in the Appendix, we show a running-time comparison demonstrating the scalability of our method.
4.1 Toy Example
In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$. This example has been used in several state-of-the-art Wasserstein barycenter works [21, 25]. Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold. We show a summary of our results in Figure 3.
Figure 2: Location-scatter family generated by a Swiss-roll measure, $Q_{0}$. Each measure $Q_{k}=T_{k,\sharp}Q_{0}$, for $T_{k}(x)=A_{k}x+b_{k}$.
First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h). Quantitatively, empirical methods, notably [14, Algorithm 2] and our WGF algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
(a) CW2B [21].
(b) NOT [23].
(c) Discrete [14].
(d) WGF.
[…]
(g) U-NOT [24].
(h) NormFlow [25].
(i) Discrete [11].
(j) WGF.
[…]
Figure 3: Comparison between Wasserstein barycenter solvers. Colored scatter plots indicate labeled barycenters. For each solver, we compute the Wasserstein distance between its solution $\hat{P}$ and the ground-truth $P^{\star}$, shown in the title of each sub figure (best seen on screen). Overall, using label information leads to barycenters that better approximate the ground-truth barycenter.
Second, we experiment with integrating labels in the ground cost as described in Section 3.3. These methods are shown in Figures 3 (i) through (l). In all cases, integrating labels produces barycenters that are closer to the ground-truth. We conclude that using the labels gives a strong inductive bias in the barycenter computation, which helps explain the gain in performance of labeled barycenters in the next section.
4.2 Multi-Source Domain Adaptation
Problem Formulation. One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [44, 45, 46]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk in the target domain measure,
$\mathcal{R}_{Q_{T}}(h)=\mathbb{E}_{(x,y)\sim Q_{T}}[\mathcal{L}(y,h(x))],$
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data. In addition, similar to [10, 12], we align the barycenter with the target through OT [47]. More details are available in Section 6 in the Appendix.
Benchmark Backbone # Samples # Domains # Dim. # Classes
Office31 ResNet50 [48] 3287 3 2048 31
BCI-CIV-2a CBraMod [49] 5184 10 200 4
TEP CNN [50] 17289 6 128 29
Office Home ResNet101 [48] 15500 4 2048 65
ISRUC CBraMod [49] 89240 100 512 5
Table 1: Overview of benchmarks used for domain adaptation, sorted by number of samples.
Experimental Setting. We run our experiments on five benchmarks: Office 31 [51], Office Home [52], BCI-CIV-2a [53], ISRUC [54], and TEP [50]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. These methods are: discrete barycenters [14, Algorithm 2], normalizing flows [25], continuous 2-Wasserstein barycenters [21], neural OT barycenters [23], and unbalanced neural OT barycenters [24]. We also include labeled barycenter strategies [11, 55]. A brief review of these methods is available in Section 2 in the Appendix. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT [56], DaDiL-R and E [11], and GMM-DaDiL [12]. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchamrk), and we measure the classification accuracy, i.e., the percentage of correct predictions.
Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP
Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN
Source-Only - 86.40 75.95 50.30 76.63 78.48
WJDOT [56] - 86.80 76.59 N/A 76.95 86.13
DaDiL-R [11] - 89.91 77.86 53.41 74.68 86.14
DaDiL-E [11] - 89.79 78.14 N/A 75.89 85.87
GMM-DaDiL [12] - 90.63 78.81 57.10 75.47 86.85
Discrete [14] 81.94 70.42 57.38 70.67 83.81
NormFlow [25] 85.91 76.73 55.45 78.04 82.89
CW2B [21] 86.37 76.44 57.25 75.84 85.83
NOT [23] 86.22 75.36 57.29 76.39 84.93
U-NOT [24] 86.97 76.85 57.17 77.06 85.43
Discrete [11] 87.93 77.09 57.43 78.20 86.09
GMM [12] 88.54 77.87 57.04 74.58 84.67
WGF (ours) 88.14 77.83 57.50 79.78 86.21
WGF-GMM (ours) 89.31 78.71 57.35 78.72 86.77
Table 2: Average classification accuracy on target domains. In total, we compare eight methods against our WGF strategy. For barycenter solvers, we indicate by ✓methods that compute barycenters on $\Omega=\mathcal{X}\times\mathcal{Y}$. Bold numbers represent the best barycenter method, and underlined numbers represent second best.
Main results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Section 6 in the Appendix. Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods. We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [47, 10, 11, 12]. Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
Figure 4: t-SNE [57] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark. Colors represent different classes from 11 to $65$. Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
Visualization. We now visualize the effect of adding the repulsion interaction energy functional on the Office-Home benchmark. Adding $\mathbb{U}$ leads to classes that are well separated. For computer vision experiments we use ResNets [48] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$. This choice encourages class clusters to live in orthogonal sub-spaces of the latent space. We show a comparison of the obtained barycenters in Figure 4. Overall, the classes tend to have less overlap when the interaction energy is added to the barycenter functional.
Ablation. We ablate the effect of $\mathbb{V}$ and $\mathbb{U}$ in the Office Home benchmark. For this benchmark, we use $\mathbb{V}(P)=\mathbb{W}_{2}(\hat{P},\hat{Q}_{T})^{2}$. The intent here is to isolate the contribution of each component to the state-of-the-art results in Table 2. Our results are summarized in Table 3, where we further compare our results with other related barycenter methods. In the empirical setting, the use of Algorithm 1 alone already presents a gain in performance compared to its discrete counterpart. Overall, the combination of $\mathbb{B}+\mathbb{V}+\mathbb{U}$ consistently gives the best performance, showing the importance of adding underlying structure to the Wasserstein barycenter in domain adaptation.
Method $\mathcal{X}\times\mathcal{Y}$ $\mathbb{B}$ $\mathbb{B}+\mathbb{V}$ $\mathbb{B}+\mathbb{U}$ $\mathbb{B}+\mathbb{V}+\mathbb{U}$
Discrete [14] 70.42 N/A N/A N/A
Discrete [11] 77.09 N/A N/A N/A
WGF (ours) 77.48 77.07 77.08 77.83
GMM [12] 77.87 N/A N/A N/A
WGF-GMM (ours) 77.28 78.28 78.12 78.71
Table 3: Ablation of average target domain classification accuracy on the Office Home benchmark.
5 Conclusion
In this paper we have introduced a new framework for computing Wasserstein barycenters, based on Wasserstein gradient flows [27, 28]. Our approach addresses scalability issues in discrete barycenter solvers [14], and is capable of regularizing the underlying barycenter through internal, potential, and interaction functionals. We presented gradient flow Algorithms 1 and 2, for empirical and Gaussian mixture measures, respectively. We further proved convergence guarantees based on the PL inequality.
We empirically tested our method against existing discrete [14, 11] and neural-network [21, 23, 24, 25]-based barycenter solvers on computer vision, neuroscience, and chemical engineering benchmarks, demonstrating that our proposed methods consistently achieve state-of-the-art performance. Our empirical findings (Table 2) further demonstrate that incorporating label information in the OT objective is key for domain adaptation. This finding highlights an interesting gap in the literature, where neural net-based solvers are not capable of exploiting this information, and thus have sub-optimal performance in domain adaptation. We leave an investigation of this issue for future work.