### v2

# Computing Wasserstein Barycenters through Gradient Flows

Abstract
Wasserstein barycenters provide a powerful tool for aggregating probability measures, while leveraging the geometry of their ambient space. Existing discrete methods suffer from poor scalability, as they require access to the complete set of samples from input measures. We address this issue by recasting the original barycenter problem as a gradient flow in the Wasserstein space. Our approach offers two advantages. First, we achieve scalability by sampling mini-batches from the input measures. Second, we incorporate functionals over probability measures, which regularize the barycenter problem through internal, potential, and interaction energies. We present two algorithms for empirical and Gaussian mixture measures, providing convergence guarantees under the Polyak-Łojasiewicz inequality. Experimental validation on toy datasets and domain adaptation benchmarks show that our methods outperform previous discrete and neural net-based methods for computing Wasserstein barycenters.
1 Introduction
(a) Static barycenter.
(b) Barycenter flow.
Figure 1: In (a), we show the usual static notion of the Wasserstein barycenter (in red), which minimizes the sum of distances to the input measures (in blue). In (b), we show our notion of barycenter as a gradient flow, flowing an initial measure $P_{0}$ (purple) to the barycenter $P^{\star}$ (yellow) of the input measures.
Wasserstein barycenters [agueh2011barycenters] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [villani2008optimal]. This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures. Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [kloeckner2010geometric]. As such, they contributed to several areas of machine learning, including model averaging [singh2020model], ensembling [coz2023barycenter], data augmentation [zhu2023interpolation], distillation [liu2023dataset, montesuma2024multisource], dictionary learning [schmitz2018wasserstein], domain adaptation [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter], and Bayesian learning [srivastava2018scalable].
Existing methods for computing Wasserstein barycenters can be divided into three categories. The first category relies on discretization of the input measures and the barycentric measure through empirical measures. These algorithms may be subdivided into fixed-support [cuturi2014fast, Algorithm 1][benamou2015iterative, janati2020debiased], and free-support [cuturi2014fast, Algorithm 2][montesuma2023multi].
The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models. In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [chewi2020gradient, kroshnin2021statistical]. For Gaussian mixtures, [delon2020wasserstein] shows connections with multi-marginal OT [pass2015multi] between the components of the input mixtures, while [montesuma2024lighter] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [cuturi2014fast, Algorithm 2].
The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples. In this sense, [korotin2021continuous] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [amos2017input]. Furthermore, [kolesov2024estimating] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [gazdieva2024robust]. More recently, [visentin2025computing] proposes computing barycenters with normalizing flows.
Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$. In this sense, neural networks have a clear advantage over discrete methods, since they can operate at the level of mini-batches. There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$. So far, only [visentin2025computing] has considered experiments in this context. In this regard, methods such as [korotin2021continuous, korotin2022wasserstein] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [ambrosio2008gradient, santambrogio2017euclidean], which is an appealing framework for optimization in the space of probability measures. This new understanding of Wasserstein barycenters has several advantages. First, we can scale the discrete barycenter problem, by drawing mini-batches from the input measures. Second, we can scale with respect to the number of input measures, as the representation of the barycentric measure is independent of $K$. Third, based on the established literature on gradient flows, we can define regularized Wasserstein barycenters with suitable functionals on the space of probability measures, which were not possible with previous methods. Lastly, we can modify the underlying metric to integrate label information between samples, leading to barycenters that respect class structure and provide more accurate and semantically meaningful estimates. We show an overview of our approach in Figure 1. Our contributions are:
  • We introduce a unified framework for computing Wasserstein barycenters based on gradient flows, which provides a principled and flexible approach to optimizing in the space of probability measures.
  • We generalize algorithms from the state-of-the-art [cuturi2014fast, montesuma2023multi, montesuma2024lighter] under a common framework.
  • We offer theoretical guarantees for the convergence of our algorithms, which also apply to previous works.
  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [delon2020wasserstein, montesuma2024lighter].
  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling. We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
Paper organization. Section 2 presents the notation and preliminaries of our paper. Section 3 covers our proposed algorithms, including the theoretical results. Section 4 describes our experiments. Finally, section 5 concludes this paper.
2 Background
Throughout this paper, $(\Omega,d)$ is a metric space. We denote by $P$ and $Q$, probability measures on $\mathcal{P}(\Omega)$, which is the space of measures over $\Omega$. For $p\in[1,+\infty)$ we further denote by $\mathcal{P}_{p,ac}(\Omega)$, the sub-set of absolutely continuous measures w.r.t. the Lebesgue measure such that $\int_{\Omega}d(z,z_{0})^{p}dP(z)<+\infty$. In our case, we have access to these measures through their samples, denoted $z^{(P)}\sim P$. The set $\Delta_{K}$ denotes the $K$-simplex, i.e., $y\in\Delta_{n}=\{y\in\mathbb{R}^{n}:\sum_{i=1}^{n}y_{i}=1\text{ and }y_{i}\geq 0\text{ }\forall i\}$. We denote functionals over $\mathcal{P}_{2,\text{ac}}$ with blackboard bold letters (e.g., $\mathbb{F}$), and their (Wasserstein) gradients by $\nabla\mkern-10.0mu\nabla$. We offer a short introduction in Section 3 in the Appendix.
2.1 Empirical Optimal Transport
In this section, given $n$ i.i.d. samples from $P$, we can approximate a probability $P$ empirically through,
$\hat{P}(z)=\dfrac{1}{n}\sum_{i=1}^{n}\delta(z-z_{i}^{(P)}),$ (1)
which gives a finite parametrization for $P$. Henceforth we use a hat to indicate empirical measures (e.g., $\hat{P}$).
OT [villani2008optimal] is a field of mathematics concerned with the transportation of mass at least effort. We refer readers to [peyre2019computational] for a computational treatment of the subject, and [montesuma2024recent] for applications in machine learning. Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost. OT was originally founded by Monge [monge1781memoire]. In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\int_{\Omega}c(z,T(z))dP(z).$ (2)
Alternatively, Kantorovich [kantorovich1942transfer] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}\int_{\Omega}c(z,z^{\prime})\gamma(z,z)$ (3)
Given $p\in[1,+\infty)$, when $c(z,z^{\prime})=d(z,z^{\prime})^{p}$, the previous problems define the $p$-Wasserstein distance,
$\mathbb{W}_{p}(P,Q)^{p}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\int_{\Omega}\int_{\Omega}d(z,z^{\prime})^{p}d\gamma(z,z^{\prime}).$ (4)
Equation 4 defines a metric in $\mathcal{P}_{p,ac}(\Omega)$. Conceptually, the Wasserstein distance lifts the metric $d$ on $\Omega$ to $\mathbb{W}_{p}$ on $\mathcal{P}_{p,ac}(\Omega)$. Based on this idea, we can define barycenters and gradient flows in the space of probability measures.
2.2 Gaussian Mixture Optimal Transport
In comparison with equation 1, we can approximate $P$ through a Gaussian Mixture Model (GMM),
$P(z)=\sum_{i=1}^{n}\pi_{i}^{(P)}\mathcal{N}(z|\mu_{i}^{(P)},\Sigma_{i}^{(P)}),$ (5)
where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM. The general theory of OT between GMMs was first presented in [delon2020wasserstein]. The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
$\omega^{\star}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$
$\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan. The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [takatsu2011wasserstein] in terms of their parameters,
$\mathbb{W}_{2}(P,P^{\prime})^{2}=\lVert\mu-\mu^{\prime}\rVert_{2}^{2}+\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{\frac{1}{2}}\Sigma^{\prime}\Sigma^{\frac{1}{2}})^{\frac{1}{2}}).$
Based on these ideas, one may define a Wasserstein-type distance between GMMs,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{inf}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$ (6)
2.3 Wasserstein Barycenter and its Variants
In the metric setting, the barycenter problem is known as Fréchet [frechet1948elements] or Karcher [karcher1977riemannian] means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
$P^{\star}=\underset{P\in\mathcal{P}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{Q}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}\biggr\}.$ (7)
In general, equation 7 does not have a closed-form solution. However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [cuturi2014fast], Gaussian measures [agueh2011barycenters, alvarez2016fixed] or Gaussian mixtures [delon2020wasserstein, montesuma2024lighter], and a further link can be made to multi-marginal OT [pass2015multi].
The work of [cuturi2014fast] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures. The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
$z_{\tau+1,i}^{(P)}=(1-\alpha)z_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(z_{\tau,i}^{(P)}),$ (8)
where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$. This strategy is reminiscent of the fixed-point approach of [alvarez2016fixed]. Furthermore, [montesuma2023multi] shows that, for feature-label joint measures, the labels can be propagated through [redko2019optimal],
$y_{\tau+1,i}^{(P)}=(1-\alpha)y_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(y_{\tau,i}^{(P)}),$ (9)
where $y_{\tau,i}^{(P)}$ are one-hot encoded vectors.
Previous work has extended equation 8 to the GMM setting, when the components are axis-aligned Gaussians, i.e., $\Sigma_{i}^{(P)}=\text{diag}(\sigma_{i}^{(P)})$. In this case,
$\mu_{i,\tau+1}^{(P)}=(1-\alpha)\mu_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\mu_{i,\tau}^{(P)}),$ (10)
$\sigma_{i,\tau+1}^{(P)}=(1-\alpha)\sigma_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\sigma_{i,\tau}^{(P)}),$
There are a few limitations with equations 8 and 10. First, the empirical iterations do not scale well to large datasets. Indeed, they assume the availability of all samples of $Q_{k}$ per iteration. Second, the GMM iterations in equation 10 are restricted to axis-aligned GMM s. This paper addresses these gaps using gradient flows.
2.4 Gradient Flows
In the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional. A gradient flow is the solution to the differential equation,
$\begin{cases}\dot{x}(t)=-\nabla F(x(t))&\text{for }t>0,\\ x(0)=x_{0},\end{cases}$ (11)
where $x_{0}$ is the initial condition. In $\mathcal{P}_{p,\text{ac}}(\Omega)$, gradient flows correspond to the continuity equation,
$\partial_{t}P_{t}=-\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})),$ (12)
where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures. Usually, these functionals take the following form [santambrogio2015optimal],
$\mathbb{F}(P)=$ $\underbrace{\int G(P(z))dz}_{\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\mathbb{V}(P)}+$ (13)
$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\mathbb{U}(P)},$
where $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ are the internal, potential, and interaction energies, respectively. Here, $G:\mathbb{R}\rightarrow\mathbb{R}$ is convex and superlinear, and $V:\Omega\rightarrow\mathbb{R}$ and $U:\Omega^{2}\rightarrow\mathbb{R}$ are convex and sufficiently smooth. Henceforth, we denote $\mathbb{F}^{\star}=\text{inf}_{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}\mathbb{F}(P)$.
3 Wasserstein Barycenters as Gradient Flows
In this section, we describe a new method for computing empirical and Gaussian mixture Wasserstein barycenters. Our main idea is using the functional:
$\mathbb{F}(P)=\mathbb{B}_{\mathcal{Q}}(P)+\mathbb{G}(P)+\mathbb{V}(P)+\mathbb{U}(P),$ (14)
where $\mathbb{B}_{\mathcal{Q}}(P)$ is the barycenter objective defined in equation 7, and $\mathbb{G},\mathbb{V}$ and $\mathbb{U}$ are the energies defined in equation 13.
3.1 Empirical Flow
In the empirical case, the barycenter is approximated through a finite set of particles (see equation 1) $\{z_{i}^{(P)}\}_{i=1}^{n}$. Hence, we may rewrite equation 14 as,
$\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{min}}\,\underbrace{\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}d(z_{i}^{(P)},z_{j}^{(Q_{k})})^{p}}_{\mathbb{B}_{\mathcal{Q}}(\hat{P})}$+
$\underbrace{\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)})}_{\mathbb{V}(\hat{P})}+\underbrace{\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)})}_{\mathbb{U}(\hat{P})}$ (15)
Here, we omit the internal energy term, $\mathbb{G}(P)$, as it is not defined for empirical measures. We solve the previous optimization problem through a block-coordinate descent strategy, that is,
$\gamma_{k}$ $=\underset{\gamma\in\Gamma(\hat{P}_{\tau},\hat{Q}_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{i,j}d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{p},$
$z_{\tau+1,i}^{(P)}$ $=z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}).$ (16)
In equation 16, $\gamma_{k}$ is used to compute $\mathbb{B}_{\mathcal{Q}}$ and is updated for each $\tau$. We provide in Algorithm 1 the pseudo-code for this strategy. Our algorithm assumes access to $Q_{k}$ through sampling, that is, we do not necessarily have access to all samples of $Q_{k}$ at once.
Algorithm 1 Empirical barycenter Wasserstein flow using Gradient Descent.
[…]
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
5:  for $k\leftarrow 1$ to $K$ do
6:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i.i.d. from $Q_{k}$.
7:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{p}(\hat{P}_{\tau},\hat{Q}_{k})^{p}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\tau})$
10:end for
Remark 3.1.
In line 5 of Algorithm 1, $\hat{Q}_{k}$ denotes the empirical measure (c.f., equation 1) supported on the mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ sampled from $Q_{k}$.
3.2 Gaussian Mixture Flow
In the case of GMMs, we flow the parameters $\theta=\{(\pi_{k},\mu_{k},\Sigma_{k})\}_{k=1}^{K}$ of $P_{\theta}=\sum\pi_{i}\mathcal{N}(\cdot|\mu_{i},\Sigma_{i})$. Similarly to equation 16, we have:
$\omega_{k}$ $=\underset{\omega\in\Gamma(\pi,\pi_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\omega_{i,j}\mathbb{W}_{2}(P_{\theta_{\tau},i},Q_{k,j})^{2},$
$\theta_{\tau+1,i}$ $=\theta_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}}).$ (17)
Here, $\omega_{k}$ is a transport plan between GMM components. In general, the energy functionals do not have closed form with respect to the GMM parameters. However, we can estimate them using Monte-Carlo and the reparametrization trick, for instance,
$z_{i}$ $=L_{i}\varepsilon+\mu_{i},\quad i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id}),$ (18)
This strategy allows us to propagate the gradients to the parameters of $P_{\tau}$. We show the overall approach in Algorithm 2.
Algorithm 2 GMM barycenter flow using Gradient Descent.
[…]
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  Sample $\{z_{i}=L_{i}\varepsilon+\mu_{i}:i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id})\}$,
5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{G}(\hat{P}_{\tau})+\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
6:  for $k\leftarrow 1$ to $K$ do
7:   $\mathbb{F}(P_{\theta_{\tau}})\leftarrow\mathbb{F}(P_{\theta_{\tau}})+\lambda_{k}\mathbb{MW}_{2}(P_{\theta_{\tau}},Q_{k})^{2}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}})$
10:end for
Covariance matrix parametrization. Optimizing $\mathbb{F}$ with respect to $\Sigma_{k}$ is challenging, due to the symmetric positive definite constraint. We enforce this constraint by writing $\Sigma_{k}=L_{k}L_{k}^{T}$, where $L_{k}$ is the lower-triangular Cholesky factor. There are two advantages to this strategy. First, optimizing with respect to $L_{k}$ is empirically more stable than $\Sigma_{k}$. Second, we can use $L_{k}$ directly in the reparametrization trick in equation 18.
3.3 Flows over Joint Measures
In [cuturi2014fast], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work. As we show in our experiments, in machine learning applications one encounters $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Whenever we deal with labeled measures (i.e., $z=(x,y)$), we adopt,
$d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (19)
where $\beta\geq 0$ is a parameter that balances the feature distance terms and the label distance terms.
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in [montesuma2021wasserstein] and [alvarez2021dataset], is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
$y_{i,c}^{(P)}=\text{softmax}(\ell_{i,1}^{(P)},\cdots,\ell_{i,n_{c}}^{(P)})=\dfrac{\exp(\ell_{i,c}^{(P)})}{\sum_{c=1}^{n_{c}}\exp(\ell_{i,c}^{(P)})},$
thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$. From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
For GMMs, we equip their components with labels $\nu_{k}\in\Delta_{n_{\text{classes}}}$. Using the ground-cost in equation 27, we are able to write the $\mathbb{MW}_{2}(P,Q)^{2}$ as,
$\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}^{\star}(\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\beta\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}^{2}),$
where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal. This approach was first proposed in [montesuma2024lighter], but the authors relied on heuristic arguments to justify this modeling choice. We present the following proposition justifying the label term,
Proposition 3.1.
Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be two GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$. Let the ground cost $c$ be,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
Functionals for joint measures. One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by [alvarez2021dataset] for transfer learning problems. Here we propose the following functionals,
$V_{E}(z)$ $=-\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (20)
$U_{R}(z,z^{\prime})$ $=\begin{cases}h(d(x,x^{\prime}))&\text{if }y\neq y^{\prime},\\ 0&\text{otherwise}\end{cases},$ (21)
where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below. For instance, in our experiments we use the hinge loss, $h(u)=\text{max}(0,\text{margin}-d)$, where $\text{margin}\geq 0$ is a margin parameter. Equations 20 and 21 correspond to entropy and repulsion terms. The first functional penalizes barycenters that have fuzzy labels. The second functional encourages classes to be well separated.
3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [santambrogio2017euclidean, Section 4.4]. As a result, we are minimizing a non-convex functional. Henceforth, we focus on $p=2$, as this is the setting where most results are available. Explicit constants and proofs are available in Section 4 in the Appendix.
Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [polyak1964gradient, karimi2016linear],
$\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (22)
which is covered in [chewi2024statistical]. Under this assumption, we have the following convergence result,
Theorem 3.1.
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball on $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Assume the inequality 47. Then, the following holds,
$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,m}}{n}}\end{aligned},$ (23)
where the expectation on the l.h.s. is taken with respect to samples from $Q_{k}$. The constants $C_{R}$ and $C_{d,m}$ depend on the radius $R$, number of barycenter samples $m$, and dimensions $d$.
The right hand side of inequality 23 is composed of two terms. The first term comes from the PL inequality, and covers the convergence towards the minimizer of $\mathbb{B}$. The second term covers the error of empirical approximations. In this sense, letting $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximation.
Next, we derive a new result for the convergence of the GMM gradient flow (cf. Algorithm 2). Here, we assume $P$ and $Q_{k}\in\mathcal{Q}$ are GMMs. Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm. Our result relies on [boite2025differentiable, Theorem 3.1]. Therefore, we assume,
  1. 1.
    bounded GMM parameters, i.e., $\lVert\mu_{i}^{(Q_{k})}\lVert_{2}\leq R_{\mu}$, $\sqrt{\text{Tr}(\Sigma_{i}^{(Q_{k})})}\leq R_{\Sigma}$ $\forall i,\forall k$ (resp. $\hat{\mu}_{i}^{(Q_{k})}$ and $\hat{\Sigma}_{i}^{(Q_{k})}$, the estimated parameters); furthermore,
  2. 2.
    Empirical convergence rates, i.e., $\mathbb{E}[\lVert\pi^{(Q_{k})}-\hat{\pi}^{(Q_{k})}\rVert_{1}]\leq\rho_{\pi}$, $\mathbb{E}[\lVert\mu_{i}^{(Q_{k})}-\hat{\mu}_{i}^{(Q_{k})}\rVert_{2}]\leq\rho_{\mu}$, and $\mathbb{E}[d_{\text{Bures}}(\Sigma_{i}^{(Q_{k})},\hat{\Sigma}_{i}^{(Q_{k})})]\leq\rho_{\Sigma}$, $\forall i$ and $\forall k$.
Here the expectations are taken with respect to samples drawn from $Q_{k}$. Under these conditions we have,
Theorem 3.2.
Let $P$ and $Q$ be two labeled GMMs with bounded parameters and satisfying the empirical convergence rates. Then, the following holds,
$\mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}\tau}}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{\lambda,\mathcal{Q}},$
where $C_{\lambda,\mathcal{Q}}$ is a constant that depends on the coordinates $\lambda$ and GMMs in $\mathcal{Q}$.
Here $C_{\lambda,\mathcal{Q}}$ depends on the $(\pi^{(Q_{k})},\mu^{(Q_{k})},\Sigma^{(Q_{k})})$ of each GMM, but not on $\nu^{(Q_{k})}$. Indeed, since we fit a GMM per class, the labels are estimated exactly (i.e., $P_{i,x}$ and $\hat{P}_{i,x}$ belong to the same class).
4 Experiments
We divide our experiments into a toy example (Section 4.1) to illustrate our methods, and multi-source domain adaptation experiments (Section 4.2). Due to length constraints, we include additional experiments in Sections 5 and 6 in the Appendix. In particular, in Table 1 in the Appendix, we show a running-time comparison demonstrating the scalability of our method.
4.1 Toy Example
In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$. This example has been used in several state-of-the-art Wasserstein barycenter works [korotin2021continuous, visentin2025computing]. Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold. We show a summary of our results in Figure 3.
Figure 2: Location-scatter family generated by a Swiss-roll measure, $Q_{0}$. Each measure $Q_{k}=T_{k,\sharp}Q_{0}$, for $T_{k}(x)=A_{k}x+b_{k}$.
First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h). Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our WGF algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
(a) CW2B [korotin2021continuous].
(b) NOT [kolesov2024estimating].
(c) Discrete [cuturi2014fast].
(d) WGF.
[…]
(g) U-NOT [gazdieva2024robust].
(h) NormFlow [visentin2025computing].
(i) Discrete [montesuma2023multi].
(j) WGF.
[…]
Figure 3: Comparison between Wasserstein barycenter solvers. Colored scatter plots indicate labeled barycenters. For each solver, we compute the Wasserstein distance between its solution $\hat{P}$ and the ground-truth $P^{\star}$, shown in the title of each sub figure (best seen on screen). Overall, using label information leads to barycenters that better approximate the ground-truth barycenter.
Second, we experiment with integrating labels in the ground cost as described in Section 3.3. These methods are shown in Figures 3 (i) through (l). In all cases, integrating labels produces barycenters that are closer to the ground-truth. We conclude that using the labels gives a strong inductive bias in the barycenter computation, which helps explain the gain in performance of labeled barycenters in the next section.
4.2 Multi-Source Domain Adaptation
Problem Formulation. One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [kouw2019review, sun2015survey, pan2009survey]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk in the target domain measure,
$\mathcal{R}_{Q_{T}}(h)=\mathbb{E}_{(x,y)\sim Q_{T}}[\mathcal{L}(y,h(x))],$
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data. In addition, similar to [montesuma2021wasserstein, montesuma2024lighter], we align the barycenter with the target through OT [courty2016optimal]. More details are available in Section 6 in the Appendix.
Benchmark Backbone # Samples # Domains # Dim. # Classes
Office31 ResNet50 [he2016deep] 3287 3 2048 31
BCI-CIV-2a CBraMod [wang2024cbramod] 5184 10 200 4
TEP CNN [montesuma2024benchmarking] 17289 6 128 29
Office Home ResNet101 [he2016deep] 15500 4 2048 65
ISRUC CBraMod [wang2024cbramod] 89240 100 512 5
Table 1: Overview of benchmarks used for domain adaptation, sorted by number of samples.
Experimental Setting. We run our experiments on five benchmarks: Office 31 [saenko2010adapting], Office Home [venkateswara2017deep], BCI-CIV-2a [brunner2008bci], ISRUC [khalighi2013automatic], and TEP [montesuma2024benchmarking]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. These methods are: discrete barycenters [cuturi2014fast, Algorithm 2], normalizing flows [visentin2025computing], continuous 2-Wasserstein barycenters [korotin2021continuous], neural OT barycenters [kolesov2024estimating], and unbalanced neural OT barycenters [gazdieva2024robust]. We also include labeled barycenter strategies [montesuma2023multi, montesuma2024multi]. A brief review of these methods is available in Section 2 in the Appendix. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT [turrisi2022multi], DaDiL-R and E [montesuma2023multi], and GMM-DaDiL [montesuma2024lighter]. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchamrk), and we measure the classification accuracy, i.e., the percentage of correct predictions.
Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP
Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN
Source-Only - 86.40 75.95 50.30 76.63 78.48
WJDOT [turrisi2022multi] - 86.80 76.59 N/A 76.95 86.13
DaDiL-R [montesuma2023multi] - 89.91 77.86 53.41 74.68 86.14
DaDiL-E [montesuma2023multi] - 89.79 78.14 N/A 75.89 85.87
GMM-DaDiL [montesuma2024lighter] - 90.63 78.81 57.10 75.47 86.85
Discrete [cuturi2014fast] 81.94 70.42 57.38 70.67 83.81
NormFlow [visentin2025computing] 85.91 76.73 55.45 78.04 82.89
CW2B [korotin2021continuous] 86.37 76.44 57.25 75.84 85.83
NOT [kolesov2024estimating] 86.22 75.36 57.29 76.39 84.93
U-NOT [gazdieva2024robust] 86.97 76.85 57.17 77.06 85.43
Discrete [montesuma2023multi] 87.93 77.09 57.43 78.20 86.09
GMM [montesuma2024lighter] 88.54 77.87 57.04 74.58 84.67
WGF (ours) 88.14 77.83 57.50 79.78 86.21
WGF-GMM (ours) 89.31 78.71 57.35 78.72 86.77
Table 2: Average classification accuracy on target domains. In total, we compare eight methods against our WGF strategy. For barycenter solvers, we indicate by ✓methods that compute barycenters on $\Omega=\mathcal{X}\times\mathcal{Y}$. Bold numbers represent the best barycenter method, and underlined numbers represent second best.
Main results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Section 6 in the Appendix. Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods. We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter]. Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
Figure 4: t-SNE [maaten2008visualizing] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark. Colors represent different classes from $1$ to $65$. Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
Visualization. We now visualize the effect of adding the repulsion interaction energy functional on the Office-Home benchmark. Adding $\mathbb{U}$ leads to classes that are well separated. For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$. This choice encourages class clusters to live in orthogonal sub-spaces of the latent space. We show a comparison of the obtained barycenters in Figure 4. Overall, the classes tend to have less overlap when the interaction energy is added to the barycenter functional.
Ablation. We ablate the effect of $\mathbb{V}$ and $\mathbb{U}$ in the Office Home benchmark. For this benchmark, we use $\mathbb{V}(P)=\mathbb{W}_{2}(\hat{P},\hat{Q}_{T})^{2}$. The intent here is to isolate the contribution of each component to the state-of-the-art results in Table 2. Our results are summarized in Table 3, where we further compare our results with other related barycenter methods. In the empirical setting, the use of Algorithm 1 alone already presents a gain in performance compared to its discrete counterpart. Overall, the combination of $\mathbb{B}+\mathbb{V}+\mathbb{U}$ consistently gives the best performance, showing the importance of adding underlying structure to the Wasserstein barycenter in domain adaptation.
Method $\mathcal{X}\times\mathcal{Y}$ $\mathbb{B}$ $\mathbb{B}+\mathbb{V}$ $\mathbb{B}+\mathbb{U}$ $\mathbb{B}+\mathbb{V}+\mathbb{U}$
Discrete [cuturi2014fast] 70.42 N/A N/A N/A
Discrete [montesuma2023multi] 77.09 N/A N/A N/A
WGF (ours) 77.48 77.07 77.08 77.83
GMM [montesuma2024lighter] 77.87 N/A N/A N/A
WGF-GMM (ours) 77.28 78.28 78.12 78.71
Table 3: Ablation of average target domain classification accuracy on the Office Home benchmark.
5 Conclusion
In this paper we have introduced a new framework for computing Wasserstein barycenters, based on Wasserstein gradient flows [ambrosio2008gradient, santambrogio2017euclidean]. Our approach addresses scalability issues in discrete barycenter solvers [cuturi2014fast], and is capable of regularizing the underlying barycenter through internal, potential, and interaction functionals. We presented gradient flow Algorithms 1 and 2, for empirical and Gaussian mixture measures, respectively. We further proved convergence guarantees based on the PL inequality.
We empirically tested our method against existing discrete [cuturi2014fast, montesuma2023multi] and neural-network [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing]-based barycenter solvers on computer vision, neuroscience, and chemical engineering benchmarks, demonstrating that our proposed methods consistently achieve state-of-the-art performance. Our empirical findings (Table 2) further demonstrate that incorporating label information in the OT objective is key for domain adaptation. This finding highlights an interesting gap in the literature, where neural net-based solvers are not capable of exploiting this information, and thus have sub-optimal performance in domain adaptation. We leave an investigation of this issue for future work.
Computing Wasserstein Barycenters through Gradient Flows
Appendix A Introduction
The goal of this supplementary material is to provide additional background on the main paper, as well as to provide missing proofs, additional experiments, and the complete experimental setting. These details were left out of the main paper due to length constraints.
Organization. We divide this supplementary material as follows. Section B provides a brief survey on the methods for computing Wasserstein barycenters. Section C.1 provides additional background on OT. Section C.2 covers additional details on the potential, internal, and interaction energies, as well as the barycenter objectives and their gradients. Section D provides missing proofs. Finally, section E and F provide additional experiments.
Appendix B A Brief Survey on Wasserstein Barycenters
B.1 Introduction
Given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, the Wasserstein barycenter problem was first introduced by [agueh2011barycenters] as the following minimization problem,
$P^{\star}=\underset{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr\}.$ (24)
This minimization problem is the analogous, in Wasserstein space, to the Euclidean barycenter problem $x\mapsto\sum_{k}\lambda_{k}d(x,x_{k})^{2}$, which is made possible since $(\mathcal{P}_{2,\text{ac}}(\Omega),\mathbb{W}_{2})$ is a metric space. Solving problem 24 is, in general, challenging. Except for the Gaussian case, there are no closed-form solutions. In the following, we review a few strategies for computing Wasserstein barycenters.
B.2 Discrete Barycenter Solvers
Discrete barycenter solvers are based on the empirical approximation of the underlying measures, that is,
$\hat{P}(z)=\sum_{i=1}^{n}a_{i}\delta(z-z_{i}^{(P)}).$ (25)
Here, $\{z_{i}^{(P)}\}_{i=1}^{n}$ is the support of $\hat{P}$ and $a=\{a_{i}:a_{i}\geq 0\text{, and }\sum a_{i}\}_{i=1}^{n}$ are the sample weights. There are two prevalent strategies when discretizing $P$. First, one can fix a grid/bins over the underlying space $\Omega$, and compute $a_{i}=P(z_{i}^{(P)})$ as the bin weight. This strategy is equivalent to estimating the density of $P$ over $\Omega$. Second, we can sample $z_{i}^{(P)}\overset{\text{i.i.d.}}{\sim}P$, in which case $a_{i}=\nicefrac{{1}}{{n}}$. In this paper, we deal primarily with free-support Wasserstein barycenters. Arguably, the first free-support algorithm was proposed by [cuturi2014fast]. Their strategy comes from plugging equation 25 into 24,
$Z^{(P),\star}=\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}\lVert z_{i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}\biggr\}.$ (26)
At this point, one should compare equation 26 with equation 17 in the main paper. Cuturi and Doucet [cuturi2014fast] propose solving this equation via a block coordinate descent strategy. Indeed, assume $\gamma_{k}=\text{OT}(P,Q_{k})$ has been calculated (e.g., through linear programming), then,
$\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}(z_{i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}.$
As a consequence, $\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=0$ leads to an update rule in terms of the barycentric projection of $P$ to $Q_{k}$, which is a discrete approximation of the Monge map [deb2021rates].
An important question comes when the samples from $P$ are feature-label joints. This is question was first raised by [montesuma2023multi], who proposed to use the ground-cost,
$d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (27)
where $y$ and $y^{\prime}$ are one-hot encoded labels, and $\beta\geq 0$ strikes a balance between the features and labels terms. As we show in the main paper, since this ground-cost is a metric, computing the OT cost with $d^{2}$ as the ground-cost is equivalent to computing the 22-Wasserstein distance on $\mathcal{X}\times\mathcal{Y}$.
The first paper to remark these computations was [montesuma2023multi], who proposed to couple these mappings,
$x_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(x_{i,\tau}^{(P)})\text{, and, }y_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)}).$ (28)
These iterations have a few nice properties with respect to the label term. First, the mappings $T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)})$ correspond to label propagation terms first defined in [redko2019optimal]. Second, as a consequence of the first point, the iterations respect the constraint $y_{i,\tau}^{(P)}\in\Delta_{n_{\text{classes}}}$. Third, these iterations can be understood as fixed-point iterations in the sense of [alvarez2016fixed]. The authors use this algorithm as the building block of their dictionary learning strategy.
The main drawback of the iterations in equation 28 is when the input measures have imbalanced classes. As we show in the additional experiments in Section E.2, OT, due the mass preservation constraints, is forced to match samples from different classes, leading to deformed barycenters with fuzzy labels (see Figure 8).
B.3 Neural Barycenter Solvers
In this section, we cover a collection of methods that use neural networks as the foundation of computing Wasserstein barycenters. These methods are oftentimes referred to as ”continuous”, thus relying on the idea that the calculated barycenter is a continuous measure. Note that this is different than claiming that these barycenters solve a continuous problem. In fact, most methods are formulated in terms of expectations of the involved measures, which are approximated through mini-batching.
Generative Modeling. One of the first methods was proposed by [cohen2020estimating], who proposed viewing equation 24 through the lens of generative modeling. Let $g_{\theta}$ denote a neural network with parameters $\theta$. Then,
$\theta^{\star}=\underset{\theta\in\Theta}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P_{\theta},Q_{k})^{2},$
where $P_{\theta}=g_{\theta,\sharp}P_{0}$ is the push-forward of a latent measure $P_{0}$ by the generator $g_{\theta}$. Therefore, the Wasserstein barycenter is parametrized via the weights of $g_{\theta}$. In comparison, methods such as [cuturi2014fast] and [montesuma2023multi] parametrize the barycenter through its samples. The underlying idea is sampling from each $Q_{k}$ and $P_{0}=\mathcal{N}(0,\sigma^{2}I)$.
Input Convex Neural Nets. More refined settings exploit the properties of the 2-Wasserstein distance with the squared Euclidean ground-cost. Indeed, note that,
$\dfrac{1}{2}\mathbb{W}_{2}(P,Q)^{2}=\underbrace{\int\dfrac{\lVert z\rVert_{2}}{2}dP(z)+\int\dfrac{\lVert z\rVert_{2}}{2}dQ(z)}_{\mathbb{C}(P,Q)}-\underset{f,g\in\Phi}{\text{min}}\,\int f(z)dP(z)+\int g(z)dQ(z),$ (29)
where $(f,g)\in\Phi=\{f(x)+g(y)\leq\lVert x-y\rVert_{2}^{2}\}$ are called potentials. The main insight of [makkuva2020optimal] is rewritting the minimization problem in the r.h.s. of equation 29 as,
$\underset{f\in\text{CVX}}{\text{sup}}\,\underset{g\in\text{CVX}}{\text{inf}}\,\underbrace{-\int f(z)dP(z)-\int(\langle z,\nabla g(z)\rangle-f(\nabla g(z)))dQ(z)}_{V_{P,Q}(f,g)}$
Based on these equations, [fan2020scalable] propose solving,
[…]
Here, samples in the Wasserstein barycenter are generated through the forward pass $h(z)$, where $z\sim P_{0}=\mathcal{N}(0,\sigma^{2}I)$. This is a min-max-min problem, which, as with many problems involving neural nets (e.g., generative adversarial nets [goodfellow2014generative]), is subject to saddle points.
An important aspect of equation 30 is the minimization with respect convex functions $f_{k}$ and $g_{k}$. Indeed, these are the potentials of the transportation problem between the barycenter $P$ and each $Q_{k}$. More specifically, In light of Brenier’s theorem [brenier1991polar], $\nabla g_{k}$ is the Monge map from $Q_{k}$ to the barycenter $P$. These elements motivate [fan2020scalable] using ICNNs, as in [makkuva2020optimal], for the optimization problem in equation 30.
Continuous 2-Wasserstein Barycenter. We now cover a family of methods relying on the fixed-point view of Wasserstein barycenters introduced by [alvarez2016fixed]. Indeed, these authors show that the Wasserstein barycenter is a fixed point of the operator $\Psi:P\mapsto(\sum_{k}\lambda_{k}T_{P\rightarrow Q_{k}})_{\sharp}P$. This is actually the principle behind most of the discrete methods described in the previous section. Now, with respect to OT potentials, if $\Psi(P)=P$, then one has,
$\sum_{k=1}^{K}\lambda_{k}T_{Q_{k}\rightarrow P}^{\star}(z)=x\overset{(a)}{\implies}\sum_{k=1}^{K}\lambda_{k}\nabla g_{k}(z)=z\overset{(b)}{\implies}\sum_{k=1}^{K}\lambda_{k}g_{k}(z)=\dfrac{\lVert z\rVert_{2}^{2}}{2}+c$
where implication (a) follows from Brenier’s theorem, and $g_{k}$ is the potential of the transportation problem from $Q_{k}$ to $P$, and implication (b) follows by integration. Here, since $c$ is an integration constant, one can conveniently take $c=0$. Now, denote $\text{Congruent}(\lambda)=\{(g_{1},\cdots,g_{K}):\sum_{k}\lambda_{k}g_{k}(z)=\nicefrac{{\lVert z\rVert_{2}^{2}}}{{2}}\}$ be the set of congruent potentials $(g_{1},\cdots,g_{K})$ with respect to coordinates $\lambda$. Using the dual formulation in equation 29 and restricting the set of potentials to congruent potentials,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]+\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\biggr(\sum_{k=1}^{K}\lambda_{k}(\mathbb{E}_{z\sim Q_{k}}[f_{k}(z)]+\mathbb{E}_{z\sim P}[g_{k}])(z)\biggr).$
Now, taking the expectation on the congruence constraint, we have,
$\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim P}[g_{k}(z)]=\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr],$
which allowed [korotin2021continuous] to simplify $\mathbb{B}_{\mathcal{Q}}(P)$ into,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}[g_{k}(z)],$ (31)
which does not involve terms with respect to the barycentric measure $P$. While the approach of [korotin2021continuous] simplifies the barycenter objective to expectations with respect to the known measures $Q_{k}$, it introduces a challenging constraint – the congruence of potentials. This constraint is challenging, as it is a functional inequality over the whole space $\Omega$. Furthermore, it adds to the already constrained setting of fitting convex functions. [korotin2021continuous] proposes solving it approximately, by penalizing potentials that do not meet the congruence condition. There is a circular dependence here, as one needs an estimate of the Wasserstein barycenter to estimate the constraint violation.
Neural (Weak) Optimal Transport. Weak OT [gozlan2017kantorovich, backhoff2019existence] comes from the interpretation of the Kantorovich formulation as,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}C(z,\gamma(\cdot|z))dP(z)\text{, where, }C(z,Q)=\int_{\Omega}c(z,z^{\prime})dQ(z^{\prime}),$ (32)
this formulation allows for more general costs $C:\Omega\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$, i.e., costs that take a sample and a probability measure as inputs. In this sense, [kolesov2024estimating] extends equation 31 to,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma(\cdot|z)}(g_{k}(z^{\prime}))]\biggr),$ (33)
which holds for congruent potentials $\sum_{k}\lambda_{k}g_{k}=0$. As with equation 31, $\mathbb{B}_{\mathcal{Q}}$ does not depend explicitly on the unknown barycentric measure $P$. As such, the authors compute the barycenter through a max-min optimization problem,
$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma_{k}(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}(g_{k}(z^{\prime}))]\biggr)\biggr\}.$
This extension opens up new possibilities for weak OT barycenters. For instance,
$C_{\text{classical}}(x,\gamma)=\dfrac{1}{n}\sum_{j=1}^{n}c(z_{i}^{(P)},z_{j}^{(Q)})\quad C_{\text{KL}}(x,\gamma)=\dfrac{1}{n}\sum_{j=1}^{n}c(z_{i}^{(P)},z_{j}^{(Q)})+\epsilon\text{KL}(\mathcal{N}(\mu(z),\sigma(z))|P_{0}),$
where $C_{\text{classical}}$ allows us to retrieve the original OT problem. $C_{\text{KL}}$ is an alternative, when modeling $\gamma(\cdot|z)=\mathcal{N}(\mu(z),\sigma(z))$. On top of this reformulation, the authors in [kolesov2024estimating] propose approximating $\gamma(\cdot|z)$ through a mapping, as in [korotin2022neural], that is,
$\gamma(\cdot|z)=T(z,s),\text{ subject to }s\sim\mathbb{S},$
where $\mathbb{S}$ is some prior measure, such as $\mathbb{S}=\mathcal{N}(0,\text{I})$. In this context, [gazdieva2024robust] extends this idea in the semi-unbalanced OT setting [liero2018optimal], where one of the mass preservation constraint of one of the marginals of $\gamma$ is relaxed. More specifically, given $\xi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$,
$D_{\xi}(P||Q)=\int_{\Omega}\xi\biggr(\dfrac{P(z)}{Q(z)}\biggr)dQ(z)\text{ if }P\ll Q\text{ and }+\infty\text{ otherwise},$
where $P\ll Q$ means that $P$ is absolutely continuous with respect to $Q$. We denote $\bar{\xi}(z^{\prime})=\text{sup}_{z\in\mathbb{R}}\{zz^{\prime}-\xi(z)\}$ to be its convex conjugate. With that in mind, [gazdieva2024robust] shows, in Theorem 1 and Corollaries 1 and 2, that minimizing,
$\underset{P\in\mathcal{P}_{2,\text{ac}}}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\biggr(\int_{\Omega}\int_{\Omega}c(z,z^{\prime})d\gamma(z,z^{\prime})+D_{\xi}(\gamma(\cdot|z)||Q_{k})\biggr),$
is equivalent to the following max-min problem,
$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}\biggr[-\bar{\xi}_{k}\biggr(-\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}[c_{k}(z,z^{\prime})-g_{k}(z^{\prime})]\biggr)\biggr]\biggr)\biggr\}.$
Normalizing Flows. This approach was proposed by [visentin2025computing], and relies on the analysis of [gangbo1996geometry]. Their main idea is re-writing the Wasserstein distance in terms of functions $f,g$,
$\mathbb{W}_{p}(P,Q)^{p}=\underset{f\in\text{Borel}(P_{0},P),g\in\text{Borel}(P_{0},Q)}{\text{inf}}\,\lVert f-g\rVert_{L_{p}(P_{0})}^{p},$ (34)
where $\text{Borel}(P,Q)$ is the set of Borel functions such that $f_{\sharp}P=Q$. Naturally, a solution $(f^{\star},g^{\star})$ to the previous problem gives the transport map $T=g^{\star}\circ f^{-1}$, where $f^{-1}$ is the inverse of $f^{\star}$. Furthermore, $P_{0}$ is an arbitrary latent measure, which further links the approach with generative modeling (i.e., take $P_{0}=\mathcal{N}(0,\text{I})$).
Based on equation 34, the insight of [visentin2025computing] is modeling $f$ and $g$ through a conditional function $f(z,s)$, $s\in\{0,1\}$. The Wasserstein barycenter problem is a straightforward generalization of the $2$ measures case. For instance, the authors consider $S=\{1,\cdots,K\}$, which means that,
$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(g_{\sharp}P_{0},Q_{k})=\sum_{k=1}^{K}\lambda_{k}\lVert g-f(z,k)\rVert_{L_{p}(P_{0})}^{p},$
where $g$ is the push-forward of $P_{0}$ to the barycenter $P^{\star}$, and $f(z,k)$ is the push-forward of $Q_{k}$ to $P^{\star}$. As the authors show in [visentin2025computing, Theorem 3.3],
$g(z)=\sum_{k=1}^{K}\lambda_{k}f(z,k).$
The main challenge in [visentin2025computing] is formulating a conditional model $f:\Omega\times S$. As the authors discuss, one possibility would be using $K=|S|$ different neural nets for modeling $f(z,\cdot)$, which is similar to what is done, for instance, in the neural net methods. The authors take a different route and use conditional normalizing flows [kobyzev2020normalizing] to model these functions.
B.4 Future Directions
From the previous discussion on neural net-based methods for computing Wasserstein barycenters, we can identify a few common design choices. For instance, [korotin2021continuous, kolesov2024estimating] and [gazdieva2024robust] use $\mathcal{O}(K)$ neural nets for computing Wasserstein barycenters. This design choice results in a parametrization of the underlying barycenter that grows with the number of input measures. The only neural net-based method that differs from this idea is [visentin2025computing], which uses a conditional model $f:\Omega\times S$.
Meanwhile, as we cover in our experiments (C.f., Section 4 in the main paper), there is a clear advantage of incorporating labels in the ground-cost when these are available. Unfortunately, it is not clear how to integrate label information in the existing methods. For instance, one could consider using the continuous parametrization in terms of logits $\ell$, as we described in Section 3.3. However, there are a few caveats,
  1. 1.
    Normalizing flows require invertible transformations. However, the softmax operation for getting the labels, i.e., $y=\text{softmax}(\ell)$, is not invertible.
  2. 2.
    Using ICNNs, one can enforce the convexity of $z=(x,\ell)$. However, applying the softmax breaks the convexity requirement.
As a result, adapting neural net based methods for the joint $\mathcal{X}\times\mathcal{Y}$ space is challenging and not straightforward. We leave this question for future work.
Appendix C Additional Background
C.1 Calculus in Wasserstein Spaces
In this section, we consider $\Omega=\mathbb{R}^{d}$, $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$ and $p=2$ fixed. While our algorithm is more general than this setting, we use these restrictions as they are the most analyzed by current literature. We leave the theoretical results for general metrics and $p$ to future work.
Notation. For the following theoretical analysis, we use the following notation. Given a measurable map $f,g:\Omega\rightarrow\Omega$. Given a measure $P\in\mathcal{P}_{2,\text{ac}}(\Omega)$, we denote,
$\lVert f\rVert_{L_{2}(P)}=\int_{\Omega}\lVert f(x)\rVert_{2}^{2}dP(x).$
Likewise, we have a notion of inner product,
$\langle f,g\rangle_{P}=\int_{\Omega}\langle f(x),g(x)\rangle_{\Omega}dP(x).$
Definition C.1.
Given a functional $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)(\mathbb{R}^{d})\rightarrow\mathbb{R}$, we call $\delta\mathbb{F}(P):\mathbb{R}^{d}\rightarrow\mathbb{R}$, if it exists, the unique (up to additive constants) function such that
$\dfrac{d}{d\varepsilon}\mathbb{F}(P+\varepsilon\chi)\biggr|_{\epsilon=0}=\int\delta\mathbb{F}(P)d\chi$
for every perturbation $\chi$ such that, at least for $\varepsilon\in[0,\varepsilon_{0}]$, the measure $P+\varepsilon\chi\in\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$. This functional is called the first variation of $\mathbb{F}$ at $P$.
Following [chewi2024statistical], given a curve of densities $\{P_{t}\}_{t\geq 0}$, we can take $\chi=\partial_{t}\mathbb{F}(P_{t})$ so that $\partial_{t}\mathbb{F}(P_{t})=\int\delta\mathbb{F}(P_{t})\partial_{t}P_{t}$. From Riemannian geometry, we can define the Wasserstein gradient to $P_{t}$ as,
$\partial_{t}\mathbb{F}(P_{t})=\langle\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}),v_{t}\rangle_{P_{t}},$ (35)
where $v_{t}\in\mathcal{T}_{P_{t}}\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$. Note, here, that $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}):\Omega\rightarrow\Omega$ is a vector field. In comparison, the first variation $\delta\mathbb{F}$ is a scalar field. These two objects are connected by the following equation [chewi2024statistical, Proposition 5.10],
$\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla\delta\mathbb{F}(P),$ (36)
Meanwhile, recall that, from dynamic OT, the WGF satisfy the continuity equation,
$\partial_{t}P_{t}+\text{div}(P_{t}v_{t})=0,$
where $\text{div}(v)=\nabla\cdot v$ is the divergence of a vector field. Therefore, taking the direction of steepest descent of the functional $\mathbb{F}$, $v_{t}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$, leads to the following connection between $\partial_{t}\mathbb{F}(P_{t})$ and $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$,
$\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$ (37)
C.2 Wasserstein Gradient of Functionals
In this section, we give an overview on the Wasserstein gradient of the terms used in equation 16. We refer readers to [chewi2024statistical, Chapter 5] and [santambrogio2017euclidean, Section 4] for further details.
Potential Energy, $\mathbb{V}$. For reference, this functional corresponds to,
$\mathbb{V}(P)=\mathbb{E}_{z\sim P}[V(z)]=\int V(z)dP(z),$
for some potential function $V:\Omega\rightarrow\mathbb{R}$. From the definition of first variation,
$\mathbb{V}(P+\epsilon\chi)=\int V(z)dP(z)+\epsilon\int V(z)d\chi(z),$
then, taking the derivative with respect to $\epsilon$,
$\dfrac{d}{d\epsilon}\mathbb{V}(P+\epsilon\chi)=\int V(z)d\chi,$
which means that $V=\delta\mathbb{V}(P)$. As a consequence, the Wasserstein gradient of $\mathbb{V}$ is,
$\nabla\mkern-10.0mu\nabla\mathbb{V}=\nabla V.$
which is a vector field over $\Omega$. Since $\mathbb{V}$ is linear in $P$, it is convex. Furthermore, its continuity with respect to the weak convergence depends on the regularity of $V$. For instance, if $V$ is bounded, $\mathbb{V}$ is continuous. If $V$ is lower semi-continuous and bounded from below, then $\mathbb{V}$ is semi-continuous.
Internal Energy, $\mathbb{G}$. The internal energy takes the form,
$\mathbb{G}(P)=\int G(P(z))dz,$
where $P(z)$ is understood as the density of the measure $P$, with respect the Lebesgue measure. Let us assume $\chi$ admits a density with respect the Lebesgue measure as well. In this case,
$\mathbb{G}(P+\epsilon\chi)=\int G(P(z)+\epsilon\chi(z))dz.$
Taking the derivative under the integral sign,
$\dfrac{d}{d\epsilon}\mathbb{G}(P+\epsilon\chi)\biggr|_{\epsilon=0}=\int G^{\prime}(P(z))\chi(z)dz=\int G^{\prime}(P(z))d\chi(z),$
which means that $\delta\mathbb{G}=G^{\prime}(P)$ and $\nabla\mkern-10.0mu\nabla\mathbb{G}=\nabla G^{\prime}(P(z))$.
Interaction Energy, $\mathbb{U}$. Here, let us assume $U(z,z^{\prime})$ symmetric for simplicity,
$\mathbb{U}(P)=\int\int U(z,z^{\prime})dP(z)dP(z^{\prime}),$
hence,
$\mathbb{U}(P+\epsilon\chi)=\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})+2\epsilon\int\int U(z,z^{\prime})dP(z)d\chi(z^{\prime})+\mathcal{O}(\epsilon^{2}).$
Again, taking derivatives with respect $\epsilon$,
$\dfrac{d}{d\epsilon}\mathbb{U}(P+\epsilon\chi)\biggr|_{\epsilon=0}=2\int\biggr(\int U(z,z^{\prime})dP(z)\biggr)d\chi(z^{\prime}),$
which means that,
$\delta\mathbb{U}(P)=2\int U(z,z^{\prime})dP(z).$
When $U(z,z^{\prime})=h(z-z^{\prime})$ for an even $h$, $\int U(z,z^{\prime})dP(z)=\int h(z-z^{\prime})P(z)dz=h\star P$, i.e., the convolution of $h$ and $P$. Furthermore, $\nabla\mkern-10.0mu\nabla\mathbb{U}=(\nabla h)\star P$
Barycenter Functional, $\mathbb{B}$. For the barycenter functional, we can derive its first variation from the 22-Wasserstein distance. Assume there exists an OT map $T_{P\rightarrow Q}^{\star}$. In this case,
$\nabla\mkern-10.0mu\nabla\mathbb{W}_{2}(P,Q)^{2}=2(\text{Id}-T_{P\rightarrow Q}^{\star})$
Therefore, given $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$, it is straightforward to obtain:
$\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}(P)$ $=\nabla\mkern-10.0mu\nabla\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr)=2\sum_{k=1}^{K}\lambda_{k}(\text{Id}-T_{P\rightarrow Q_{k}}^{\star}).$
A Connection with [cuturi2014fast, Algorithm 2]. Let us now focus on the empirical case with an Euclidean ground-cost. Assume $\gamma^{\star}$ computed between $\hat{P}$ and $\hat{Q}$, which means that,
$\mathbb{W}_{2}(\hat{P},\hat{Q})^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}^{\star}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}\text{, and }\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}\lVert z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}.$
We can compute the Wasserstein gradient $\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}$ by computing $\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}$, that is,
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}(z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{\tau,i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{K}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})}.$
From this last equation, we can use the barycentric mapping definition,
$T_{P\rightarrow Q_{k}}^{\star}(z_{i,\tau}^{P})=n\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})},$ (38)
to write,
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\dfrac{2}{n}\biggr(z_{\tau,i}^{(P)}-\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{\tau,i}^{(P)})\biggr),$
which means we can write the gradient flow update rule as,
$z_{\tau+1,i}^{(P)}=\biggr(1-\dfrac{2\alpha}{n}\biggr)z_{i,\tau}^{(P)}+\dfrac{2\alpha}{n}\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{i,\tau}^{(P)}),$
Which is the update rule in [cuturi2014fast, Algorithm 2] with $\theta=\nicefrac{{2\alpha}}{{n}}$. Furthermore, we can also instantiate the fixed-point algorithm of [montesuma2023multi], for $\alpha=n/2$. We now highlight two differences of our work with respect these two strategies. First, our setting is more general. In fact, our algorithm is capable of handling ground metrics other than the squared Euclidean metric. Furthermore, our framework couples the barycenter objective with other functionals. Second, while [cuturi2014fast] assume that all samples of $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ are available at once, our algorithm assumes that these measures are available through sampling. This treatment is oftentimes called continuous in the literature [korotin2021continuous, korotin2022wasserstein], however, one still relies on samples from these measures rather than some truly continuous form.
C.3 Domain Adaptation
In this section, we offer a brief introduction to domain adaptation in classification problems which are the main application we explore in the main paper.
Statistical learning theory [vapnik1999overview] deals with the problem of learning a function $h:\mathcal{X}\rightarrow\mathcal{Y}$, such that it minimizes the risk of making a mistake, that is,
$h^{\star}=\underset{h\in\mathcal{H}}{\text{arg inf}}\,\biggr\{\mathcal{R}_{Q}(h)=\int_{\mathcal{X}\times\mathcal{Y}}\mathcal{L}(y,h(x))dQ(x,y)\biggr\},$ (39)
where $P\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$ is a probability measure, and $\mathcal{H}$ is a family of functions from $\mathcal{X}$ to $\mathcal{Y}$. $\mathcal{R}_{Q}(h)$ denotes the risk of $h$ under $Q$. Equation 39 interprets the learning problem as a minimization problem.
In practical scenarios, one has a dataset $\{x_{i}^{(Q)},y_{i}^{(P)}\}_{i=1}^{n}$, $(x_{i}^{(P)},y_{i}^{(Q)})\overset{i.i.d.}{\sim}P$. As a result, one can estimate the empirical risk, and consequently minimize it,
$\hat{h}=\underset{h\in\mathcal{H}}{\text{arg inf}}\,\biggr\{\hat{\mathcal{R}}_{Q}(h)=\dfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(y_{i}^{(Q)},h(x_{i}^{(Q)}))\biggr\},$ (40)
which is the foundation of the Empirical Risk Minimization (ERM) framework. In this sense, a solution $\hat{h}$ to equation 40 is said to generalize if it achieves low risk $\mathcal{R}_{Q}(h)$.
There is a fundamental assumption at play in the theory of generalization under the ERM framework. Indeed, both $\mathcal{R}_{P}$ and $\hat{\mathcal{R}}_{P}$ evaluate the risk under a fixed probability measure $P$. However, machine learning models are oftentimes confronted with data following different, but related probability measures [quinonero2022dataset], a process known as distribution shift.
The main tool for tackling distribution shift is domain adaptation. This problem considers now two datasets, namely, a labeled source dataset $\{x_{i}^{(Q_{S})},y_{i}^{(Q_{S})}\}_{i=1}^{n_{S}}$, and an unlabeled target dataset $\{x_{j}^{(Q_{T})}\}_{j=1}^{n_{T}}$. The goal is to obtain a function $h$ that achieves low risk under $Q$, with the available data. A natural extension to this problem is multi-source domain adaptation, where multiple source domains are available, i.e., $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, each with its own associated dataset $\{(x_{i}^{(Q_{k})},y_{i}^{(Q_{k})})\}_{i=1}^{n_{k}}$.
Source-Only Training. A straightforward baseline in domain adaptation is the source-only training. Let $h=g\circ f$ be the composition of a feature extractor $g:\mathcal{X}\rightarrow\mathcal{U}$ and a feature classifier $f:\mathcal{U}\rightarrow\mathcal{Y}$. We can train $h$ end-to-end by minimizing,
$\mathcal{L}(\theta)=\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathcal{L}(y_{i}^{(Q_{k})},f(g(x_{i}^{(Q_{k})}))).$
This training produces an embedding space $\mathcal{U}$ that is discriminative of the classes in the domain adaptation problem. Following previous work [montesuma2021wasserstein, montesuma2023multi, montesuma2024distillation], we perform adaptation in this space.
Wasserstein Barycenters and Domain Adaptation. The motivation for using Wasserstein barycenters in multi-source domain adaptation comes from a theoretical result of [redko2019optimal]. Indeed, by [redko2019optimal, Theorem 2],
$\mathcal{R}_{Q_{T}}(h)\leq\mathcal{R}_{Q_{S}}(h)+\mathbb{W}_{1}(\hat{Q}_{S},\hat{Q}_{T})+\mathcal{O}(n_{S}^{-1/2}+n_{T}^{-1/2})+\mathbb{R}(Q_{S},Q_{T}),$ (41)
where $\mathbb{R}(Q_{S},Q_{T})=\text{inf}_{h\in\mathcal{H}}\mathcal{R}_{Q_{S}}(h)+\mathcal{R}_{Q_{T}}(h)$ is the joint risk. This result establishes that when source and target measures are close, their risks are close as well. Now, assume equation 41 holds for each pair $(Q_{k},Q_{T})$, then, summing over $k$ weighted by $\lambda_{k}$,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
using the triangle inequality on $\mathbb{W}_{1}(\hat{Q}_{k},\hat{Q}_{T})$
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},P)+\mathbb{W}_{1}(P,\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
A reasonable choice to make this bound tight is choosing $P$ that minimizes $P\mapsto\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},P)+\mathbb{W}_{1}(P,\hat{Q}_{T})$.
From $p=1$ to $p>1$. The argument for $p>1$ is a bit more intricate. Since $\mathbb{W}_{1}$ is the weakest of the Wasserstein distances [villani2008optimal, Remark 6.6], we can directly move from $p=1$ to $p>1$,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
which, from the triangle inequality, results in,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)+\mathbb{W}_{p}(P,\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$ (42)
The term $\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)+\mathbb{W}_{p}(P,\hat{Q}_{T})$ does not correspond to the $p$-Wasserstein barycenter due the missing power of $p$. We can retrieve it through Young’s inequality for products. Let $a,b>0$ be real numbers, and $p,q>1$. Then $ab\leq\frac{a^{p}}{p}+\frac{b^{q}}{q}.$ We can get the inequality,
$\mathbb{W}_{p}(\hat{P},\hat{Q})\leq\dfrac{\mathbb{W}_{p}(\hat{P},\hat{Q})^{p}}{p}+\dfrac{p-1}{p},$
[…]
$\mathbb{W}_{p}(\hat{Q}_{k},P)\leq\frac{1}{p}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\frac{p-1}{p}\text{, and }\mathbb{W}_{p}(P,\hat{Q}_{T})\leq\frac{1}{p}\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}+\frac{p-1}{p}$
Plugging this back into inequality 42,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\dfrac{1}{p}\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}\biggr)+\dfrac{2(p-1)}{p}+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}).$
Modulo some additional constants depending on $p>1$, we have the same reasoning as $p=1$.
Algorithm 3 Wasserstein Barycenter Transport
[…]
2:Labeled target domain data $\{T_{P\rightarrow Q_{T}}(x_{i}^{(P)}),y_{i}^{(P)}\}_{i=1}^{(P)}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  Compute $P=\text{Bar}(\lambda,\mathcal{Q})$
5:  Compute $\gamma^{\star}$ between barycenter and target using eq. 43
6:  Transport $(x_{i}^{(P)},y_{i}^{(P)})$ Using eq. 44
7:end for
Wasserstein Barycenter Transport $(p=2)$. [montesuma2021wasserstein] uses a two-step procedure for finding $P$. First, one solves,
$P^{\star}=\underset{P\in\mathcal{P}_{1,\text{ac}}(\Omega)}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(Q_{k},P)^{2},$
which is the 22-Wasserstein barycenter problem. Then, one aligns the barycenter with the target domain through OT. This is done through the barycenter mapping,
$\gamma^{\star}$ $=\underset{\gamma\in\Gamma(P,\hat{Q}_{T})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{T}}\gamma_{ij}\lVert x_{i}^{(P)}-x_{j}^{(Q_{T})}\rVert_{2}^{2},$ (43)
$T_{P\rightarrow Q_{T}}(x_{i}^{(P)})$ $=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$ (44)
which is the minimizer of $P\mapsto\mathbb{W}_{2}(P,\hat{Q}_{T})^{2}$. We provide an algorithm for this strategy in Algorithm 3. Note that since this procedure generates a labeled dataset on data following the target domain measure, we can train a classifier with the pairs $\{T_{P\rightarrow Q_{T}}(x_{i}^{(P)}),y_{i}^{(P)}\}$.
Appendix D Missing Proofs and Derivations
D.1 Connecting equations 12 and 16
In this section, we the empirical flow in equation 16 from the continuity equation. As a reminder, the continuity equation is,
$\partial_{t}P_{t}=\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})).$
Here we want to plug $\hat{P}_{t}=n^{-1}\sum_{i=1}^{n}\delta_{z_{t,i}}$ into this equation. On the one hand, plugging the formula for the empirical measure into the r.h.s. of this equation gives us,
$\text{div}(\hat{P}_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t}))=\text{div}\biggr(\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\nabla(\delta\mathbb{F}(\hat{P}_{t}))(z_{t,i})\biggr).$ (45)
Here, we used the definition for the Wasserstein gradient (c.f., equation 36), i.e., $\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla(\delta\mathbb{F}(P))(z_{t,i})$. On the other hand, due to the linearity of derivatives and the chain rule,
$\partial_{t}\hat{P}_{t}=\dfrac{1}{n}\sum_{i=1}^{n}\partial_{t}\delta_{z_{t,i}^{(P)}}=-\dfrac{1}{n}\sum_{i=1}^{n}\text{div}(\partial_{t}z_{t,i}\delta_{z_{t,i}})=\text{div}\biggr(-\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\partial_{t}z_{t,i}\biggr).$ (46)
Equating the r.h.s. of equations 45 and 46 gives us the continuous-time update $\partial_{t}z_{t,i}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t})$. We can discretize this differential equation in time, using the Forward Euler scheme,
$z_{\tau+1,i}=z_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}),$
where $\alpha>0$ is the time-step and $\tau$ is the discrete time index.
D.2 Proposition 3.1
As a recap, in this section we provide a formal proof that when,
$P=\sum_{i}\pi_{i}^{(P)}(\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}}),$
and under the ground cost,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime}),$
the $\mathbb{MW}_{2}$ distance becomes,
$\mathbb{MW}_{2}(P,Q)^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\omega^{\star}_{ij}(\mathbb{W}_{2}(P_{i},Q_{j})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}).$
As a reminder, we have the following proposition,
Proposition D.1.
Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be 2 GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$. Let the ground cost $c$ be,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
Proof.
Our first step is decomposing the Wasserstein distance $\mathbb{W}_{2}(P_{i},Q_{j})^{2}$ into $\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}$. Indeed,
$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int c(z,z^{\prime})d\gamma(z,z^{\prime}),$
where $z=(x,y)$ is the feature-label pair. Note that since $P_{i}=P_{i,x}\otimes\delta_{\nu_{i}^{(P)}}$, $\gamma$ is supported on the set $\{(x,y,x,y^{\prime}):y=\nu_{i}^{(P)},y^{\prime}=\nu_{j}^{(Q)}\}$. Thus, the OT plan has the structure:
$\gamma=\gamma_{x}\otimes\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}$
where $\gamma_{x}$ is the OT plan between Gaussians $P_{i,x}$ and $Q_{j,x}$. This means that the previous integral can be decomposed into,
$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int(\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2})d\gamma(z,z^{\prime}),$
which we can decompose two terms. The first,
$\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma(z,z^{\prime})$ $=\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma_{x}(x,x^{\prime})=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2},$
and,
$\int\int\rho(y,y^{\prime})^{2}d\gamma(z,z^{\prime})$ $=\int\int\rho(y,y^{\prime})^{2}\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}dydy^{\prime}=\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2},$
which gives the desired decomposition. ∎
The result in the previous proposition stems from the special structure we defined for the GMMs. Since the result is agnostic with respect the label metric, it holds for $\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})=\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}$
D.3 Theorem 3.1
In this section, we present the proof for theorem 3.1. We mainly use two tools to obtain this result,
  1. 1.
    The exponential convergence given by the PL inequality,
    $\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (47)
  2. 2.
    An uniform bound on $P\mapsto\mathbb{B}_{\mathcal{Q}}(P)$ controlling the error of approximating $Q_{k}$ with $\hat{Q}_{k}$.
These results are available, respectively, in [chewi2024statistical] and [portales2025sample]. For completeness, we re-state and demonstrate them. We start with a modified version of [chewi2024statistical, Corollary 5.17],
Proposition D.2.
(Measure theoretic Gronwall’s Lemma [chewi2024statistical]) Let $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ be a functional that satisfies the PL inequality (c.f. 47) with constant $C_{\text{PL}}>0$, then,
$\mathbb{F}(P_{\tau})-\mathbb{F}^{\star}\leq e^{-2C_{\text{PL}}\tau}(\mathbb{F}(P_{0})-\mathbb{F}^{\star}).$ (48)
Proof.
Define $\phi(t)=\mathbb{F}(P_{t})-\mathbb{F}^{\star}$. Then,
$\partial_{t}\phi(t)$ $=\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$
$\geq-C_{\text{PL}}(\mathbb{F}(P_{0})-\mathbb{F}^{\star})=-C_{\text{PL}}\phi(t).$
From Grönwall’s Lemma [chewi2024statistical, Lemma 5.16],
$\underbrace{\mathbb{F}(P_{t})-\mathbb{F}^{\star}}_{\partial_{t}\phi(t)}\leq e^{-C_{\text{PL}}}\underbrace{(\mathbb{F}(P_{0})-\mathbb{F}(P)^{\star})}_{\phi(0)}$
We now present the uniform approximation bound of [portales2025sample] for empirical barycenters,
Theorem D.1.
(Barycenter Empirical Approximation error [portales2025sample]) Let $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ be a finite family of measures on $\mathcal{P}_{p,\text{ac}}(\mathcal{B}(0,R))$ for $R>0$. Let $\{z_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}$, $i=1,\cdots,n_{k}$ be a set of $K$ i.i.d. samples from each $Q_{k}$. Let $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}$. Let $p\in[1,+\infty)$. Then,
$\mathbb{E}\biggr[\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{sup}}\,|\mathbb{B}_{\mathcal{Q}}(\hat{P})-\widehat{\mathbb{B}}_{\mathcal{Q}}(\hat{P})|\biggr]\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$ (49)
where,
$C_{p,R}$ $=8\sqrt{2}\int_{0}^{3(2R)^{p}}\sqrt{\log\biggr(2+\dfrac{32p(2R)^{p}}{s}\biggr)}ds,$
$C_{d,m}$ $=m(d+1).$
With these bounds established, we can then prove our Theorem 3.1,
Proof.
Let $\tau>0$ be the current optimization step. We have the following,
[…]
From equation 50, we can bound the expected empirical optimization gap,
$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq$ $\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]+\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|]+\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}^{\star}-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|].$ (51)
Now, we can bound the first and third terms using the approximation error of Portales et al. [portales2025sample] (i.e., equation 49),
$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$ (52)
$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}-\mathbb{B}_{\mathcal{Q}}^{\star}|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$
here, we recall that $\mathbb{B}_{\mathcal{Q}}^{\star}=\mathbb{B}_{\mathcal{Q}}(P^{\star})=\text{inf}_{P}\mathbb{B}_{\mathcal{Q}}(P)$. Coming back to the second term, note that the true optimization gap $\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}^{\star}_{\mathcal{Q}}$ is deterministic, as it does not involve sampling from the measures in $\mathcal{Q}$. This means that,
$\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|]=|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|\leq e^{-C_{\text{PL}}\tau}(\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{0})-\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}).$ (53)
Combining equations 5352 and 51, we get,
$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}_{\mathcal{Q}}(\hat{P}_{0})-\mathbb{B}^{\star}_{\mathcal{Q}})+C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$
which is the desired inequality. ∎
D.4 Theorem 3.2
The goal of this section is to stablish the proof of theorem 3.2. The proof roughly follows the same idea of the empirical case. However, since all the involved measures are GMMs by hypothesis, we do not have empirical approximation gaps per se. In this regard, we do consider the stochasticity of fitting GMMs, provided data from each underlying $Q_{k}$. As a result, we need an approximation tool, in the spirit of theorem D.1 of [portales2025sample]. We do so through the following result of [boite2025differentiable],
Proposition D.3.
(GMM approximation bound [boite2025differentiable]) Let $P$ and $Q$ be GMMs with parameters $\{(\pi^{(P)}_{i},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}_{i=1}^{n}$ (resp. $Q$), and let $\hat{P}$ and $\hat{Q}$ be the corresponding GMMs empirically estimated by the expectation-maximization algorithm. Assume the following,
  • Hypothesis 1 (bounded parameters): There are constants $R_{\mu}>0$ and $R_{\Sigma}>0$ such that,
    $\lVert\mu_{i}^{(P)}\rVert_{2}\leq R_{\mu},\sqrt{\text{Tr}(\Sigma_{i}^{(P)})}\leq R_{\Sigma},\lVert\hat{\mu}_{i}^{(P)}\rVert_{2}\leq R_{\mu},\sqrt{\text{Tr}(\hat{\Sigma}_{i}^{(P)})}\leq R_{\Sigma}\text{, resp. }Q,$
  • Hypothesis 2 (empirical convergence): Assume that exist $\rho_{\pi}>0$, $\rho_{\mu}$, and $\rho_{\Sigma}>0$ such that,
    $\mathbb{E}[\lVert\pi^{(P)}-\hat{\pi}^{(P)}\rVert_{1}]\leq\rho_{\pi}\quad\mathbb{E}[\lVert\mu_{i}^{(P)}-\hat{\mu}^{(P)}_{i}\rVert_{2}]$ $\leq\rho_{\mu}\quad\mathbb{E}[d_{\text{bures}}(\Sigma_{i}^{(P)},\hat{\Sigma}_{i}^{(P)})]$ $\leq\rho_{\Sigma}$
    $i=1,\cdots,n$ (resp. $Q$), where,
    $d_{\text{bures}}(\Sigma,\Sigma^{\prime})=\sqrt{\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{1/2}\Sigma^{\prime}\Sigma^{1/2})^{1/2})}.$
In these conditions,
$\mathbb{E}[|\mathbb{MW}_{2}(\hat{P},\hat{Q})^{2}-\mathbb{MW}_{2}(P,Q)^{2}|]\leq\underbrace{8R_{\mu}\rho_{\mu}+8R_{\Sigma}\rho_{\Sigma}+8(R_{\mu}^{2}+R_{\Sigma}^{2})\rho_{\pi}}_{C_{\pi,\mu,\Sigma}},$ (54)
Note that the same result applies for labeled GMM. Indeed, since we fit a GMM per class, the labels are deterministic. As a result, $\nu_{i}^{(P)}=\hat{\nu}_{i}^{(P)}$. These previous theorems allow us to bound the barycenter functional,
Corollary D.1.
Let $\{Q_{k}\}_{k=1}^{K}$ be $K>0$ GMMs over $\Omega$, and let $\{\hat{Q}_{k}\}_{k=1}^{K}$ be their corresponding empirical approximations. Assume Hypothesis 1 and 2 of Proposition 4.3. Let $P$ be a GMM, and $\lambda\in\Delta_{K}$, then,
$\mathbb{E}[|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|]\leq\sum_{k=1}^{K}\lambda_{k}C_{k}=C_{\lambda,\mathcal{Q}},$
where $C_{k}=8R_{\mu_{k}}+8R_{\Sigma_{k}}\rho_{\Sigma_{k}}+8(R_{\mu_{k}}^{2}+R_{\Sigma_{k}}^{2})\rho_{\pi_{k}}$.
Proof.
The proof relies on the triangle inequality:
$|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|$ $\leq\biggr|\sum_{k=1}^{K}\lambda_{k}\mathbb{MW}_{2}(P,Q_{k})^{2}-\sum_{k}\lambda_{k}\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}\biggr|$
$\leq\sum_{k=1}^{K}\lambda_{k}|\mathbb{MW}_{2}(P,Q_{k})^{2}-\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}|,$
then, taking the expectation on both sides,
$\mathbb{E}[|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|]\leq\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[|\mathbb{MW}_{2}(P,Q_{k})^{2}-\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}|]\leq\sum_{k=1}^{K}\lambda_{k}C_{k}=C_{\lambda,\mathcal{Q}}.$
Furthermore, note that the constant can be written explicitly as,
$C_{\lambda,\mathcal{Q}}=\sum_{k=1}^{K}\lambda_{k}(8R_{\mu_{k}}+8R_{\Sigma_{k}}\rho_{\Sigma_{k}}+8(R_{\mu_{k}}^{2}+R_{\Sigma_{k}}^{2})\rho_{\pi_{k}})$ (55)
With these results established, we now prove Theorem 3.2,
Proof.
Our arguments largely follow those in section D.3. For instance, the steps taken in equations 50 and 51 directly apply to the GMM setting. Now, in place of equation 52, we have,
$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]\leq C_{\lambda,\mathcal{Q}},\quad\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}-\mathbb{B}_{\mathcal{Q}}^{\star}|]\leq C_{\lambda,\mathcal{Q}}.$ (56)
Furthermore, equation 53 also holds in the GMM case. Indeed, the optimization gap is deterministic. Therefore, combining equations 56 with 51 and 53, we get,
$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}_{\mathcal{Q}}(\hat{P}_{0})-\mathbb{B}^{\star}_{\mathcal{Q}})+C_{\lambda,\mathcal{Q}},$
which is the desired inequality.
Appendix E Additional Toy Experiments
E.1 Two Moons
Figure 5: Two moons samples alongside the fitted GMMs. Stars and ellipses denote the means and covariance matrices.
In this experiment, similarly to the Swiss-roll example we explored in the main paper, we deal with Location-Scatter families of measures. The idea is that we are given $Q_{0}\in\mathcal{P}_{2,\text{ac}}$, and we define a finite family of measures,
$\mathcal{Q}=\{T_{k,\sharp}Q_{0}:T_{k}(x)=A_{k}x+b_{k},A_{k}\in\text{PD}(d)\}_{k=1}^{K}$ (57)
where $\text{PD}(d)$ denotes the set of $d\times d$ positive-definite matrices. The barycenter of $Q_{1},\cdots,Q_{K}$ is itself a push-forward of $Q_{0}$, noted $T$. Furthermore, $T$ has the form,
$T(x)=\sum_{k=1}^{K}\lambda_{k}T_{k}(x)=\biggr(\sum_{k=1}^{K}\lambda_{k}A_{k}\biggr)x+\sum_{k=1}^{K}\lambda_{k}b_{k}$
For this experiment, we use $Q_{0}$ as the two moons dataset [^2] More specifically, we use https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_moons.html to generate data from the two moons measure, and https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_swiss_roll.html to generate data from the Swiss roll measure.. For the transformations, we use a family of rotation-shift affine functions, i.e.,
$A_{k}=\begin{bmatrix}\cos\theta_{k}&-\sin\theta_{k}\\ \sin\theta_{k}&\cos\theta_{k}\end{bmatrix},\quad b_{k}\in\mathbb{R}^{2}.$
where $\theta_{k}\in\{-30,-15,15,30\}$ degrees, and,
$b_{k}\in\{\begin{bmatrix}0,0\end{bmatrix},\begin{bmatrix}10,0\end{bmatrix},\begin{bmatrix}0,10\end{bmatrix},\begin{bmatrix}10,10\end{bmatrix}\}$.
The generated family of measures is shown in Figure 5.
Figure 6: From left to right, comparison of the empirical barycenter with and without an interaction energy term, and the Gaussian mixture barycenter. Overall, due its clustering effect, the GMM barycenter better respects the underlying geometry of the input measures.
We show a comparison of the empirical and GMM barycenters in Figure 6. Note that, without an interaction term, the empirical barycenter mixes a few samples from opposing classes. This artifact comes from the rotations in the input measures. Adding the interaction energy term $U_{R}$ solves this problem, at the cost of some deformation of the class clusters. Finally, the GMM flow is able to compute an accurate barycenter due its clustering effect on the Gaussian components. We refer readers to [montesuma2024lighter] and [montesuma2025optimal] for more details on this effect.
E.2 Handling Class Imbalance
A challenge classic Wasserstein barycenter algorithms face is handling class imbalance. For instance, since the algorithm in [montesuma2023multi] relies on label propagation, the soft-labels in the Wasserstein barycenter might end up mixing classes. This phenomenon leads to Wasserstein barycenters with mixed classes and fuzzy labels. We tackle these problems with the functionals described in Section 3.3 in the main paper. We show, in Figure 7, an example of the imbalance problem. Here, $\pi=[\pi_{0},\pi_{1}]$ denotes the class proportions.
Figure 7: Imbalanced Gaussian blobs for the class imbalance example. We draw $\pi_{0}\sim\mathcal{U}[0,1]$ and let $\pi_{1}=1-\pi_{0}$.
We start by comparing the joint Wasserstein barycenter algorithm of [montesuma2023multi] (a) with our mini-batch strategy (b). Note that using mini-batches already helps with the fuzzy label problem, but many labels are still uncertain in the final barycenter. Adding the label entropy penalty $\mathbb{V}$, with underlying functional $V$ defined in equation 20 in the main paper, leads to sharp labels, as we show in (c). Finally, by adding the interaction energy term we can separate the classes, as show in Figure 8 (d).
(a) [montesuma2023multi].
(b) $\mathbb{B}_{\mathcal{Q}}$ (ours).
(c) $\mathbb{B}_{\mathcal{Q}}+\mathbb{V}$ (ours).
(d) $\mathbb{B}_{\mathcal{Q}}+\mathbb{V}+\mathbb{U}$ (ours).
Figure 8: Discrete (a), and mini-batch Wasserstein barycenters (b and c). Overall, the discrete algorithm of [montesuma2023multi] struggles with class imbalances, since the transport plan matches points from different classes. This phenomenon leads to fuzzy samples near the class boundary. This issue is partly solved through mini-batching (b), but a definitive solution comes with adding the entropy functional defined in equation 20 in the main paper.
Appendix F Additional details on domain adaptation
F.1 Evaluation Protocol, Benchmarks, Fine-Grained Results
Benchmarks. In total, we compare five domain adaptation benchmarks, divided into computer vision, neuroscience and chemical engineering. An overview of the benchmarks is presented in Table 1 in our main paper. Here, we give further details on each of the benchmarks. For computer vision, we use the Office 31 [saenko2010adapting] and Office Home [venkateswara2017deep] benchmarks. For neuroscience, we use the BCI-CIV-2a [brunner2008bci] and ISRUC [khalighi2013automatic] benchmarks. For chemical engineering, we use the TEP benchmark [montesuma2024benchmarking].
Evaluation Protocol. As we describe in the following, all benchmarks deal with multi-class classification. We measure performance using the standard classification accuracy,
$\text{accuracy}(y,\hat{y})=\dfrac{1}{n}\sum_{i=1}^{n}\delta(y_{i}-\hat{y}_{i})\times 100\%.$
For the Office 31, Office Home, Tennessee Eastman Process, and BCI-CIV-2a, we adopt the so-called leave-one-domain-out evaluation. For instance, given $3$ domains $\{A,B,C\}$, we perform $3$ experiments, leaving one domain as the target domain at each time. For instance, one of these experiments would be $\{A,B\}\rightarrow C$, where $\{A,B\}$ are the sources, and $C$ is the target.
Backbones. As we discuss in the main paper, we perform adaptation over embeddings. To that end, we use the same experimental setting as [montesuma2024lighter] for the Office 31, Office Home, and TEP benchmarks. More specifically, we train a ResNet 50 and ResNet 101 for the Office 31 and Office Home benchmarks, and a CNN for the TEP benchmark. For this latter backbone, we refer readers to [montesuma2024benchmarking] for further details. For neuroscience benchmarks, we follow the experimental protocol of [wang2024cbramod], where we fine-tune CBraMod, a transformer-based foundation model, on available source domain data. We use the exact same training settings as these authors.
Algorithm $\mathcal{X}\times\mathcal{Y}$ A D W Avg. $\uparrow$
ResNet50 - 67.50 95.00 96.83 86.40
WJDOT [turrisi2022multi] - 67.77 97.32 95.32 86.80
DaDiL-E [montesuma2023multi] - 70.55 100.00 98.83 89.79
DaDiL-R [montesuma2023multi] - 70.90 100.00 98.83 89.91
GMM-DaDiL [montesuma2024lighter] - 72.47 100.00 99.41 90.63
Discrete [cuturi2014fast] 64.45 88.39 92.98 81.94
NormFlow [visentin2025computing] 65.67 97.32 94.73 85.91
CW2B [korotin2021continuous] 69.16 94.64 95.32 86.37
NOT [kolesov2024estimating] 69.33 92.85 96.49 86.22
U-NOT [gazdieva2024robust] 68.29 97.32 95.32 86.97
Discrete [montesuma2023multi] 67.94 98.21 97.66 87.93
GMM [montesuma2024lighter] 70.13 99.11 96.49 88.54
WGF (ours) 71.25 98.21 97.07 88.84
WGF-GMM (ours) 71.77 99.11 97.07 89.34
Table 4: Office 31.
Office 31 [saenko2010adapting] contains, in total, 4652 images among 31 classes of objects. These images are divided into 3 different domains: Amazon, dSLR, and Webcam. Images in the Amazon domain represent products, usually with an homogeneous background, medium resolution, and studio lightning. There are 2817 images in this domain. The dSLR domain contains high-resolution images obtained through a digital SLR camera in an office environment. There are 498 images in this domain. Finally, the Webcam domain contains images of objects captured with a webcam, also in an office environment. There 795 images in this domain. As analyzed in [ringwald2021adaptiope], this benchmark is one of the most commonly used in the literature. Despite its moderate size, it presents challenges of its own. For instance, since the Amazon domain images were collected automatically, it has label noise.
We show our results in the Office 31 benchmark in Table 4. Overall, dictionary learning methods such as DaDiL [montesuma2023multi] or GMM-DaDiL [montesuma2024lighter] remain state-of-the-art for this benchmark. However, our gradient flow algorithm is able to improve previous methods, closing the gap between these two kinds of algorithms. In comparison with unsupervised barycenters, using labels in the ground-cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) has a clear advantage.
Algorithm $\mathcal{X}\times\mathcal{Y}$ Ar Cl Pr Rw Avg. $\uparrow$
ResNet101 - 72.90 62.20 83.70 85.00 75.95
WJDOT [turrisi2022multi] - 74.28 63.80 83.78 84.52 76.59
DaDiL-E [montesuma2023multi] - 77.16 64.95 85.47 84.97 78.14
DaDiL-R [montesuma2023multi] - 75.92 64.83 85.36 85.32 77.86
GMM-DaDiL [montesuma2024lighter] - 77.16 66.21 86.15 85.32 78.81
Discrete [cuturi2014fast] 68.11 59.79 75.56 78.21 70.42
NormFlow [visentin2025computing] 77.16 62.54 83.10 84.63 76.86
CW2B [korotin2021continuous] 74.69 63.46 82.54 85.09 76.44
NOT [kolesov2024estimating] 72.22 63.57 82.77 82.91 75.36
U-NOT [gazdieva2024robust] 75.30 64.94 82.88 84.28 76.85
Discrete [montesuma2023multi] 75.72 63.80 84.23 84.63 77.09
GMM [montesuma2024lighter] 75.31 64.26 86.71 85.21 77.87
WGF (ours) 74.48 65.17 85.92 85.77 77.84
WGF-GMM (ours) 74.89 65.97 87.16 86.81 78.71
Table 5: Office-Home.
Office-Home [venkateswara2017deep] contains, 15,500 images among 65 classes of objects. There are four domains: Art, Clipart, Product, and Real-World. The Art domain contains artistic depictions of objects, such as sketches, paintings, and ornamentation. It has 2427 images. The Clipart domain is a collection of graphic art representing the objects. This domain has 4365 images. Product contains images of objects without a background. It has 4439 images. Real-World contains images of objects captured with a camera. It contains 4357 images. In this sense, the Product domain is similar to the Amazon domain in the Office 31 benchmark, and the Real-World is similar to dSLR or Webcam. In comparison with the Amazon benchmark, the Office-Home benchmark adds another factor of domain variation: style (e.g., Art vs. Real-World). Furthermore, due the larger number of classes, one might expect a harder classification problem.
We show our results in Table 5. In comparison with the Office 31 benchmark, our methods get even closer to the state-of-the-art. For instance, WGF-GMM has only a $0.10\%$ gap in performance with GMM-DaDiL, while showing state-of-the-art performance on domains Product and Real-World. Again, we verify that using labels in the ground cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) offers a clear advantage.
Remark F.1.
We follow previous works on visual domain adaptation, especially [saenko2010adapting] and [venkateswara2017deep], and report the classification accuracy on fixed partitions of the target domains. For that reason, Tables 4 and 5 do not have confidence intervals. While it would be possible to run these experiments with random partitions (e.g., $k$-fold cross validation), this would mean that our results would not be comparable to previous established results in [montesuma2024lighter], for instance.
Tennessee-Eastman Process (TEP) [montesuma2024benchmarking] is a benchmark in chemical engineering. This benchmark comprises a set of simulations of a chemical reaction that occurs inside a reactor. In total, 34 variables are measured over time. As a result, we have a set of time-series representing how the reaction progresses over each simulation. In some of these simulations, one of 28 possible faults can be introduced. The goal is, based on the readings, to determine which kind of fault, or its absence, has occurred. As a result, there are 29 classes. Each simulation is also associated with a production mode, which drastically changes the statistical properties of the data generation process, thus introducing a distribution shift. As a result, each domain corresponds to one of these modes of operation. We refer readers to [montesuma2024benchmarking] and [reinartz2021extended] for more details.
Algorithm $\mathcal{X}\times\mathcal{Y}$ Mode 1 Mode 2 Mode 3 Mode 4 Mode 5 Mode 6 Avg. $\uparrow$
CNN 80.82 $\pm$ 0.96 63.69 $\pm$ 1.71 87.47 $\pm$ 0.99 79.96 $\pm$ 1.07 74.44 $\pm$ 1.52 84.53 $\pm$ 1.12 78.48
WJDOT [turrisi2022multi] - 89.06 $\pm$ 1.34 75.60 $\pm$ 1.84 89.99 $\pm$ 0.86 89.38 $\pm$ 0.77 85.32 $\pm$ 1.29 87.43 $\pm$ 1.23 86.13
DaDiL-E [montesuma2023multi] - 90.45 $\pm$ 1.02 77.08 $\pm$ 1.21 86.79 $\pm$ 2.14 89.01 $\pm$ 1.35 84.04 $\pm$ 3.16 87.85 $\pm$ 1.06 85.87
DaDiL-R [montesuma2023multi] - 91.97 $\pm$ 1.22 77.15 $\pm$ 1.32 85.41 $\pm$ 1.69 89.39 $\pm$ 1.03 84.49 $\pm$ 1.95 88.44 $\pm$ 1.29 86.14
GMM-DaDiL [montesuma2024lighter] - 91.72 $\pm$ 1.41 76.41 $\pm$ 1.89 89.68 $\pm$ 1.49 89.18 $\pm$ 1.17 86.05 $\pm$ 1.46 88.02 $\pm$ 1.12 86.85
Discrete [cuturi2014fast] 90.38 $\pm$ 1.29 71.17 $\pm$ 1.26 85.27 $\pm$ 1.09 87.29 $\pm$ 0.99 83.83 $\pm$ 1.74 84.94 $\pm$ 1.87 83.81
NormFlow [visentin2025computing] 87.06 $\pm$ 0.57 66.25 $\pm$ 1.11 88.58 $\pm$ 1.00 89.35 $\pm$ 0.34 80.99 $\pm$ 2.21 85.08 $\pm$ 1.45 82.89
CW2B [korotin2021continuous] 91.72 $\pm$ 0.79 73.71 $\pm$ 1.42 88.37 $\pm$ 1.31 89.32 $\pm$ 1.14 84.66 $\pm$ 1.75 87.19 $\pm$ 1.24 85.83
NOT [kolesov2024estimating] 90.31 $\pm$ 1.08 72.37 $\pm$ 0.94 87.99 $\pm$ 1.31 89.28 $\pm$ 0.81 84.35 $\pm$ 1.68 85.29 $\pm$ 2.02 84.93
U-NOT [gazdieva2024robust] 90.82 $\pm$ 0.65 72.68 $\pm$ 1.58 88.65 $\pm$ 0.95 89.00 $\pm$ 1.28 85.22 $\pm$ 1.07 86.19 $\pm$ 1.70 85.43
Discrete [montesuma2023multi] 92.38 $\pm$ 0.66 73.74 $\pm$ 1.07 88.89 $\pm$ 0.85 89.38 $\pm$ 1.26 85.53 $\pm$ 1.35 86.60 $\pm$ 1.63 86.09
GMM [montesuma2024lighter] 92.23 $\pm$ 0.70 71.81 $\pm$ 1.78 84.72 $\pm$ 1.92 89.28 $\pm$ 1.55 87.51 $\pm$ 1.73 82.49 $\pm$ 1.81 84.67
WGF (ours) 92.41 $\pm$ 0.90 74.16 $\pm$ 1.29 88.78 $\pm$ 0.91 89.84 $\pm$ 0.62 84.66 $\pm$ 2.37 87.39 $\pm$ 2.11 86.21
WGF-GMM (ours) 92.55 $\pm$ 1.12 75.95 $\pm$ 1.51 89.54 $\pm$ 1.11 89.21 $\pm$ 1.10 85.63 $\pm$ 1.17 87.74 $\pm$ 1.48 86.77
Table 6: TEP.
We report our results in Table 6. For this benchmark, the gap between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) barycenter methods is closer. For instance, CW2B [korotin2021continuous] achieves 85.83 for the average target domain performance versus 86.09 for the discrete barycenter. Furthermore, it surpasses the GMM method of [montesuma2024lighter]. In comparison, our methods establish a new state-of-the-art for barycenter-based domain adaptation in this benchmark. Note that our WGF-GMM is able to get the second best performance overall, close to GMM-DaDiL.
Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 Avg. $\uparrow$
CBraMod - 52.22 40.45 64.23 40.41 47.74 48.34 53.64 50.52 55.03 50.30
DaDiL-R [montesuma2023multi] - 58.68 40.27 66.14 41.45 59.54 51.38 53.12 57.46 55.21 53.69
GMM-DaDiL [montesuma2024lighter] - 60.59 42.36 70.31 48.12 61.11 50.34 62.15 58.85 60.00 57.10
Discrete [cuturi2014fast] 60.93 43.40 69.96 50.00 61.28 50.69 61.63 61.45 57.11 57.38
NormFlow [visentin2025computing] 57.11 43.22 67.53 47.71 60.24 51.04 56.25 60.76 55.20 55.45
CW2B [korotin2021continuous] 60.00 43.57 70.13 48.95 60.93 51.21 61.28 61.63 57.46 57.25
NOT [kolesov2024estimating] 61.28 43.22 70.31 47.91 60.59 51.21 61.46 61.80 57.29 57.23
U-NOT [gazdieva2024robust] 60.59 44.09 69.44 47.70 61.11 50.34 61.97 61.11 58.15 57.17
Discrete [montesuma2023multi] 60.76 42.53 70.66 48.54 61.28 50.69 61.63 61.28 59.54 57.43
GMM [montesuma2024lighter] 60.59 42.53 69.79 47.08 61.45 50.34 62.32 59.54 59.72 57.04
WGF (ours) 61.63 43.40 71.00 47.91 61.45 51.38 61.97 61.63 57.12 57.50
WGF-GMM (ours) 60.59 43.05 70.48 47.29 61.80 51.56 62.15 61.80 57.46 57.35
Table 7: BCI-CIV-2a.
BCI-CIV-2a [brunner2008bci] Contains electroencephalogram (EEG) data from nine subjects. In brief, an EEG measures the electrical activity of the brain over time. As such, like the TEP benchmark, this benchmark has a collection of time series data. The BCI-CIV-2a data is concerned with motor imagery tasks, that is, the imagination of movement of the left hand, right hand, both feet, and tongue. These correspond to class 1, 2, 3, and 4. Therefore, in this benchmark we have four classes. Our goal here is to perform cross-subject adaptation, namely, we use data from a given set of source subjects, and try to predict on a target subject. Therefore, each domain is a subject, so there are nine domains in total. There are, on average, $560$ samples per subject, with a total of $5088$ samples.
We show our results in Table 7. For cross-subject studies, we highlight that, besides improving individual domain performance, it is desirable to not degrade the source-only performance. This is indeed the case for almost all methods. For instance, DaDiL [montesuma2023multi] degrades performance of subjects 2 and 7 by a small margin. For this benchmark, the difference between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) is not as marked as before. We hypothesize that this benchmark falls into the covariate shift hypothesis, namely, $Q_{S}(X)\neq Q_{T}(X)$, rather than the more general joint shift $Q_{S}(X,Y)\neq Q_{T}(X,Y)$.
Algorithm 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
CBraMod 78.97 80.35 70.46 85.59 80.69 72.85 77.00 73.29 83.41 68.95 76.63
WJDOT [turrisi2022multi] 80.41 78.25 65.69 82.85 81.74 77.74 76.33 68.75 86.95 70.81 76.95
DaDiL-E [montesuma2023multi] 76.12 77.90 65.93 83.21 81.28 68.45 75.66 72.61 79.75 65.93 75.14
DaDiL-R [montesuma2023multi] 78.06 78.13 66.97 84.64 81.51 72.02 76.33 72.84 80.85 67.55 75.89
GMM-DaDiL [montesuma2024lighter] 76.63 78.95 66.51 82.73 77.44 73.81 77.00 71.70 83.05 66.86 75.47
Discrete [cuturi2014fast] 78.77 80.58 69.76 85.83 79.06 73.45 78.33 72.50 84.14 68.48 77.09
NormFlow [visentin2025computing] 81.12 82.90 75.58 84.28 80.34 78.69 75.88 70.34 80.48 70.81 78.05
CW2B [korotin2021continuous] 79.08 78.25 69.07 86.54 78.48 70.35 76.88 70.90 82.80 66.04 75.84
NOT [kolesov2024estimating] 78.57 79.88 67.32 85.59 78.72 73.21 77.77 72.50 83.41 67.21 76.44
U-NOT [gazdieva2024robust] 80.20 83.48 73.72 87.50 76.51 72.62 76.66 70.22 78.17 71.51 77.06
Discrete [montesuma2023multi] 78.87 80.69 72.09 86.07 78.60 78.45 78.00 72.95 85.48 70.81 78.20
GMM [montesuma2024lighter] 75.20 75.46 69.88 80.71 78.60 71.90 76.33 68.41 78.17 68.14 74.28
WGF (ours) 80.20 81.62 76.86 86.42 79.30 80.00 75.88 76.13 86.95 74.42 79.78
WGF-GMM (ours) 79.89 81.74 72.32 86.42 80.69 79.04 77.55 72.72 85.73 71.04 78.72
Table 8: ISRUC.
ISRUC [khalighi2013automatic] is a benchmark for EEG data for sleep staging. In brief, sleep staging is the classification of physiological data, such as EEG, recorded during sleep into different sleep stages. In this context, the ISRUC benchmark is the most large scale benchmark used in this paper. This study comprises 100 subjects, with a total of 89240 samples (approximately 900 samples per domain). The sleep staging problem in this benchmark has five classes: awake, non-rapid eye movement sleep (subdivided into N1, N2, N3), and rapid eye movement sleep.
Due the elevated number of subjects in this benchmark, we adopt a slightly different evaluation protocol. Instead of using the leave-of-domain-out strategy, we fix the first $90$ subjects as training subjects and evaluate the performance of adapting to the remaining $10$ subjects. This means that we compute Wasserstein barycenters of $K=90$ input measures.
In this benchmark, using our gradient flow strategy yields the best result overall (e.g., $79.78$ average domain performance for our empirical gradient flow). Furthermore, we see an important gap between the GMM barycenter algorithm of [montesuma2024lighter] and ours, which is similar to the gap found in the TEP benchmark (c.f. Table 6). There are mainly two reasons for this. First, we can freely choose the learning rate for updating the GMM components, which leads to an overall more stable algorithm. Second, we can regularize the barycenter problem, with additional functionals.
F.2 Adapting Neural Network Barycenter Methods to Domain Adaptation
In this section, we cover how we adapted previous neural network-based barycenters to perform domain adaptation. There is one main challenge with using unlabeled barycenters for domain adaptation: we cannot directly move the barycenter to the target domain, as we would end up with more unlabeled data. A straightforward solution to this problem is moving the available labeled data that we have, that is, the source domain data. For this scenario, the barycenter domain works as a pivot domain, as we transport the source to the target through the composition,
$T_{Q_{k}\rightarrow Q_{T}}(x_{i}^{(Q_{k})})=T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})).$
In comparison, by computing a labeled barycenter, we are synthesizing data in $P^{\star}$. As we demonstrated in Tables 4 through 8, this is generally beneficial for domain adaptation. In this way, we can acquire a labeled dataset in the target domain through $\{\{T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})),y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$.
To isolate the effect of the quality of the computed barycenters, we estimate the mappings as we did in equations 43 and $4$, that is,
$T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})=n\sum_{j=1}^{n}\gamma_{k,i,j}^{\star}x_{j}^{(P)}\quad T_{P\rightarrow Q_{k}}(x_{i}^{(P)})=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$
where $\gamma_{k}^{\star}$ is the OT plan from $Q_{k}$ to $P$, and $\gamma^{\star}$ is the OT plan from $P$ to $Q_{T}$.
F.3 Complexity Analysis and Running Time
Complexity Analysis. We start with analyzing the complexity of evaluating the functionals used in our work. We recall that,
$\mathbb{V}(P)=\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)}),\quad\mathbb{U}(P)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)}),$
where, assuming that the complexity of evaluating $V$ and $U$ is $\mathcal{O}(1)$, the complexity of evaluating $\mathbb{V}$ is $\mathcal{O}(n)$ and for $\mathbb{U}$ is $\mathcal{O}(n^{2})$. For $\mathbb{B}_{\mathcal{Q}}$, the complexity depends on the number of elements used in OT, and the number of input measures. In other words, we have $\mathcal{O}(Kn_{b}^{3}\log n_{b})$ for the empirical case, and $\mathcal{O}(Kn_{\text{components}}\log n_{\text{components}})$ for WGF-GMM. We assume this cost dominates over the evaluation of other functionals.
Method Complexity per iteration Parameter Complexity Running time per iteration (s) # Iterations Running Time (s)
Discrete [cuturi2014fast, montesuma2023multi] $Kn^{3}\log n$ $n$ 4.28 50 214.37
NormFlow [visentin2025computing] N/A $|f|$ 0.17 5000 878.32
CW2B [korotin2021continuous] N/A $K(|f|+|g|)$ 2.18 1000 2181.00
NOT [kolesov2024estimating] N/A $K(|f|+|g|)$ 41.63 25 1040.75
U-NOT [gazdieva2024robust] N/A $K(|f|+|g|)$ 13.29 100 1329.77
WGF (ours) $Kn_{b}^{3}\log n_{b}$ $n$ 0.58 200 117.24
WGF-GMM (ours) $Kn_{\text{components}}^{3}\log n_{\text{components}}$ $n_{\text{components}}$ 0.11 150 17.14
Table 9: Complexity and running time (in seconds) of different barycenter calculation strategies on the ISRUC benchmark. $n_{b}$ denotes the batch size, $n$ number of samples, $K$ number of input measures, $d$ number of dimensions, $n_{\text{components}}$ number of components in GMMs. For NormFlow, $f$ denotes the neural net implementing the normalizing flow. For CW2b, NOT and U-NOT, $f$ and $g$ denote the neural networks approximating the Kantorovich potentials. $|f|$ and $|g|$ denote the number of their parameters.
Running Time. We perform a running time analysis of the tested methods on the ISRUC benchmark, which contains the most number of samples and domains. For all neural net methods, we fix a batch size of $256$ samples. Overall, our proposed methods run significantly faster than previous methods. For instance, in comparison with the discrete solver of [cuturi2014fast], our algorithm is approximately seven times faster per iteration. It is noteworthy that some neural net methods are relatively faster per iteration (e.g. NormFlow). However, we empirically verified that these methods sometimes need a large number of iterations to converge to a good barycenter (e.g. 5000 iterations).
F.4 Hyperparameter Setting
In this section, we detail the hyperparameters of our methods. For both flows, we have mainly three parameters,
  • Number of samples in the barycenter support, $n$. This hyperparameter is replaced by the number of components in the GMM barycenter, $K$.
  • Learning rate $\alpha$ for the gradient flow,
  • Number of iterations $n_{\text{iter}}$,
additionally, for the empirical flow, we have the batch size $n_{b}$ for the input measures. We show the complete list of hyperparameters in Table 10. Additionally to these parameters, we also have the choice of regularizing functionals, $\mathbb{V}$ and $\mathbb{U}$.
Remark F.2.
In all of our domain adaptation experiments, we use a diagonal GMM. We empirically verified that optimizing with respect to complete covariances is numerically unstable and time consuming.
WGF WGF-GMM
Benchmark $n$ $n_{b}$ $\alpha$ $n_{\text{iter}}$ $\mathbb{V}$ $\mathbb{U}$ $K$ $\alpha$ $n_{\text{iter}}$ $\mathbb{V}$ $\mathbb{U}$
Office31 1550 1550 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion 124 1.0 150 - -
Office-Home 1300 650 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion 975 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion
TEP 870 290 1.0 150 $\mathbb{W}_{2}(P,Q_{T})$ - 145 0.1 200 - -
BCI-CIV-2a 800 200 1.0 - - Repulsion 400 0.01 200 Entropy Repulsion
ISRUC 1000 1000 0.1 200 - - 40 0.1 150 Entropy Repulsion
Table 10: Hyperparameters used in each benchmark. $n$ denotes number of samples, $\alpha$ denotes learning rate, $n_{\text{iter}}$ denotes number of iterations.
We offer some additional reasoning on how to choose these parameters. The number of samples and components controls the complexity of the calculated barycenter. We express these values in terms of the number of classes. For instance, $n=1550$ in the Office 31 benchmark translates to $n=31\times 50$ samples, that is, $50$ samples per class. The same reasoning applies for the batch size. Furthermore, our method is fairly robust to the choice of learning rate. For instance, as we discussed in section C.1, the fixed-point iterations in the discrete barycenter are equivalent to setting $\alpha=n/2$, where $n$ is the number of samples in the Wasserstein barycenter. This generally strikes a balance on how fast the algorithm converges.