### v3

# Wasserstein Gradient Flows for Scalable and Regularized Barycenter Computation

Abstract
Wasserstein barycenters provide a principled approach for aggregating probability measures, while preserving the geometry of their ambient space. Existing discrete methods are not scalable as they assume access to the complete set of samples from the input measures. Meanwhile, neural network approaches do scale well, but rely on complex optimization problems and cannot easily incorporate label information. We address these limitations through gradient flows in the space of probability measures. Through time discretization, we achieve a scalable algorithm that i) relies on mini-batch optimal transport, ii) accepts modular regularization through task-aware functions, and iii) seamlessly integrates supervised information into the ground-cost. We empirically validate our approach on domain adaptation benchmarks that span computer vision, neuroscience, and chemical engineering. Our method establishes a new state-of-the-art barycenter solver, with labeled barycenters consistently outperforming unlabeled ones.
1 Introduction
Defining the mean or center of a set of probability measures is a fundamental primitive in geometric probability theory [nielsen2020]. Given $K\in\mathbb{N}$ probability measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ on a metric space $(\Omega,d)$, their Wasserstein barycenter [agueh2011barycenters] is given by,
$P^{\star}=\underset{P\in\mathcal{P}_{2}(\Omega)}{\text{arg min}}\,\biggl\{\mathbb{B}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr\},$ (1)
where $\mathcal{P}_{2}(\Omega)$ is the set of measures over $\Omega$ with finite second moments, and $\lambda=(\lambda_{1},\cdots\lambda_{k})$, $\lambda_{k}\geq 0$, $\forall k$, $\sum_{k}\lambda_{k}=1$ is an array of barycentric coordinates. Equation 1 defines a notion of average of the input measures in $\mathcal{Q}$. The main advantage of these barycenters stems from Optimal Transport (OT) [villani2008optimal], that is, they lift the geometry of the space $(\Omega,d)$ to the space $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ [kloeckner2010geometric].
Wasserstein barycenters have wide applications to machine learning, such as posterior aggregation in Bayesian inference [srivastava2018scalable], model fusion [singh2020model], ensembling in time series forecasting [coz2023barycenter], data augmentation [zhu2023interpolation], and Domain Adaptation (DA) [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter]. In each application, the main advantage is geometric fidelity, that is, the Wasserstein barycenter captures and preserves the underlying structure of samples. Figure 1 shows a conceptual illustration.
(a) Conceptual Illustration.
(b) $\mathbb{B}$.
(c) $\mathbb{B}+\mathbb{V}$.
(d) $\mathbb{B}+\mathbb{V}+\mathbb{U}$.
Figure 1: In (a), we show an illustration of our flow. We show the initial barycenter measure $P_{0}$ which flow towards the solution $P^{\star}$. In red, green, and orange, we show the input measures’ mini-batches as solid dots. The flow is given by the combination of gradients $v_{\tau,i}=\sum_{k=1}^{K}\lambda_{k}\nabla\mathbb{W}_{2}(P_{\tau},Q_{k})$. In parallel, (b) shows a pathological barycenter calculation, with fuzzy labels and unclear class boundaries. Adding regularizers $\mathbb{V}$ (b) and $\mathbb{U}$ (c) progressively resolves this issue.
Despite its advantages, Wasserstein barycenters remain challenging beyond the Gaussian setting [altschuler2022wasserstein]. We have identified three gaps in the current literature. First, the seminal algorithm of cuturi2014fast requires having access to the complete set of input measures’ samples at once, making it intractable for large datasets. Neural network methods [fan2020scalable, korotin2021continuous, kolesov2024estimating, gazdieva2024robust] alleviate this issue by operating on mini-batches, but parametrize the barycentric measure with $\mathcal{O}(K)$ neural networks, coupling model complexity with the number of input measures. Second, neural methods cannot seamlessly incorporate label information into their ground-cost, limiting their performance in supervised tasks (cf. Table 2). Third, the objective function in equation 1 only accounts for distributional fit. In practice, the barycentric measure must satisfy additional structural properties, such as class separation. Current barycenter solvers do not offer a principled way of enforcing these properties.
Recent methods study the barycenter problem through the gradient flow perspective. First, chewi2020gradient established a gradient flow algorithm on the Bures-Wasserstein manifold, i.e., on Gaussian measures. Second, chizat2025doubly describes the doubly regularized barycenter problem through the same lens. Here, in addition to entropic regularization at the level of OT, they proposed an outer regularization that penalizes the entropy of the barycentric measure, which leads to a Noisy Particle Gradient Descent (NPGD) algorithm for empirical measures. Although this algorithm incorporates outer regularization, it has the same limitation as cuturi2014fast, as it assumes full-batch access to the samples of input measures.
We address the aforementioned challenges through the lens of gradient flows in the space of probability measures [ambrosio2008gradient, santambrogio2017euclidean]. We therefore conceptualize the barycenter problem as the flow from an initial measure $P_{0}=\mathcal{N}(0,\text{Id})$ following the Wasserstein gradient of the functional,
$\mathbb{F}(P)=\mathbb{B}(P)+\mathbb{R}(P),$ (2)
where $\mathbb{B}$ denotes the barycenter functional (cf. equation 1), and $\mathbb{R}$ denotes additional regularizing functionals (see Sections 2.3 and 3.3, and Appendix C.2).
Following the established gradient flow literature [santambrogio2017euclidean], we decompose $\mathbb{R}$ into internal, potential, and interaction energies, denoted $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$, respectively. This decomposition generalizes the double regularization strategy of chizat2025doubly. Indeed, setting $\mathbb{V}=\mathbb{U}=0$ and $\mathbb{G}(P)=\eta_{G}H(P)$, where $H(P)$ denotes the entropy of $P$, corresponds to their exact framework. Similarly, alvarez2021dataset devised an approach for flowing a single dataset towards a known target. Our setting differs structurally from that. First, we go beyond the single measure objective, optimizing the multi-measure barycentric objective (cf. equation 1). Second, while alvarez2021dataset transforms an existing dataset towards a known target, ours synthesizes a measure (i.e., the barycentric measure) from noise.
Our contributions are as follows. (i) A mini-batch, time discretized gradient flow algorithm for regularized empirical Wasserstein barycenters (Algorithm 1). By randomly sampling the input measures, we reduce the computational complexity of traditional barycenter methods [cuturi2014fast, montesuma2023multi, chizat2025doubly] achieving $2\times$ to $50\times$ speedups with respect to discrete solvers (Figure 5). (ii) Modular, task-aware regularizing functionals for Wasserstein barycenters (Section 3.3). Following previous gradient flow literature [santambrogio2017euclidean, alvarez2021dataset], we decompose $\mathbb{R}$ into internal, potential, and interaction energies, allowing incorporation of plug-and-play regularizers into the barycenter problem. This goes beyond chizat2025doubly, who considered a particular internal energy as an outer regularization strategy. (iii) Extensive DA benchmarking (Section 4). We empirically show that our algorithm produces barycenters that respect class structure, by incorporating both appropriate functionals (Section 3.3), and labels into the ground-cost of optimal transport (Section 3.2). On 5 DA benchmarks, spanning computer vision, neuroscience, and chemical engineering, we establish a comparison between empirical and neural barycenter solvers, showing that structured costs are essential for DA performance.
Paper organization. Section 2 reviews OT, Wasserstein barycenters and gradient flows. Section 3 develops our empirical, time-discretized gradient flow framework and algorithms. Section 4 presents our experiments. Finally, Section 5 concludes this paper.
2 Background
Throughout this paper $(\Omega,d)$ is a metric space, and its elements are denoted $z\in\Omega$. We denote $\mathcal{P}_{2}(\Omega)$ as the set of measures with finite second moments [villani2008optimal, Definition 6.4]. We work with empirical measures,
$\text{Emp}_{n}(\Omega)=\biggr\{\hat{P}\in\mathcal{P}_{2}(\Omega):\hat{P}=\frac{1}{n}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}\biggr\},$ (3)
where $\delta_{z_{0}}(z)=\delta(z-z_{0})$ denotes the Dirac measure centered at $z_{0}\in\Omega$. Given a measure $P\in\mathcal{P}_{2}(\Omega)$, we denote by $\hat{P}\in\text{Emp}_{n}(\Omega)$ its empirical approximation based on i.i.d. samples $z_{1}^{(P)},\cdots,z_{n}^{(P)}\sim P$. The superscript $\cdot^{(P)}$ indicates the measure from which the samples originate.
2.1 Optimal Transport
OT is a field of mathematics [villani2008optimal] concerned with mass transportation at least effort [peyre2019computational, montesuma2024recent]. It was originally founded by monge1781memoire, and seeks a mapping such that,
$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\mathbb{E}_{z\sim P}[d(z,T(z))^{2}].$ (4)
$T_{P\rightarrow Q}^{\star}$ is an OT mapping between $P$ and $Q$, and $T_{\sharp}$ is the pushforward mapping $P$ to $Q$. Alternatively, kantorovich1942transfer proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\mathbb{E}_{(z,z^{\prime})\sim\gamma}[d(z,z^{\prime})^{2}],$ (5)
where $\gamma^{\star}$ denotes an OT plan. OT defines the so-called 22-Wasserstein distance, through the infimum value of the optimization problem in equations 4 and 5,
$\mathbb{W}_{2}(P,Q)^{2}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\mathbb{E}_{(z,z^{\prime})\sim\gamma}[d(z,z^{\prime})^{2}].$ (6)
Equation 6 defines a metric in $\mathcal{P}_{2}(\Omega)$ which bridges the notions of Euclidean (i.e., center of mass) and Wasserstein barycenters. An alternative consists of regularizing the OT problem with the entropy of $\gamma$, resulting in the Sinkhorn divergence cuturi2013sinkhorn,
$\mathbb{W}_{2,\epsilon}(P,Q)^{2}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\mathbb{E}[d(z,z^{\prime})^{2}]+\lambda H(\gamma),$ (7)
where $H(\gamma)=\text{KL}(\gamma|P\otimes Q)$, for $(P\otimes Q)(z)=P(z)Q(z)$, and KL is the Kullback Leibler divergence. Equation 7 enjoys many desirable properties. First, by relying on the matrix scaling algorithm of sinkhorn1964relationship, it has complexity $\mathcal{O}(n^{2})$ per iteration. Second, it can be vectorized in modern GPUs, leading to dramatic speed ups. Third, since $H(\gamma)$ is strictly convex, the overall optimization problem in equation 7 is convex, leading to a unique solution.
2.2 Wasserstein Barycenter
In the metric setting, the barycenter problem is known as frechet1948elements or karcher1977riemannian means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}=\{a\in\mathbb{R}^{K}_{\geq 0}:\sum_{k}a_{k}=1\}$, we define the Wasserstein barycenter over $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ through the following optimization problem,
$\hat{P}^{\star}=\underset{\hat{P}\in\text{Emp}_{n}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{2,\epsilon}(\hat{P}|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2,\epsilon}(\hat{P},\hat{Q}_{k})^{2}\biggr\}.$ (8)
[…]
2.3 Gradient Flows
Our method relies on the gradient flow interpretation of the minimization in equation 8. Starting from the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional. A gradient flow is the solution to the Ordinary Differential Equation (ODE),
$\dot{x}(t)=-\nabla F(x(t))\text{, subject to }x(0)=x_{0}.$ (9)
This framework extends to probability measures. Let $\mathbb{F}:\mathcal{P}_{2}(\Omega)\rightarrow\mathbb{R}$ be a functional over $\mathcal{P}_{2}(\Omega)$. The evolution of a curve $\{P_{t}\}_{t\geq 0}$ in $\mathcal{P}_{2}(\Omega)$ is given by the continuity equation
$\partial_{t}P_{t}=-\text{div}(P_{t}v_{t}),$ (10)
where $v_{t}:\Omega\rightarrow\Omega$ is a velocity field, and div is the divergence operator, understood in a measure theoretic sense. Choosing $v_{t}=-\nabla_{W}\mathbb{F}(P_{t})$ yields the gradient flow of $\mathbb{F}$, that is, the measure theoretic analogue of equation 9. Here, $\nabla_{W}$ denotes the Wasserstein gradient of the functional $\mathbb{F}$ (c.f. [chewi2024statistical, Section 5.4] and appendix C.1).
The structure of $\mathbb{F}$ determines the dynamics of $P_{t}$. Following the classical OT theory [santambrogio2015optimal, Chapter 7], we consider the functional $\mathbb{F}(P)=\mathbb{B}(P)+\mathbb{R}(P)$, where,
$\mathbb{R}(P)=$ $\underbrace{\int G(p(z))dz}_{\text{Internal energy }\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\text{Potential energy }\mathbb{V}(P)}+$ (11)
$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\text{Interaction energy }\mathbb{U}(P)}$
where $p:\Omega\rightarrow\mathbb{R}$ denotes the density of $P$, $G:\mathbb{R}\rightarrow\mathbb{R}$, $V:\Omega\rightarrow\mathbb{R}$, and $U:\Omega\times\Omega\rightarrow\mathbb{R}$. For example, when $(s)=s\log s, 𝔾\mathbb{G}$ is the entropy functional of [chizat2025doubly], associated with diffusion dynamics. $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ introduce diffusion, drift, and pairwise behaviors in the flow, respectively. See Section 3.3 for further examples.
3 Wasserstein Barycenters as Gradient Flows
In this section, we describe a new method for computing empirical OT barycenters. We use,
[…]
First, we derive the empirical version of the continuity equation (c.f., Equation 10). For $\hat{P}_{t}\in\text{Emp}_{n}(\Omega)$:
$\partial_{t}\hat{P}_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\partial_{t}\delta_{z_{t,i}^{(P)}}=-\dfrac{1}{n}\sum_{i=1}^{n}\text{div}(\dot{z}_{t,i}^{(P)}\delta_{z_{t,i}^{(P)}}),$ (13)
$=-\text{div}(\hat{P}_{t}v_{t}),$
where we identified $v_{t}(z_{t,i}^{(P)})=\dot{z}_{t,i}^{(P)}$. Second, we discretize the gradient flow in time, through an index $\tau=\alpha t$, $\alpha>0$. Using the forward Euler scheme,
$\dot{z}_{t,i}^{(P)}=\frac{z_{\tau+1,i}^{(P)}-z_{\tau,i}^{(P)}}{\alpha}\rightarrow z_{\tau+1,i}^{(P)}=z_{\tau,i}^{(P)}+\alpha v_{\tau,i},$ (14)
where $v_{\tau,i}=v_{\tau}(z_{\tau,i}^{(P)})$. We can obtain different update strategies based on the velocity field,
$v_{\tau,i}$ $=-\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau}),$ (15)
$v_{\tau,i}$ $=-\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau})+\sqrt{2\alpha\eta_{G}}\eta_{\tau,i},$ (16)
$v_{\tau,i}$ $=u_{\tau,i},\quad u_{\tau+1,i}=\beta u_{\tau,i}-\alpha\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau}).$ (17)
Equations 15, 16 and 17 correspond to the steepest descent, Langevin dynamics, and momentum descent strategies, respectively. $\eta_{G}\geq 0$ denotes the diffusion coefficient. The Langevin dynamics version of the flow corresponds to the NPGD dynamics of chizat2025doubly. Furthermore, for empirical measures, the functionals in Equation 12 are tractable, and take the following form,
[…]
where $\hat{p}$ denotes the density of $\hat{P}$. Since $\hat{P}$ is empirical it is, in general, difficult to compute $\mathbb{G}$. A notable exception is the entropy functional (i.e., $G(s)=s\log s$), which adds a diffusion term to the gradient flow dynamics. This corresponds to Langevin dynamics [chizat2025doubly] (Equation 16). For general choices of $G$, one needs to estimate the density $\hat{p}$ at each iteration of the empirical flow.
Algorithm 1 Empirical Time Discretized Wasserstein Gradient Flow
[…]
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
[…]
4:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})+\mathbb{G}(\hat{P}_{\tau})$
6:  for $k\leftarrow 1$ to $K$ do
7:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i .i.d. from $Q_{k}$.
8:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{2,\epsilon}(\hat{P}_{\tau},\hat{Q}_{k})^{2}$
9:  end for
10:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}+\alpha v_{\tau,i}$
11:end for
An immediate consequence of Equation 18 is that our gradient flow formulation encompasses several previous works. For instance, with $G=U=V=0$, we have the classical free-support barycenters of cuturi2014fast. Furthermore, for the entropy functional and $\epsilon>0$, we have the NPGD algorithm of chizat2025doubly. Finally, as we show in Section 3.2 below, with a specific metric on $\Omega=\mathcal{X}\times\mathcal{Y}$, we retrieve the algorithm of montesuma2023multi.
3.1 Scaling with Mini-Batch OT
The computation of $\mathbb{B}_{\epsilon}$ in the discrete approach [cuturi2014fast] involves an OT problem between the complete set of samples in the input measures. Therefore, when the input measures are large scale, the barycenter problem becomes infeasible. We circumvent this issue with mini-batch OT [fatras2021minibatch].
The main idea behind our scalable barycenter algorithm is accessing $\hat{Q}_{k}\in\mathcal{Q}$ through sampling. At each iteration of Algorithm 1, we sample a mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ of $m$ samples from each input measure (cf. line 5 in Algorithm 1). Since we sample the same number of samples from each measure, we can actually vectorize the computation of the $K$ OT problems in the evaluation of $\mathbb{W}_{2,\epsilon}$.
Let $\{z_{\tau,i}^{(P)}\}_{i=1}^{n}$ be the barycenter support, and $\{z_{j}^{(Q_{k})}\}_{j=1}^{m}$ be a mini-batch from $Q_{k}$. The Gibbs kernel $E_{k,i,j}=e^{-(C_{k,i,j}/\epsilon)}$, and $C_{k,i,j}=d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{2}$ are arrays with shape $(K,n,m)$. The Sinkhorn iterations become,
$\psi_{k}^{(\ell+1)}=\dfrac{b}{E_{k}^{T}\phi_{k}^{(\ell)}}\text{, and }\phi_{k}^{(\ell+1)}=\dfrac{a_{k}}{E_{k}\psi_{k}^{(\ell+1)}},$ (19)
where the division operation should be understood coordinate-wise. Furthermore, $b_{j}=n^{-1}$ and $a_{k,i}=m^{-1}$ is the importance of each sample in the barycenter, and input measure support, respectively. For simplicity, we assume uniform weights. For $L\in\mathbb{N}$ iterations, we can retrieve $\gamma_{\epsilon,k}^{\star}=\text{diag}(\phi_{k}^{(L)})E_{k}\text{diag}(\psi_{k}^{(L)})$. Since each $\gamma_{\epsilon,k}^{\star}$ has shape $(n,m)$, we can actually stack it in an array of shape $(K,n,m)$, similarly to $E$ and $C$.
The main advantage of equation 19 is that the $K$ OT problems can be vectorized. This is possible because we draw the same number of samples from each input measure, making it possible to vectorize the Sinkhorn iterations. For instance, a similar strategy is possible for measures over a fixed grid [peyre2019computational, Remark 4.16]. The OT plan $\gamma_{\epsilon}$, obtained at the end of the Sinkhorn iterations, is used to compute $\mathbb{B}_{\epsilon}$, in lines 4 through 7 in Algorithm 1.
3.2 Flows over Joint Measures
An interesting advantage of the gradient flow formalism is that the only necessary ingredient is a differentiable metric over $\Omega$. Therefore, we generalize the setting of cuturi2014fast, who assumed $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. As we show in our experiments, in machine learning applications one has $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Let $z=(x,y),x\in\mathcal{X},y\in\mathcal{Y}$,
$d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (20)
where $\beta\geq 0$ is a parameter that balances the feature distance terms and the label distance terms.
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in montesuma2021wasserstein and alvarez2021dataset, is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{n_{\text{classes}}}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
$y_{i,c}^{(P)}=\text{softmax}(\ell_{i,1}^{(P)},\cdots,\ell_{i,n_{c}}^{(P)})=\dfrac{\exp(\ell_{i,c}^{(P)})}{\sum_{c=1}^{n_{c}}\exp(\ell_{i,c}^{(P)})},$
thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$, $x\in\mathbb{R}^{d}$, $\ell\in\mathbb{R}^{n_{\text{classes}}}$. From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{\text{hard},i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
3.3 Task-Aware Regularizing Functionals
One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by alvarez2021dataset for transfer learning problems. Here we propose the following functionals,
$V_{E}(z)$ $=-\eta_{V}\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (21)
$U_{R}(z,z^{\prime})$ $=\begin{cases}\eta_{U}h(d(x,x^{\prime}))&\text{if }y_{\text{hard}}\neq y_{\text{hard}}^{\prime},\\ 0&\text{otherwise}\end{cases},$ (22)
where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below, and $\eta_{V}>0$ and $\eta_{U}>0$ are scaling parameters. For instance, in our experiments we use the hinge loss, $h(u)=\text{max}(0,\text{margin}-d)$, where $\text{margin}\geq 0$ is a margin parameter. Equations 21 and 22 correspond to entropy and repulsion, respectively. The first functional penalizes barycenters that have fuzzy labels. The second functional encourages classes to be well separated.
3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 12 comes from the fact that $P\mapsto\mathbb{W}_{2}(P,Q)^{2}$ is not geodesically convex in $\mathcal{P}_{2}(\Omega)$ [santambrogio2017euclidean, Section 4.4]. As a result, we are minimizing a non-convex functional. Explicit constants and proofs are available in Appendix D. Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [polyak1964gradient, karimi2016linear],
$\lVert\nabla_{W}\mathbb{B}(P)\rVert_{L_{2}(P)}^{2}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}).$ (23)
When this inequality holds, we have the following convergence result.
Theorem 3.1.
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball in $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2}(\Omega)$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2}(\Omega)$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Approximate each $Q_{k}$ with $m$ i.i.d. samples $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$, $z_{i}^{(Q_{k})}\sim Q_{k}$ and $\{z_{\tau,i}^{(P)}\}_{i=1}^{n}$ be the barycenter support. Under inequality 42 the following holds,
$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(\hat{P}_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,n}}{m}}\end{aligned},$ (24)
where the expectation is taken with respect to samples from $Q_{k}$. The constants $C_{R}$ and $C_{d,n}$ depend on the radius $R$, number of barycenter samples $n$, and dimensions $d$. $m$ denotes the number of samples drawn from each $Q_{k}.$
The first term in the r.h.s. of equation 24 decays exponentially with the iterations $\tau$, and is a direct consequence of the PL inequality in equation 23. The second term captures the empirical approximation error of the involved measures, that is, $Q_{k}\in\mathcal{Q}$ and $P_{\tau}$. In practice, $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximations.
On the PL Inequality. Equation 23 is a standard tool used in proving the convergence of non-convex optimization algorithms, however, it is an open question whether it holds in general settings. For instance, chewi2020gradient shows that this inequality holds for Guassian measures under specific spectral conditions. In the Appendix, we show that the same result holds for location-scatter families of measures (Appendix D.2), based on the isometry between these sets, and the Bures-Wasserstein manifold established by [alvarez2016fixed, Theorem 2.3]. On top of this theoretical justification, we provide an experiment validating the bound on the Swiss roll measure (Appendix D.1).
(a) $Q_{0}$.
(b) $Q_{1}$.
(c) $Q_{2}$.
(d) $Q_{3}$.
(e) $Q_{4}$.
(f) $P^{\star}$.
Figure 2: Swiss roll measures in a location-scatter family, where a source measure is transformed under an affine function $T_{k}(x)=A_{k}x+b_{k}$. Colors reflect the relative position of points in the manifold.
4 Experiments
In the following, all experiments were conducted on a virtual machine with 8 CPUs (AMD EPYC 7413, 48GB of RAM) and an NVIDIA L4 GPU (24GB of VRAM). Appendix F includes additional details, fine-grained results, complexity analysis and ablations.
4.1 Swiss Roll Measures
In Figure 2, we show the Swiss roll measure [korotin2021continuous] $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$, $T_{k}(x)=A_{k}x+b_{k}$, for which the ground-truth barycenter is known in closed-form [alvarez2016fixed], that is, $P^{\star}=(\sum_{k}\lambda_{k}T_{k})_{\sharp}Q_{0}$. These measures are shown in Figure 2.
(a) Discrete.
(b) CW2B.
(c) U-NOT.
(d) NormFlow.
(e) WGF.
[…]
Figure 3: Comparison of unsupervised barycenter methods. Solvers (b, c, d) are neural, and (e, f) are ours.
We compare 5 barycenter solvers, including discrete solvers [cuturi2014fast, montesuma2023multi], and neural solvers: CW2B [korotin2021continuous], U-NOT gazdieva2024robust, and NormFlow visentin2025computing. First, we benchmark unsupervised solvers in Figures 3 (a) through (f). Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our Wasserstein Gradient Flow (WGF) algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. Overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization is more complicated and very sensitive to hyper-parameters.
Second, we experiment with integrating labels in the ground cost as described in Section 3.2. These methods are shown in Figures 4 (a), (b) and (c). In all cases, integrating labels produces barycenters that are closer to the ground-truth. We conclude that using the labels gives a strong inductive bias in the barycenter computation, which explains the gain in performance of labeled barycenters in the next section.
(a) Discrete.
(b) WGF.
[…]
Figure 4: Comparison of supervised barycenter methods.
Next, we compare the running time of our WGF solver with the discrete solver of cuturi2014fast, for an increasing barycenter support size $n\in\{2^{10},\cdots,2^{16}\}$ and batch size $m\in\{2^{8},2^{9},2^{10}\}$ for $200$ iterations. The discrete solver is run for $n\in\{2^{10},\cdots,2^{14}\}$. For this algorithm, larger support sizes results in memory overflow. Our WGF is able to compute larger barycenters via mini-batching.
We summarize our results in Figure 5. First, in Figure 5 (a), our WGF leads to speedups ranging from $2\times$ $(n=2^{12},m=2^{10})$ to $50\times$ $(n=2^{14},m=2^{8})$. Similar gains were obtained using Exact OT on CPU (Appendix F.3). Second, in Figure 5 (b), we isolate the gain of using GPU acceleration and our vectorization strategy (Section 3.1), showing consistent gains ranging from $2.1\times$ to $8.3\times$. Overall, our WGF strategy provides compounding speedups from three sources: (i) minibatching, (ii) entropic regularization, and (iii) GPU parallelism.
(a) Scaling with $m$ ($\epsilon>0$)
(b) Scaling with $\epsilon$ ($m=256$)
Figure 5: Running time analysis for our WGF algorithm. In (a), we compare the running time of our technique in comparison with the discrete solver of cuturi2014fast. In (b), we isolate the effect of GPU acceleration, showing consitent gains with increasing support size.
4.2 Multi-Source Domain Adaptation
One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [kouw2019review, sun2015survey, pan2009survey]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk or error in the target domain measure,
$\mathcal{R}_{Q_{T}}(h)=\mathbb{E}_{(x,y)\sim Q_{T}}[\mathcal{L}(y,h(x))],$
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data (Appendix E.2).
Experimental Setup. The unsupervised WGF is fit with unlabeled measures, i.e., $Q_{k}=n_{k}^{-1}\sum_{i}\delta_{x_{i}^{(Q_{k})}}$, where $x_{i}^{(Q_{k})}$ are embedding vectors. The supervised WGF is fit with labeled measures, that is, $Q_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}\delta_{(x_{i}^{(Q_{k})},y_{i}^{(Q_{k})})}$. In both cases, $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{(x_{i}^{(P)},y_{i}^{(P)})}$ which allows us to apply the regularizing functionals defined in section 3.3.
Wasserstein Barycenter Transport. montesuma2021wasserstein proposed a technique based on Wasserstein barycenters for domain adaptation. The idea is to map the synthesized barycenter support to the target domain through the barycentric mapping [courty2016optimal]. Then, a classifier can be learned on the target domain based on the transported data. We adapt this idea to unsupervised barycenter algorithms by using the computed barycenter as a pivot domain. See appendix E.4 for more details.
Benchmarks. We run our experiments on five benchmarks: Office 31 [saenko2010adapting], Office Home [venkateswara2017deep], BCI-CIV-2a [brunner2008bci], ISRUC [khalighi2013automatic], and TEP [montesuma2024benchmarking]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Benchmark Backbone # Samples # Domains # Dim. # Classes
Office31 ResNet50 3287 3 2048 31
BCI-CIV-2a CBraMod 5184 10 200 4
TEP CNN 17289 6 128 29
Office Home ResNet101 15500 4 2048 65
ISRUC CBraMod 89240 100 512 5
Table 1: Overview of benchmarks used for domain adaptation, sorted by number of samples.
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. Besides those used in the previous section, we include the NOT approach of kolesov2024estimating, and the Gaussian Mixture Model (GMM) barycenter of montesuma2024lighter. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT turrisi2022multi, DaDiL-R, and E montesuma2023multi, and GMM-DaDiL montesuma2024lighter. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchmark), and we measure the classification accuracy, i.e., the percentage of correct predictions.
Main Results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Appendix F. In general, labeled barycenter methods (the 3 last rows in Table 2) have a clear advantage over unsupervised methods. We argue that label information is essential in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter]. Among barycenter methods, our WGF method achieves the best performance in all benchmarks, outperforming previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC, BCI-CIV, and TEP benchmarks.
Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP Avg. Rank
Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN -
Source-Only - 86.40 75.95 50.30 76.63 78.48 -
WJDOT - 86.80 76.59 N/A 76.95 86.13 7.8
DaDiL-R - 89.91 77.86 53.41 74.68 86.14 6.4
DaDiL-E - 89.79 78.14 N/A 75.89 85.87 5.0
GMM-DaDiL - 90.63 78.81 57.10 75.47 86.85 4.4
Discrete 87.82 77.17 57.38 70.67 83.81 8.8
NormFlow 85.91 76.73 55.45 78.04 82.89 10.0
CW2B 86.37 76.44 57.25 75.84 85.83 9.0
NOT 86.22 75.36 57.29 76.39 84.93 9.4
U-NOT 86.97 76.85 57.17 77.06 85.43 7.8
[…]
Discrete 87.93 77.09 57.43 78.20 86.09 5.0
GMM 88.54 77.87 57.04 74.58 84.67 8.2
[…]
Table 2: Average classification accuracy on target domains. The top group (WJDOT through GMM-DaDiL) include general domain adaptation methods. The middle and bottom groups show barycenter solvers without and with access to source labels ($\mathcal{X}\times\mathcal{Y}=$✗, $\mathcal{X}\times\mathcal{Y}=$✓), respectively. Bold indicates best among barycenter methods.
Unsupervised vs. Supervised Barycenters. An intriguing finding of Table 2 is that, on some benchmarks, unsupervised WGF achieves decent performance, even outperforming supervised baselines (e.g., WJDOT of turrisi2022multi). We probe the reason why these methods have unreasonably good performance, by performing a t-SNE [maaten2008visualizing] visualization of the barycenter against the source domain samples, shown in Figure 6. We see that even without labels, the barycenter samples cluster around the classes. This effect stems from the geometry of the embeddings, which were optimized (through supervision on source domains) to separate classes. As a result, unsupervised OT already captures, to some extent, the cluster structure of the barycentric measure. Figures 6 (b) and (c) investigate the remaining performance gap, as we transport the measures towards the target. Figure 6 (b) shows the transported sources to the barycentric domain (left) and the synthesized barycenter (right), and Figure 6 shows the final transported samples in the target domain for the unsupervised strategy (left) and supervised one (right). We conclude that using a synthesized supervised barycenter is better than using an unsupervised barycenter as pivot domain (Appendix E.4). Therefore, the inductive bias of using labels in the ground cost is essential for domain adaptation. This further explains the performance of DaDiL methods, which directly synthesize the target domain through an approximate barycenter.
(a) Unsupervised vs. Supervised barycenters.
(b) Transported sources vs. supervised barycenter.
(c) Transported samples in the target domain.
Figure 6: Comparison of unsupervised vs. supervised barycenter transports on the TEP benchmark.
5 Conclusion
We have presented a new Wasserstein barycenter algorithm (Algorithm 1), inspired by the gradient flow literature santambrogio2015optimal, that addresses two persistent challenges: scalability and regularization. We have derived a new algorithm that is amenable to mini-batch OT fatras2021minibatch (Section 3.1). Under entropic OT, we further demonstrate that OT computations can be vectorized across the $K$ input measures, leading to further speedups (Figure 5).
Our framework is flexible, and accommodates modular plug-in regularization from internal, potential, and interaction energies (Section 3.3). Extensive experiments on domain adaptation demonstrate the value of our approach and that incorporating labels in the OT ground-cost is necessary for state-of-the-art performance (Table 2).
Finally, our work opens interesting research questions. First, our framework can be extended to more complex differentiable structures, such as Riemannian manifolds. Second, it opens the question on how to effectively integrate labels and other regularization schemes into neural network solvers.
Appendix A Introduction
The goal of this supplementary material is to provide additional background on the main paper, as well as to provide missing proofs, additional experiments, and the complete experimental setting. These details were left out of the main paper due to length constraints.
Organization. We divide this supplementary material as follows. Section B provides a brief survey on the methods for computing Wasserstein barycenters. Section C.1 provides additional background on OT. Section C.2 covers additional details on the potential, internal, and interaction energies, as well as the barycenter objectives and their gradients. Section D provides missing proofs. Finally, section F provide additional experiments.
Appendix B A Brief Survey on Wasserstein Barycenters
B.1 Introduction
Given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, the Wasserstein barycenter problem was first introduced by agueh2011barycenters as the following minimization problem,
$P^{\star}=\underset{P\in\mathcal{P}_{2}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{\epsilon}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2,\epsilon}(P,Q_{k})^{2}\biggr\},$ (25)
where $\epsilon\geq 0$ denotes the entropic regularization weight. This minimization problem is the analogous, in Wasserstein space, to the Euclidean barycenter problem $x\mapsto\sum_{k}\lambda_{k}d(x,x_{k})^{2}$, which is made possible since $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ is a metric space. Solving problem 25 is, in general, challenging. Except for the Gaussian case, there are no closed-form solutions. In the following, we review a few strategies for computing Wasserstein barycenters.
B.2 Discrete Barycenter Solvers
Discrete barycenter solvers are based on the empirical approximation of the underlying measures, that is,
$\hat{P}(z)=\sum_{i=1}^{n}a_{i}\delta(z-z_{i}^{(P)})\in\text{Emp}_{n}(\Omega).$ (26)
Here, $\{z_{i}^{(P)}\}_{i=1}^{n}$ is the support of $\hat{P}$ and $a=\{a_{i}:a_{i}\geq 0\text{, and }\sum a_{i}\}_{i=1}^{n}$ are the sample weights. There are two prevalent strategies when discretizing $P$. First, one can fix a grid/bins over the underlying space $\Omega$, and compute $a_{i}=P(z_{i}^{(P)})$ as the bin weight. This strategy is equivalent to estimating the density of $P$ over $\Omega$. Second, we can sample $z_{i}^{(P)}\overset{\text{i.i.d.}}{\sim}P$, in which case $a_{i}=\nicefrac{{1}}{{n}}$. In this paper, we deal primarily with free-support Wasserstein barycenters. Arguably, the first free-support algorithm was proposed by cuturi2014fast. Their strategy comes from plugging equation 26 into 25,
$Z^{(P),\star}$ $=\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{arg min}}\,\biggr\{\mathbb{B}_{\epsilon}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{\epsilon,k,i,j}\lVert z_{i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}\biggr\},$ (27)
subject to $\gamma_{\epsilon,k}^{\star}=\underset{\gamma\in\Gamma(P,Q_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{ij}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}+\epsilon\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{ij}\log\gamma_{ij}$
At this point, one should compare equation 27 with equation 18 in the main paper. cuturi2014fast propose solving this equation via a block coordinate descent strategy. Indeed, assume $\gamma_{k}=\text{OT}(P,Q_{k})$ has been calculated (e.g., through linear programming), then,
$\nabla_{z_{i}^{(P)}}\mathbb{B}_{\epsilon}(P|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}(z_{i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}.$
As a consequence, $\nabla_{z_{i}^{(P)}}\mathbb{B}_{\epsilon}(P|\mathcal{Q})=0$ leads to an update rule in terms of the barycentric projection of $P$ to $Q_{k}$, which is a discrete approximation of the Monge map deb2021rates.
An important question comes when the samples from $P$ are feature-label joints, first raised by montesuma2023multi, who proposed to use the ground-cost,
$d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (28)
where $y$ and $y^{\prime}$ are one-hot encoded labels, and $\beta\geq 0$ strikes a balance between the features and labels terms. As we show in the main paper, since this ground-cost is a metric, computing the OT cost with $d^{2}$ as the ground-cost is equivalent to computing the 22-Wasserstein distance on $\mathcal{X}\times\mathcal{Y}$.
The first paper to remark these computations was montesuma2023multi, who proposed to couple these mappings,
$x_{\tau+1,i}^{(P)}=\sum_{k=1}^{K}\lambda_{k}\biggr(\underbrace{n\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}x_{j}^{(Q_{k})}}_{T_{P\rightarrow Q_{k}}^{\star}(x_{\tau,i}^{(P)})}\biggr)\text{, and, }y_{\tau+1,i}^{(P)}=\sum_{k=1}^{K}\lambda_{k}\biggr(\underbrace{n\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}y_{j}^{(Q_{k})}}_{T_{P\rightarrow Q_{k}}^{\star}(y_{\tau,i}^{(P)})}\biggr).$ (29)
These iterations have a few nice properties with respect to the label term. First, the mappings $T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)})$ correspond to label propagation terms first defined in redko2019optimal. Second, since the summations in equation 29 are convex combinations of the labels, the update rules naturally suffice $y_{\tau,i}^{(P)}\in\Delta_{n_{\text{classes}}}$. Third, these iterations can be understood as fixed-point iterations in the sense of alvarez2016fixed. The authors use this algorithm as the building block of their dictionary learning strategy.
B.3 Neural Barycenter Solvers
In this section, we cover a collection of methods that use neural networks as the foundation of computing Wasserstein barycenters. These methods are oftentimes referred to as "continuous", thus relying on the idea that the calculated barycenter is a continuous measure. Note that this is different than claiming that these barycenters solve a continuous problem. In fact, most methods are formulated in terms of expectations of the involved measures, which are approximated through mini-batching. Indeed, mini-batching is the main tool for scalability in Wasserstein barycenters.
Generative Modeling. One of the first methods was proposed by cohen2020estimating, who proposed viewing equation 25 through the lens of generative modeling. Let $g_{\theta}$ denote a neural network with parameters $\theta$. Then,
$\theta^{\star}=\underset{\theta\in\Theta}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P_{\theta},Q_{k})^{2},$
where $P_{\theta}=g_{\theta,\sharp}P_{0}$ is the push-forward of a latent measure $P_{0}$ by the generator $g_{\theta}$. Therefore, the Wasserstein barycenter is parametrized via the weights of $g_{\theta}$. In comparison, methods such as cuturi2014fast and montesuma2023multi parametrize the barycenter through its samples. The underlying idea is sampling from each $Q_{k}$ and $P_{0}=\mathcal{N}(0,\sigma^{2}I)$.
Input Convex Neural Nets. More refined settings exploit the properties of the 2-Wasserstein distance with the squared Euclidean ground-cost. Indeed, from [makkuva2020optimal, Equation 4], exploiting the decomposition
$\lVert z-z^{\prime}\rVert_{2}^{2}=\lVert z\rVert_{2}^{2}+\lVert z^{\prime}\rVert_{2}^{2}-2\langle z,z^{\prime}\rangle,$
we can obtain,
$\dfrac{1}{2}\mathbb{W}_{2}(P,Q)^{2}=\underbrace{\int\dfrac{\lVert z\rVert_{2}}{2}dP(z)+\int\dfrac{\lVert z\rVert_{2}}{2}dQ(z)}_{\mathbb{C}(P,Q)}-\underset{f,g\in\Phi}{\text{min}}\,\int f(z)dP(z)+\int g(z)dQ(z),$ (30)
where $(f,g)\in\Phi=\{f(x)+g(y)\leq\lVert x-y\rVert_{2}^{2}\}$ are called potentials. The main insight of makkuva2020optimal is rewritting the minimization problem in the r.h.s. of equation 30 as,
$\underset{f\in\text{CVX}}{\text{sup}}\,\underset{g\in\text{CVX}}{\text{inf}}\,\underbrace{-\int f(z)dP(z)-\int(\langle z,\nabla g(z)\rangle-f(\nabla g(z)))dQ(z)}_{V_{P,Q}(f,g)}$
Based on these equations, fan2020scalable propose solving,
[…]
Here, samples in the Wasserstein barycenter are generated through the forward pass $h(z)$, where $z\sim P_{0}=\mathcal{N}(0,\sigma^{2}I)$. This is a min-max-min problem, which, as with many problems involving neural nets (e.g., generative adversarial nets [goodfellow2014generative]), is subject to saddle points.
An important aspect of equation 31 is the minimization with respect convex functions $f_{k}$ and $g_{k}$. Indeed, these are the potentials of the transportation problem between the barycenter $P$ and each $Q_{k}$. More specifically, In light of Brenier’s theorem brenier1991polar, $\nabla g_{k}$ is the Monge map from $Q_{k}$ to the barycenter $P$. These elements motivate fan2020scalable using Input Convex Neural Networks (ICNNs), as in makkuva2020optimal, for the optimization problem in equation 31.
Continuous 2-Wasserstein Barycenter. We now cover a family of methods relying on the fixed-point view of Wasserstein barycenters introduced by alvarez2016fixed. Indeed, these authors show that the Wasserstein barycenter is a fixed point of the operator $\Psi:P\mapsto(\sum_{k}\lambda_{k}T_{P\rightarrow Q_{k}})_{\sharp}P$. This is actually the principle behind most of the discrete methods described in the previous section. Now, with respect to OT potentials, if $\Psi(P)=P$, then one has,
$\sum_{k=1}^{K}\lambda_{k}T_{Q_{k}\rightarrow P}^{\star}(z)=z\overset{(a)}{\implies}\sum_{k=1}^{K}\lambda_{k}\nabla g_{k}(z)=z\overset{(b)}{\implies}\sum_{k=1}^{K}\lambda_{k}g_{k}(z)=\dfrac{\lVert z\rVert_{2}^{2}}{2}+c$
where implication (a) follows from Brenier’s theorem, and $g_{k}$ is the potential of the transportation problem from $Q_{k}$ to $P$, and implication (b) follows by integration. Here, since $c$ is an integration constant, one can conveniently take $c=0$. Now, denote $\text{Congruent}(\lambda)=\{g_{1:K}=(g_{1},\cdots,g_{K}):\sum_{k}\lambda_{k}g_{k}(z)=\nicefrac{{\lVert z\rVert_{2}^{2}}}{{2}}\}$ be the set of congruent potentials $g_{1:K}=(g_{1},\cdots,g_{K})$ with respect to coordinates $\lambda$. Using the dual formulation in equation 30 and restricting the set of potentials to congruent potentials,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]+\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{g_{1:K}\in\text{Congruent}(\lambda)}{\text{min}}\,\biggr(\sum_{k=1}^{K}\lambda_{k}(\mathbb{E}_{z\sim Q_{k}}[f_{k}(z)]+\mathbb{E}_{z\sim P}[g_{k}])(z)\biggr),$
where $g_{1:K}=(g_{1},\cdots,g_{K})$. Now, taking the expectation on the congruence constraint, we have,
$\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim P}[g_{k}(z)]=\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr],$
which allowed korotin2021continuous to simplify $\mathbb{B}(P|\mathcal{Q})$ into,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{g_{1:K}\in\text{Congruent}(\lambda)}{\text{min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}[g_{k}(z)],$ (32)
which does not involve terms with respect to the barycentric measure $P$. While the approach of korotin2021continuous simplifies the barycenter objective to expectations with respect to the known measures $Q_{k}$, it introduces a challenging constraint – the congruence of potentials. This constraint is challenging, as it is a functional inequality over the whole space $\Omega$. Furthermore, it adds to the already constrained setting of fitting convex functions. korotin2021continuous proposes solving it approximately, by penalizing potentials that do not meet the congruence condition. There is a circular dependence here, as one needs an estimate of the Wasserstein barycenter to estimate the constraint violation.
Neural (Weak) Optimal Transport. Weak OT gozlan2017kantorovich, backhoff2019existence comes from the interpretation of the Kantorovich formulation as,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}C(z,\gamma(\cdot|z))dP(z)\text{, where, }C(z,Q)=\int_{\Omega}c(z,z^{\prime})dQ(z^{\prime}),$ (33)
this formulation allows for more general costs $C:\Omega\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$, i.e., costs that take a sample and a probability measure as inputs. In this sense, kolesov2024estimating extends equation 32 to,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma(\cdot|z)}(g_{k}(z^{\prime}))]\biggr),$ (34)
which holds for congruent potentials $\sum_{k}\lambda_{k}g_{k}=0$. As with equation 32, $\mathbb{B}$ does not depend explicitly on the unknown barycentric measure $P$. As such, the authors compute the barycenter through a max-min optimization problem,
$\underset{\{g_{1:K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma_{k}(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}(g_{k}(z^{\prime}))]\biggr)\biggr\}.$
This extension opens up new possibilities for weak OT barycenters. For instance,
$C_{\text{classical}}(x,\gamma)=\dfrac{1}{n}\sum_{j=1}^{n}c(z_{i}^{(P)},z_{j}^{(Q)})\quad C_{\text{KL}}(x,\gamma)=\dfrac{1}{n}\sum_{j=1}^{n}c(z_{i}^{(P)},z_{j}^{(Q)})+\epsilon\text{KL}(\mathcal{N}(\mu(z),\sigma(z))|P_{0}),$
where $C_{\text{classical}}$ allows us to retrieve the original OT problem. $C_{\text{KL}}$ is an alternative, when modeling $\gamma(\cdot|z)=\mathcal{N}(\mu(z),\sigma(z))$. On top of this reformulation, the authors in kolesov2024estimating propose approximating $\gamma(\cdot|z)$ through a mapping, as in korotin2022neural, that is,
$\gamma(\cdot|z)=T(z,s),\text{ subject to }s\sim\mathbb{S},$
where $\mathbb{S}$ is some prior measure, such as $\mathbb{S}=\mathcal{N}(0,\text{I})$. In this context, gazdieva2024robust extends this idea in the semi-unbalanced OT setting liero2018optimal, where one of the mass preservation constraint of one of the marginals of $\gamma$ is relaxed. More specifically, given $\xi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$,
$D_{\xi}(P||Q)=\int_{\Omega}\xi\biggr(\dfrac{P(z)}{Q(z)}\biggr)dQ(z)\text{ if }P\ll Q\text{ and }+\infty\text{ otherwise},$
where $P\ll Q$ means that $P$ is absolutely continuous with respect to $Q$. We denote $\bar{\xi}(z^{\prime})=\text{sup}_{z\in\mathbb{R}}\{zz^{\prime}-\xi(z)\}$ to be its convex conjugate. With that in mind, gazdieva2024robust shows, in Theorem 1 and Corollaries 1 and 2, that minimizing,
$\underset{P\in\mathcal{P}_{2}}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\biggr(\int_{\Omega}\int_{\Omega}c(z,z^{\prime})d\gamma(z,z^{\prime})+D_{\xi}(\gamma(\cdot|z)||Q_{k})\biggr),$
is equivalent to the following max-min problem,
$\underset{\{g_{1:K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}\biggr[-\bar{\xi}_{k}\biggr(-\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}[c_{k}(z,z^{\prime})-g_{k}(z^{\prime})]\biggr)\biggr]\biggr)\biggr\}.$
Normalizing Flows. This approach was proposed by visentin2025computing, and relies on the analysis of gangbo1996geometry. Their main idea is re-writing the Wasserstein distance in terms of functions $f,g$,
$\mathbb{W}_{2}(P,Q)^{2}=\underset{f\in\text{Borel}(P_{0},P),g\in\text{Borel}(P_{0},Q)}{\text{inf}}\,\lVert f-g\rVert_{L_{2}(P_{0})}^{2},$ (35)
where $\text{Borel}(P,Q)$ is the set of Borel functions such that $f_{\sharp}P=Q$. Naturally, a solution $(f^{\star},g^{\star})$ to the previous problem gives the transport map $T=g^{\star}\circ f^{-1}$, where $f^{-1}$ is the inverse of $f^{\star}$. Furthermore, $P_{0}$ is an arbitrary latent measure, which further links the approach with generative modeling (i.e., take $P_{0}=\mathcal{N}(0,\text{I})$).
Based on equation 35, the insight of visentin2025computing is modeling $f$ and $g$ through a conditional function $f(z,s)$, $s\in\{0,1\}$. The Wasserstein barycenter problem is a straightforward generalization of the $2$ measures case. For instance, the authors consider $S=\{1,\cdots,K\}$, which means that,
$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(g_{\sharp}P_{0},Q_{k})^{2}=\sum_{k=1}^{K}\lambda_{k}\lVert g-f(z,k)\rVert_{L_{p}(P_{0})}^{2},$
where $g$ is the push-forward of $P_{0}$ to the barycenter $P^{\star}$, and $f(z,k)$ is the push-forward of $Q_{k}$ to $P^{\star}$. As the authors show in [visentin2025computing, Theorem 3.3],
$g(z)=\sum_{k=1}^{K}\lambda_{k}f(z,k).$
The main challenge in visentin2025computing is formulating a conditional model $f:\Omega\times S$. As the authors discuss, one possibility would be using $K=|S|$ different neural nets for modeling $f(z,\cdot)$, which is similar to what is done, for instance, in the neural net methods. The authors take a different route and use conditional normalizing flows kobyzev2020normalizing to model these functions.
B.4 Future Directions
From the previous discussion on neural net-based methods for computing Wasserstein barycenters, we can identify a few common design choices. For instance, korotin2021continuous, kolesov2024estimating and gazdieva2024robust use $\mathcal{O}(K)$ neural nets for computing Wasserstein barycenters. This design choice results in a parametrization of the underlying barycenter that grows with the number of input measures. The only neural net-based method that differs from this idea is visentin2025computing, which uses a conditional model $f:\Omega\times S$.
Meanwhile, as we cover in our experiments (C.f., Section 4 in the main paper), there is a clear advantage of incorporating labels in the ground-cost when these are available. Unfortunately, it is not clear how to integrate label information in the existing methods. For instance, one could consider using the continuous parametrization in terms of logits $\ell$, as we described in Section 3.3. However, there are a few caveats,
  1. 1.
    Normalizing flows require invertible transformations. However, the softmax operation for getting the labels, i.e., $y=\text{softmax}(\ell)$, is not invertible.
  2. 2.
    Using ICNNs, one can enforce the convexity of $z=(x,\ell)$. However, applying the softmax breaks the convexity requirement.
As a result, adapting neural net based methods for the joint $\mathcal{X}\times\mathcal{Y}$ space is challenging and not straightforward. We leave this question for future work.
Appendix C Additional Background
C.1 Calculus in Wasserstein Spaces
In this section, we consider $\Omega=\mathbb{R}^{d}$, $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$ and $p=2$ fixed. While our algorithm is more general than this setting, we use these restrictions as they are the most analyzed by current literature. We leave the theoretical results for general metrics and $p$ to future work.
Notation. For the following theoretical analysis, we use the following notation. Given a measurable map $f,g:\Omega\rightarrow\Omega$. Given a measure $P\in\mathcal{P}_{2}(\Omega)$, we denote,
$\lVert f\rVert_{L_{2}(P)}^{2}=\int_{\Omega}\lVert f(x)\rVert_{2}^{2}dP(x).$
Likewise, we have a notion of inner product,
$\langle f,g\rangle_{P}=\int_{\Omega}\langle f(x),g(x)\rangle_{\Omega}dP(x).$
Definition C.1.
Given a functional $\mathbb{F}:\mathcal{P}_{2}(\Omega)(\mathbb{R}^{d})\rightarrow\mathbb{R}$, we call $\delta\mathbb{F}(P):\mathbb{R}^{d}\rightarrow\mathbb{R}$, if it exists, the unique (up to additive constants) function such that
$\dfrac{d}{d\varepsilon}\mathbb{F}(P+\varepsilon\chi)\biggr|_{\epsilon=0}=\int\delta\mathbb{F}(P)d\chi,$ (36)
for every perturbation $\chi$ such that, at least for $\varepsilon\in[0,\varepsilon_{0}]$, the measure $P+\varepsilon\chi\in\mathcal{P}_{2}(\mathbb{R}^{d})$. This functional is called the first variation of $\mathbb{F}$ at $P$.
We refer readers to [chewi2024statistical, Chapter 5] and [ambrosio2008gradient, Chapter 8] for further details on gradient flows in Wasserstein spaces. The main tool we will use here is the continuity equation,
$\partial_{t}P_{t}=-\text{div}(P_{t}v_{t}),$
where the divergence operator should be understood in the measure theoretic sense. Here, $v_{t}:\Omega\rightarrow\Omega$ is a vector field. Now, given $\mathbb{F}:\mathcal{P}_{2}(\Omega)\to\mathbb{R}$ with first variation $\delta\mathbb{F}(P):\Omega\to\mathbb{R}$, the time derivative of $\mathbb{F}$ along the curve $\{P_{t}\}_{t\geq 0}$ can be computed as,
$\partial_{t}\mathbb{F}(P_{t})\overset{(1)}{=}\int\delta\mathbb{F}(P_{t})d(\partial_{t}P_{t})\overset{(2)}{=}-\int\delta\mathbb{F}(P_{t})\text{div}(P_{t}v_{t})dz\overset{(3)}{=}\int\langle\nabla\delta\mathbb{F}(P_{t}),v_{t}\rangle dP_{t}\overset{(4)}{=}\langle\nabla\delta\mathbb{F}(P_{t}),v_{t}\rangle_{P_{t}},$
where (1) is obtained by taking $\chi=\partial_{t}P_{t}$ in equation 36; (2) is obtained by applying the continuity equation; (3) is obtained by using integration by parts; finally, (4) is obtained via the definition of inner product. This motivates the equality,
$\nabla_{W}\mathbb{F}(P)=\nabla\delta\mathbb{F}(P).$ (37)
We refer readers to [chewi2024statistical, Proposition 5.10], for a more formal derivation. This discussion characterizes the Wasserstein gradient $\nabla_{W}\mathbb{F}(P):\Omega\to\Omega$ as a vector field. Therefore, taking the direction of steepest descent of the functional $\mathbb{F}$, $v_{t}=-\nabla_{W}\mathbb{F}(P_{t})$, leads to the following connection between $\partial_{t}\mathbb{F}(P_{t})$ and $\nabla_{W}\mathbb{F}(P_{t})$,
$\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla_{W}\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$ (38)
C.2 Wasserstein Gradient of Functionals
In this section, we give an overview on the Wasserstein gradient of the terms used in equation 16. We refer readers to [chewi2024statistical, Chapter 5] and [santambrogio2017euclidean, Section 4] for further details.
Potential Energy, $\mathbb{V}$. For reference, this functional corresponds to,
$\mathbb{V}(P)=\mathbb{E}_{z\sim P}[V(z)]=\int V(z)dP(z),$
for some potential function $V:\Omega\rightarrow\mathbb{R}$. From the definition of first variation,
$\mathbb{V}(P+\epsilon\chi)=\int V(z)dP(z)+\epsilon\int V(z)d\chi(z),$
then, taking the derivative with respect to $\epsilon$,
$\dfrac{d}{d\epsilon}\mathbb{V}(P+\epsilon\chi)=\int V(z)d\chi,$
which means that $V=\delta\mathbb{V}(P)$. As a consequence, the Wasserstein gradient of $\mathbb{V}$ is,
$\nabla_{W}\mathbb{V}=\nabla V.$
which is a vector field over $\Omega$. Since $\mathbb{V}$ is linear in $P$, it is convex. Furthermore, its continuity with respect to the weak convergence depends on the regularity of $V$. For instance, if $V$ is bounded, $\mathbb{V}$ is continuous. If $V$ is lower semi-continuous and bounded from below, then $\mathbb{V}$ is semi-continuous.
Internal Energy, $\mathbb{G}$. The internal energy takes the form,
$\mathbb{G}(P)=\int G(P(z))dz,$
where $P(z)$ is understood as the density of the measure $P$, with respect the Lebesgue measure. Let us assume $\chi$ admits a density with respect the Lebesgue measure as well. In this case,
$\mathbb{G}(P+\epsilon\chi)=\int G(P(z)+\epsilon\chi(z))dz.$
Taking the derivative under the integral sign,
$\dfrac{d}{d\epsilon}\mathbb{G}(P+\epsilon\chi)\biggr|_{\epsilon=0}=\int G^{\prime}(P(z))\chi(z)dz=\int G^{\prime}(P(z))d\chi(z),$
which means that $\delta\mathbb{G}=G^{\prime}(P)$ and $\nabla_{W}\mathbb{G}=\nabla G^{\prime}(P(z))$.
Interaction Energy, $\mathbb{U}$. Here, let us assume $U(z,z^{\prime})$ symmetric for simplicity,
$\mathbb{U}(P)=\int\int U(z,z^{\prime})dP(z)dP(z^{\prime}),$
hence,
$\mathbb{U}(P+\epsilon\chi)=\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})+2\epsilon\int\int U(z,z^{\prime})dP(z)d\chi(z^{\prime})+\mathcal{O}(\epsilon^{2}).$
Again, taking derivatives with respect $\epsilon$,
$\dfrac{d}{d\epsilon}\mathbb{U}(P+\epsilon\chi)\biggr|_{\epsilon=0}=2\int\biggr(\int U(z,z^{\prime})dP(z)\biggr)d\chi(z^{\prime}),$
which means that,
$\delta\mathbb{U}(P)=2\int U(z,z^{\prime})dP(z).$
When $U(z,z^{\prime})=h(z-z^{\prime})$ for an even $h$, $\int U(z,z^{\prime})dP(z)=\int h(z-z^{\prime})P(z)dz=h\star P$, i.e., the convolution of $h$ and $P$. Furthermore, $\nabla_{W}\mathbb{U}=(\nabla h)\star P$
Barycenter Functional, $\mathbb{B}$. For the barycenter functional, we can derive its first variation from the 22-Wasserstein distance. Assume there exists an OT map $T_{P\rightarrow Q}^{\star}$. In this case,
$\nabla_{W}\mathbb{W}_{2}(P,Q)^{2}=2(\text{Id}-T_{P\rightarrow Q}^{\star}).$ (39)
Therefore, given $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$, it is straightforward to obtain:
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ $=\nabla_{W}\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr)=2\sum_{k=1}^{K}\lambda_{k}(\text{Id}-T_{P\rightarrow Q_{k}}^{\star}).$ (40)
A Connection with [cuturi2014fast, Algorithm 2]. Let us now focus on the empirical case with an Euclidean ground-cost. Assume $\gamma^{\star}$ computed between $\hat{P}$ and $\hat{Q}$, which means that,
$\mathbb{W}_{2}(\hat{P},\hat{Q})^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}^{\star}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}\text{, and }\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}\lVert z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}.$
We can compute the Wasserstein gradient $\nabla_{W}\mathbb{B}$ by computing $\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}$, that is,
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}(z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{\tau,i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{K}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})}.$
From this last equation, we can use the barycentric mapping definition,
$T_{P\rightarrow Q_{k}}^{\star}(z_{i,\tau}^{P})=n\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})},$ (41)
to write,
$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=\dfrac{2}{n}\biggr(z_{\tau,i}^{(P)}-\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{\tau,i}^{(P)})\biggr),$
which means we can write the gradient flow update rule as,
$z_{\tau+1,i}^{(P)}=\biggr(1-\dfrac{2\alpha}{n}\biggr)z_{i,\tau}^{(P)}+\dfrac{2\alpha}{n}\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{i,\tau}^{(P)}),$
Which is the update rule in [cuturi2014fast, Algorithm 2] with $\theta=\nicefrac{{2\alpha}}{{n}}$. Furthermore, we can also instantiate the fixed-point algorithm of montesuma2023multi, for $\alpha=n/2$. We now highlight two differences of our work with respect these two strategies. First, our setting is more general. In fact, our algorithm is capable of handling ground metrics other than the squared Euclidean metric. Furthermore, our framework couples the barycenter objective with other functionals. Second, while cuturi2014fast assume that all samples of $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ are available at once, our algorithm assumes that these measures are available through sampling. This treatment is oftentimes called continuous in the literature korotin2021continuous, korotin2022wasserstein, however, one still relies on samples from these measures rather than some truly continuous form.
Appendix D Proof of Theorem 3.1
In this section, we present the proof for theorem 3.1. We mainly use two tools to obtain this result,
  1. 1.
    The exponential convergence given by the PL inequality,
    $\lVert\nabla_{W}\mathbb{B}(P)\rVert_{L_{2}(P)}^{2}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (42)
  2. 2.
    An uniform bound on $P\mapsto\mathbb{B}(P|\mathcal{Q})$ controlling the error of approximating $Q_{k}$ with $\hat{Q}_{k}$.
These results are available, respectively, in chewi2024statistical and portales2025sample. For completeness, we re-state and demonstrate them. We start with a modified version of [chewi2024statistical, Corollary 5.17],
Proposition D.1.
(Measure theoretic Gronwall’s Lemma chewi2024statistical) Let $\mathbb{F}:\mathcal{P}_{2}(\Omega)\rightarrow\mathbb{R}$ be a functional that satisfies the PL inequality (c.f. 42) with constant $C_{\text{PL}}>0$, then,
$\mathbb{F}(P_{\tau})-\mathbb{F}^{\star}\leq e^{-2C_{\text{PL}}\tau}(\mathbb{F}(P_{0})-\mathbb{F}^{\star}).$ (43)
Proof.
Define $\phi(t)=\mathbb{F}(P_{t})-\mathbb{F}^{\star}$. Then,
$\partial_{t}\phi(t)$ $=\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla_{W}\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$
$\geq-C_{\text{PL}}(\mathbb{F}(P_{0})-\mathbb{F}^{\star})=-C_{\text{PL}}\phi(t).$
From Grönwall’s Lemma [chewi2024statistical, Lemma 5.16],
$\underbrace{\mathbb{F}(P_{t})-\mathbb{F}^{\star}}_{\partial_{t}\phi(t)}\leq e^{-C_{\text{PL}}}\underbrace{(\mathbb{F}(P_{0})-\mathbb{F}(P)^{\star})}_{\phi(0)}$
We now present the uniform approximation bound of portales2025sample for empirical barycenters,
Theorem D.1.
(Barycenter Empirical Approximation error portales2025sample) Let $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ be a finite family of measures on $\mathcal{P}_{2}(\mathcal{B}(0,R))$ for $R>0$. Approximate each $Q_{k}$ with $m$ i.i.d. samples $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$. Let $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}$. Let $p\in[1,+\infty)$ and $\epsilon\geq 0$. Then,
$\mathbb{E}\biggr[\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{sup}}\,|\mathbb{B}_{\epsilon}(\hat{P}|\mathcal{Q})-\widehat{\mathbb{B}}_{\epsilon}(\hat{P}|\mathcal{Q})|\biggr]\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$ (44)
where,
$C_{p,R}$ $=\begin{cases}8\sqrt{2}\int_{0}^{3(2R)^{p}}\sqrt{\log\biggr(2+\dfrac{32p(2R)^{p}}{s}\biggr)}ds&\epsilon=0,\\ 8\sqrt{2}\int_{0}^{(4p+1)(2R)^{p}}\sqrt{\log(2+\frac{64p(2R)^{p}}{s})}ds&\epsilon>0,\end{cases},$
$C_{d,n}$ $=n(d+1).$
With these bounds established, we can then prove our Theorem 3.1.
Proof.
Let $\tau>0$ be the current optimization step. We have the following,
[…]
From Equation 45 we can bound the expected empirical optimization gap,
$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}|]\leq$ $\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\mathbb{B}(\hat{P}_{\tau})|]+\mathbb{E}[|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|]+\mathbb{E}[|\mathbb{B}^{\star}-\hat{\mathbb{B}}^{\star}|].$ (46)
Now, we can bound the first and third terms using the approximation error of portales2025sample (i.e., Equation 44),
$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\mathbb{B}(\hat{P}_{\tau})|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$ (47)
$\mathbb{E}[|\hat{\mathbb{B}}^{\star}-\mathbb{B}^{\star}|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$
here, we recall that $\mathbb{B}^{\star}=\mathbb{B}(P^{\star})=\text{inf}_{P}\mathbb{B}(P)$. Returning to the second term, note that the true optimization gap $\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})-\mathbb{B}^{\star}$ is deterministic, as it does not involve sampling from the measures in $\mathcal{Q}$. This means that,
$\mathbb{E}[|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|]=|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|\leq e^{-C_{\text{PL}}\tau}(\hat{\mathbb{B}}(\hat{P}_{0})-\hat{\mathbb{B}}^{\star}).$ (48)
Combining Equations 46, 47 and 48 we get,
$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(\hat{P}_{0})-\mathbb{B}^{\star})+C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$
which is the desired inequality. ∎
D.1 Empirical Verification of the PL Inequality
An open question in the literature is the verification of the PL inequality (Equation 42) for general measures, such as empirical ones. To give further context, current results prove that it indeed holds for Gaussian measures, under specific spectral conditions chewi2020gradient. Therefore, we do an empirical validation on whether this bound holds in the empirical case. In a nutshell, to test Equation 42, we need to compute the PL ratio,
$\text{PL}_{\text{ratio}}(\hat{P}_{\tau})=\dfrac{\lVert\nabla_{W}\mathbb{B}(\hat{P}_{\tau})\rVert_{L_{2}(\hat{P}_{\tau})}}{\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}},$ (49)
and verify that it indeed remains strictly positive over the course of our optimization. Computing the PL ratio in Equation 49 requires knowledge of $\mathbb{B}^{\star}$, which is, in general, intractable. However, for location-scatter families, it is known in closed form, since $P^{\star}$ itself is known. Therefore, we verify it in the context of the Swiss-roll measure, where all quantities are known. We show our results in Figure 7, which shows that the ratio converges with iterations to approximately $0.56$.
Figure 7: Empirical test of the PL inequality. From left to right, we show: (i) the optimality gap, (ii) the gradient norm, and (iii) the PL ratio as specified in Equation 49.
D.2 PL Inequality on Location-Scatter Family
An open question in the literature is the verification of the PL inequality (Equation 42) for general measures. Current results establish that it holds for Gaussian measures under specific spectral conditions [chewi2020gradient]. In this section, we show that the PL inequality transfers from Gaussians to the broader class of location-scatter families, which includes our Swiss roll experiments (Section 4).
Location-Scatter Families. Following alvarez2016fixed, let $Q_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be a measure with mean $\mu_{0}$ and covariance $\Sigma_{0}$. The location-scatter family generated by $Q_{0}$ is,
$\mathcal{L}(Q_{0}):=\left\{T_{\sharp}Q_{0}:T(x)=Ax+m,b\in\mathbb{R}^{d},A\in\mathcal{S}^{++}_{d}\right\}$, (50)
where $\mathcal{S}^{++}_{d}$ denotes the cone of $d\times d$ symmetric positive definite matrices. Each $Q\in\mathcal{L}(Q_{0})$ has mean $b$ and covariance $A$, but inherits the higher-order structure of $Q_{0}$.
Isometry to the Bures-Wasserstein Manifold. The key insight comes from alvarez2016fixed: for any $Q_{1},Q_{2}\in\mathcal{L}(Q_{0})$, the 2-Wasserstein distance equals
$\mathbb{W}_{2}^{2}(Q_{1},Q_{2})=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{2}^{1/2}\Sigma_{1}\Sigma_{2}^{1/2})^{1/2}),$ (51)
This is precisely the Wasserstein distance on the Bures-Wasserstein manifold: $\mathcal{BW}=\{\mathcal{N}(\mu,\Sigma):\mu\in\mathbb{R}^{d},\Sigma\in\mathcal{S}_{d}^{++}\}$ [takatsu2011wasserstein]. Henceforth, we note the elements of $\mathcal{BW}$ by $G_{\mu,\Sigma}$. In that sense $G_{\mu_{1},\Sigma_{1}}$ and $G_{\mu_{2},\Sigma_{2}}$ denote the Gaussian measures matching the moments of $Q_{1},Q_{2}\in\mathcal{L}(Q_{0})$. As a consequence, the restriction of $\mathbb{W}_{2}$ to $\mathcal{L}(Q_{0})$ is isometric to the Bures-Wasserstein manifold. Moreover, the optimal transport map between elements of $\mathcal{L}(Q_{0})$ takes the same form as for Gaussians:
$T^{\star}_{Q_{1}\to Q_{2}}(x)=\mu_{2}+A(x-\mu_{1}),\quad A=\Sigma_{1}^{-1/2}\left(\Sigma_{1}^{1/2}\Sigma_{2}\Sigma_{1}^{1/2}\right)^{1/2}\Sigma_{1}^{-1/2}.$ (52)
Barycenter Functional Equivalence. Let $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}\subset\mathcal{L}(Q_{0})$ with $Q_{k}=T_{k}Q_{0}$, such that $\mu_{k}$ and $\Sigma_{k}$ are the mean and covariance of $Q_{k}$. Let $\mathcal{G}=\{G_{\mu_{k},\Sigma_{k}}\}_{k=1}^{K}$ be the corresponding Gaussian measures. Equation 51 establishes:
$\mathbb{B}(P|\mathcal{Q})=\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\text{, where }P\in\mathcal{L}(Q_{0})\text{, and }G_{\mu,\Sigma}\in\mathcal{BW},$ (53)
where $\mu,\Sigma$ are the mean and covariance of $P\in\mathcal{L}(Q_{0})$. Furthermore, by alvarez2016fixed, the barycenter $P^{\star}\in\mathcal{L}(Q_{0})$ corresponds to the Bures-Wasserstein barycenter: if $P^{\star}$ is the minimizer of $P\mapsto\mathbb{B}(P|\mathcal{Q})$ over $\mathcal{L}(Q_{0})$ with mean $\mu^{\star}$ and covariance $\Sigma^{\star}$, then $G_{\mu^{\star},\Sigma^{\star}}$ is the minimizer of $G_{\mu,\Sigma}\mapsto\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})$.
Wasserstein Gradient Structure. The Wasserstein gradient of $\mathbb{B}$ at $P$ with mean and covariance $\mu$ and $\Sigma$ is (cf. Equation 40)
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\left(\mathrm{Id}-T^{\star}_{P\to Q_{k}}\right).$ (54)
Since each $T^{\star}_{P\to Q_{k}}$ is affine (Equation 52), the gradient $\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ is itself an affine map:
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})(x)$ $=2\sum_{k=1}^{K}\lambda_{k}(x-(A_{k}x+b_{k})),$
$=2\biggr(x-(\sum_{k=1}^{K}\lambda_{k}A_{k})(x)+\sum_{k=1}^{K}\lambda_{k}b_{k}\biggr),$
$=\underbrace{\biggr(2(\text{Id}-\sum_{k=1}^{K}\lambda_{k}A_{k})\biggr)}_{M}(x)+\underbrace{2\sum_{k=1}^{K}\lambda_{k}b_{k}}_{c},$
$=Mx+c,$ (55)
where $M=2(\text{Id}-\bar{A})$, $c=2(\bar{A}\mu-\bar{\mu})$, where $\bar{A}=\sum_{k=1}^{K}\lambda_{k}A_{k}$ and $\bar{\mu}=\sum_{k=1}^{K}\lambda_{k}\mu_{k}$. More importantly, by Equation (52), this is the same affine map as for the Gaussian case, resulting in,
$\nabla_{W}\mathbb{B}(P|\mathcal{Q})=\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\quad\text{as vector fields }\mathbb{R}^{d}\to\mathbb{R}^{d}.$ (56)
We now establish that the $L^{2}$-norms of these gradients are also equal, which is the key step for transferring the PL inequality.
Lemma D.1 ($L^{2}$-norm of Affine Maps).
Let $T(x)=Mx+c$ be an affine map with $M\in\mathbb{R}^{d\times d}$ and $c\in\mathbb{R}^{d}$. For any probability measure $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with mean $\mu_{P}$ and covariance $\Sigma_{P}$,
$\|T\|_{L^{2}(P)}^{2}:=\int_{\mathbb{R}^{d}}\|T(x)\|^{2}\,dP(x)=\text{Tr}(M^{T}M\Sigma_{P})+\|M\mu_{P}+c\|^{2}.$ (57)
In particular, $\|T\|_{L^{2}(P)}^{2}$ depends on $P$ only through its first two moments.
Proof.
Expanding the squared norm:
$\|T\|_{L^{2}(P)}^{2}$ $=\int\|Mx+c\|^{2}\,dP(x)$
$=\int\|M(x-\mu_{P})+(M\mu_{P}+c)\|^{2}\,dP(x)$
$=\int\|M(x-\mu_{P})\|^{2}\,dP(x)+2\langle M\mu_{P}+c,M\underbrace{\int(x-\mu_{P})\,dP(x)}_{=0}\rangle+\|M\mu_{P}+c\|^{2}$
$=\int(x-\mu_{P})^{T}M^{T}M(x-\mu_{P})\,dP(x)+\|M\mu_{P}+c\|^{2}$
$=\text{Tr}(M^{T}M\Sigma_{P})+\|M\mu_{P}+c\|^{2},$
where the last step uses $\mathbb{E}_{x\sim P}[(x-\mu_{P})(x-\mu_{P})^{\top}]=\Sigma_{P}$. ∎
Lemma D.2 (Gradient Norm Equality).
For any $P\in\mathcal{L}(Q_{0})$ and its corresponding Gaussian $G_{\mu,\Sigma}$:
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}=\|\nabla_{W}\mathbb{B}(G|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}.$ (58)
Proof.
First, define $T_{\text{LS}}=\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ and $T_{\text{Gaussian}}=\nabla_{W}\mathbb{B}(G_{m,\Sigma}|\mathcal{G})$. By construction, both of these mappings are affine, and by Equation 56 $T_{\text{LS}}=T_{\text{Gaussian}}$. From Lemma D.2, since $P$ and $G_{\mu,\Sigma}$ share the first and second order moments by construction, they are equal, which means,
$\lVert\nabla_{W}\mathbb{B}(P|\mathcal{Q})\rVert_{L_{2}(P)}^{2}=T_{\text{LS}}=T_{\text{Gaussian}}=\lVert\nabla_{W}\mathbb{B}(G|\mathcal{G})\rVert_{L_{2}(G_{\mu,\Sigma})}^{2},$
which is the desired result. ∎
We can now state and prove the following result.
Proposition D.2 (PL Inequality Transfer).
Let $\mathcal{L}(Q_{0})$ be a location-scatter family and $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}\subset\mathcal{L}(Q_{0})$. Let $\mathcal{G}=\{G_{\mu_{k},\Sigma_{k}}\}_{k=1}^{K}$ be the corresponding Gaussian with matching moments. If the Bures-Wasserstein barycenter functional $\mathbb{B}(\cdot|\mathcal{G})$ satisfies the PL inequality with constant $C_{\text{PL}}>0$, i.e.,
$\|\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right),$ (59)
then $\mathbb{B}(\cdot|\mathcal{Q})$ satisfies the PL inequality on $\mathcal{L}(Q_{0})$ with the same constant,
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P_{m,\Sigma})}^{2}\geq C_{\mathrm{PL}}\left(\mathbb{B}(P|\mathcal{Q})-\mathbb{B}(P^{\star}|\mathcal{Q})\right).$ (60)
Proof.
Starting from the PL inequality for Gaussians (Equation 59):
$\|\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right).$
Applying Lemma D.2 to the left-hand side:
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right).$
Applying Equation (53) and barycenter correspondence to the right-hand side:
$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}\geq C_{\mathrm{PL}}\left(\mathbb{B}(P|\mathcal{Q})-\mathbb{B}(P^{\star}|\mathcal{Q})\right).\qed$
Explicit PL Constant. By chewi2020gradient, the PL inequality holds for Gaussians under $\zeta$-regularity: if $\|\Sigma_{k}\|_{\mathrm{op}}\leq 1$ and $\det\Sigma_{k}\geq\zeta$ for all $k$, then $C_{\mathrm{PL}}=\zeta^{2}/4$. Then, Proposition D.2 extends this PL constant to location-scatter families satisfying the same spectral bounds.
Scope and Limitations. Proposition D.2 applies to location-scatter families, which are generated by affine transformations of a base measure. This includes elliptical distributions (e.g., multivariate $t$-distributions, uniform distributions on ellipsoids) and, importantly, the Swiss roll measures used in our experiments. This result does not directly apply to arbitrary measures.
Appendix E Domain Adaptation
E.1 Introduction
Statistical learning theory vapnik1999overview deals with the problem of learning a function $h:\mathcal{X}\rightarrow\mathcal{Y}$, such that it minimizes the risk of making a mistake, that is,
$h^{\star}=\underset{h\in\mathcal{H}}{\text{arg inf}}\,\biggr\{\mathcal{R}_{Q}(h)=\int_{\mathcal{X}\times\mathcal{Y}}\mathcal{L}(y,h(x))dQ(x,y)\biggr\},$ (61)
where $P\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$ is a probability measure, and $\mathcal{H}$ is a family of functions from $\mathcal{X}$ to $\mathcal{Y}$. $\mathcal{R}_{Q}(h)$ denotes the risk of $h$ under $Q$. Equation 61 interprets the learning problem as a minimization problem.
In practical scenarios, one has a dataset $\{x_{i}^{(Q)},y_{i}^{(P)}\}_{i=1}^{n}$, $(x_{i}^{(P)},y_{i}^{(Q)})\overset{i.i.d.}{\sim}P$. As a result, one can estimate the empirical risk, and consequently minimize it,
$\hat{h}=\underset{h\in\mathcal{H}}{\text{arg inf}}\,\biggr\{\hat{\mathcal{R}}_{Q}(h)=\dfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(y_{i}^{(Q)},h(x_{i}^{(Q)}))\biggr\},$ (62)
which is the foundation of the Empirical Risk Minimization (ERM) framework. In this sense, a solution $\hat{h}$ to equation 62 is said to generalize if it achieves low risk $\mathcal{R}_{Q}(h)$.
There is a fundamental assumption at play in the theory of generalization under the ERM framework. Indeed, both $\mathcal{R}_{P}$ and $\hat{\mathcal{R}}_{P}$ evaluate the risk under a fixed probability measure $P$. However, machine learning models are oftentimes confronted with data following different, but related probability measures quinonero2022dataset, a process known as distribution shift.
The main tool for tackling distribution shift is domain adaptation. This problem considers now two datasets, namely, a labeled source dataset $\{x_{i}^{(Q_{S})},y_{i}^{(Q_{S})}\}_{i=1}^{n_{S}}$, and an unlabeled target dataset $\{x_{j}^{(Q_{T})}\}_{j=1}^{n_{T}}$. The goal is to obtain a function $h$ that achieves low risk under $Q$, with the available data. A natural extension to this problem is multi-source domain adaptation, where multiple source domains are available, i.e., $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, each with its own associated dataset $\{(x_{i}^{(Q_{k})},y_{i}^{(Q_{k})})\}_{i=1}^{n_{k}}$.
E.2 Source-Only Training
A straightforward baseline in domain adaptation is the source-only training. Let $h=g\circ f$ be the composition of a feature extractor $g:\mathcal{X}\rightarrow\mathcal{U}$ and a feature classifier $f:\mathcal{U}\rightarrow\mathcal{Y}$. We can train $h$ end-to-end by minimizing,
$\mathcal{L}(\theta)=\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathcal{L}(y_{i}^{(Q_{k})},f(g(x_{i}^{(Q_{k})}))).$
This training produces an embedding space $\mathcal{U}$ that is discriminative of the classes in the domain adaptation problem. Following previous work montesuma2021wasserstein, montesuma2023multi, montesuma2024distillation, we perform adaptation in this space.
E.3 Wasserstein Barycenter Transport
The motivation for using Wasserstein barycenters in multi-source domain adaptation comes from a theoretical result of redko2019optimal. Indeed, by [redko2019optimal, Theorem 2],
$\mathcal{R}_{Q_{T}}(h)\leq\mathcal{R}_{Q_{S}}(h)+\mathbb{W}_{1}(\hat{Q}_{S},\hat{Q}_{T})+\mathcal{O}(n_{S}^{-1/2}+n_{T}^{-1/2})+\mathbb{R}(Q_{S},Q_{T}),$ (63)
where $\mathbb{R}(Q_{S},Q_{T})=\text{inf}_{h\in\mathcal{H}}\mathcal{R}_{Q_{S}}(h)+\mathcal{R}_{Q_{T}}(h)$ is the joint risk. This result establishes that when source and target measures are close, their risks are close as well. Now, assume equation 63 holds for each pair $(Q_{k},Q_{T})$, then, summing over $k$ weighted by $\lambda_{k}$,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
using the triangle inequality on $\mathbb{W}_{1}(\hat{Q}_{k},\hat{Q}_{T})$
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},P)+\mathbb{W}_{1}(P,\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
A reasonable choice to make this bound tight is choosing $P$ that minimizes $P\mapsto\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},P)+\mathbb{W}_{1}(P,\hat{Q}_{T})$.
From $p=1$ to $p>1$. The argument for $p>1$ is a bit more intricate. Since $\mathbb{W}_{1}$ is the weakest of the Wasserstein distances [villani2008optimal, Remark 6.6], we can directly move from $p=1$ to $p>1$,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
which, from the triangle inequality, results in,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)+\mathbb{W}_{p}(P,\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$ (64)
The term $\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)+\mathbb{W}_{p}(P,\hat{Q}_{T})$ does not correspond to the $p$-Wasserstein barycenter due the missing power of $p$. We can retrieve it through Young’s inequality for products. Let $a,b>0$ be real numbers, and $p,q>1$. Then $ab\leq\frac{a^{p}}{p}+\frac{b^{q}}{q}.$ We can get the inequality,
$\mathbb{W}_{p}(\hat{P},\hat{Q})\leq\dfrac{\mathbb{W}_{p}(\hat{P},\hat{Q})^{p}}{p}+\dfrac{p-1}{p},$
[…]
$\mathbb{W}_{p}(\hat{Q}_{k},P)\leq\frac{1}{p}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\frac{p-1}{p}\text{, and }\mathbb{W}_{p}(P,\hat{Q}_{T})\leq\frac{1}{p}\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}+\frac{p-1}{p}$
Plugging this back into inequality 64,
$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+$ $\dfrac{1}{p}\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}\biggr)+\dfrac{2(p-1)}{p}+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}).$
Modulo some additional constants depending on $p>1$, we have the same reasoning as $p=1$.
Algorithm 2 Wasserstein Barycenter Transport
[…]
2:Labeled target domain data $\{T_{P\rightarrow Q_{T}}(x_{i}^{(P)}),y_{i}^{(P)}\}_{i=1}^{(P)}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  Compute $P^{\star}=\text{Bar}(\lambda,\mathcal{Q})$
5:  Compute $\gamma^{\star}$ between $P^{\star}$ and $Q_{T}$ (eq. 65)
6:  Transport $(x_{i}^{(P)},y_{i}^{(P)})$ (eq. 66)
7:end for
Wasserstein Barycenter Transport $(p=2)$. montesuma2021wasserstein uses a two-step procedure for finding $P$. First, one solves,
$P^{\star}=\underset{P\in\mathcal{P}_{1,\text{ac}}(\Omega)}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(Q_{k},P)^{2},$
which is the 22-Wasserstein barycenter problem. Then, one aligns the barycenter with the target domain through OT. This is done through the barycenter mapping,
$\gamma^{\star}$ $=\underset{\gamma\in\Gamma(P,\hat{Q}_{T})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{T}}\gamma_{ij}\lVert x_{i}^{(P)}-x_{j}^{(Q_{T})}\rVert_{2}^{2},$ (65)
and,
$T_{P\rightarrow Q_{T}}(x_{i}^{(P)})$ $=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$ (66)
which is the minimizer of $P\mapsto\mathbb{W}_{2}(P,\hat{Q}_{T})^{2}$. We provide an algorithm for this strategy in Algorithm 2. Note that since this procedure generates a labeled dataset on data following the target domain measure, we can train a classifier with the pairs $\{T_{P\rightarrow Q_{T}}(x_{i}^{(P)}),y_{i}^{(P)}\}$.
E.4 Adapting Neural Network Barycenter Solvers to DA
Here, we cover how we adapted previous neural network-based barycenters to perform domain adaptation. There is one main challenge with using unlabeled barycenters for domain adaptation: we cannot directly move the barycenter to the target domain, as we would end up with more unlabeled data. A straightforward solution to this problem is moving the available labeled data that we have, that is, the source domain data. For this scenario, the barycenter domain works as a pivot domain, as we transport the source to the target through the composition,
$T_{Q_{k}\rightarrow Q_{T}}(x_{i}^{(Q_{k})})=T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})).$ (67)
In comparison, by computing a labeled barycenter, we synthesize data in $P^{\star}$.
As we demonstrated in Tables 3 through 7, this is generally beneficial for domain adaptation. In this way, we can acquire a labeled dataset in the target domain through $\{\{T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})),y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$. To isolate the effect of the quality of the computed barycenters, we estimate the mappings as we did in equations 65 and 66, that is,
$T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})=n\sum_{j=1}^{n}\gamma_{k,i,j}^{\star}x_{j}^{(P)}\quad T_{P\rightarrow Q_{k}}(x_{i}^{(P)})=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$ (68)
where $\gamma_{k}^{\star}$ is the OT plan from $Q_{k}$ to $P$, and $\gamma^{\star}$ is the OT plan from $P$ to $Q_{T}$.
Appendix F Additional Experiments
In total, we compare five domain adaptation benchmarks, divided into computer vision, neuroscience and chemical engineering. An overview of the benchmarks is presented in Table 1 in our main paper. Here, we give further details on each of the benchmarks. For computer vision, we use the Office 31 saenko2010adapting and Office Home venkateswara2017deep benchmarks. For neuroscience, we use the BCI-CIV-2a brunner2008bci and ISRUC khalighi2013automatic benchmarks. For chemical engineering, we use the TEP benchmark montesuma2024benchmarking.
F.1 Evaluation Protocol and Backbones
As we describe in the following, all benchmarks deal with multi-class classification. We measure performance using the standard classification accuracy,
$\text{accuracy}(y,\hat{y})=\dfrac{1}{n}\sum_{i=1}^{n}\delta(y_{i}-\hat{y}_{i})\times 100\%.$
For the Office 31, Office Home, Tennessee Eastman Process, and BCI-CIV-2a, we adopt the so-called leave-one-domain-out evaluation. For example, given $3$ domains $\{A,B,C\}$, we perform $3$ experiments, leaving one domain as the target domain at each time. Then, one of these experiments would be $\{A,B\}\rightarrow C$, where $\{A,B\}$ are the sources, and $C$ is the target.
As we discuss in the main paper, we perform adaptation over embeddings. To that end, we use the same experimental setting as montesuma2024lighter for the Office 31, Office Home, and TEP benchmarks. More specifically, we train a ResNet 50 and ResNet 101 for the Office 31 and Office Home benchmarks, and a CNN for the TEP benchmark. For this latter backbone, we refer readers to montesuma2024benchmarking for further details. For neuroscience benchmarks, we follow the experimental protocol of wang2024cbramod, where we fine-tune CBraMod, a transformer-based foundation model, on available source domain data. We use the exact same training settings as these authors.
F.2 Fine-Grained Results
Algorithm $\mathcal{X}\times\mathcal{Y}$ A D W Avg. $\uparrow$
ResNet50 - 67.50 95.00 96.83 86.40
WJDOT turrisi2022multi - 67.77 97.32 95.32 86.80
DaDiL-E montesuma2023multi - 70.55 100.00 98.83 89.79
DaDiL-R montesuma2023multi - 70.90 100.00 98.83 89.91
GMM-DaDiL montesuma2024lighter - 72.47 100.00 99.41 90.63
Discrete cuturi2014fast 70.55 96.43 96.49 87.82
NormFlow visentin2025computing 65.67 97.32 94.73 85.91
CW2B korotin2021continuous 69.16 94.64 95.32 86.37
NOT kolesov2024estimating 69.33 92.85 96.49 86.22
U-NOT gazdieva2024robust 68.29 97.32 95.32 86.97
WGF (ours) 69.16 98.21 96.49 87.96
Discrete montesuma2023multi 67.94 98.21 97.66 87.93
GMM montesuma2024lighter 70.13 99.11 96.49 88.54
WGF (ours) 70.03 99.11 99.41 89.52
Table 3: Classification accuracy per domain for the Office 31 benchmark.
Office 31 saenko2010adapting contains, in total, 4652 images among 31 classes of objects. These images are divided into 3 different domains: Amazon, dSLR, and Webcam. Images in the Amazon domain represent products, usually with an homogeneous background, medium resolution, and studio lightning. There are 2817 images in this domain. The dSLR domain contains high-resolution images obtained through a digital SLR camera in an office environment. There are 498 images in this domain. Finally, the Webcam domain contains images of objects captured with a webcam, also in an office environment. There 795 images in this domain. As analyzed in ringwald2021adaptiope, this benchmark is one of the most commonly used in the literature. Despite its moderate size, it presents challenges of its own. For instance, since the Amazon domain images were collected automatically, it has label noise.
We show our results in the Office 31 benchmark in Table 3. Overall, dictionary learning methods such as DaDiL montesuma2023multi or GMM-DaDiL montesuma2024lighter remain state-of-the-art for this benchmark. However, our gradient flow algorithm is able to improve previous methods, closing the gap between these two kinds of algorithms. In comparison with unsupervised barycenters, using labels in the ground-cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) has a clear advantage.
Algorithm $\mathcal{X}\times\mathcal{Y}$ Ar Cl Pr Rw Avg. $\uparrow$
ResNet101 - 72.90 62.20 83.70 85.00 75.95
WJDOT turrisi2022multi - 74.28 63.80 83.78 84.52 76.59
DaDiL-E montesuma2023multi - 77.16 64.95 85.47 84.97 78.14
DaDiL-R montesuma2023multi - 75.92 64.83 85.36 85.32 77.86
GMM-DaDiL montesuma2024lighter - 77.16 66.21 86.15 85.32 78.81
Discrete cuturi2014fast 76.13 63.91 83.11 85.55 77.17
NormFlow visentin2025computing 77.16 62.54 83.10 84.63 76.86
CW2B korotin2021continuous 74.69 63.46 82.54 85.09 76.44
NOT kolesov2024estimating 72.22 63.57 82.77 82.91 75.36
U-NOT gazdieva2024robust 75.30 64.94 82.88 84.28 76.85
WGF (ours) 74.89 64.60 83.89 85.21 77.56
Discrete montesuma2023multi 76.54 63.91 83.78 84.74 77.25
GMM montesuma2024lighter 75.31 64.26 86.71 85.21 77.87
WGF (ours) 76.13 65.52 85.81 86.24 78.42
Table 4: Classification accuracy per domain for the Office-Home benchmark.
Office-Home venkateswara2017deep contains, 15,500 images among 65 classes of objects. There are four domains: Art, Clipart, Product, and Real-World. The Art domain contains artistic depictions of objects, such as sketches, paintings, and ornamentation. It has 2427 images. The Clipart domain is a collection of graphic art representing the objects. This domain has 4365 images. Product contains images of objects without a background. It has 4439 images. Real-World contains images of objects captured with a camera. It contains 4357 images. In this sense, the Product domain is similar to the Amazon domain in the Office 31 benchmark, and the Real-World is similar to dSLR or Webcam. In comparison with the Office31 benchmark, the Office-Home benchmark adds another factor of domain variation: style (e.g., Art vs. Real-World). We show our results in Table 4. In comparison with the Office 31 benchmark, our methods get even closer to the state-of-the-art. Again, using labels in the ground cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) offers a clear advantage.
Remark F.1.
We follow previous works on visual domain adaptation, especially saenko2010adapting and venkateswara2017deep, and report the classification accuracy on fixed partitions of the target domains. For that reason, Tables 3 and 4 do not have confidence intervals. While it would be possible to run these experiments with random partitions (e.g., $k$-fold cross validation), this would mean that our results would not be comparable to previous established results in montesuma2024lighter, for instance.
Tennessee-Eastman Process (TEP) [reinartz2021extended, montesuma2024benchmarking] is a benchmark in chemical engineering. This benchmark comprises a set of simulations of a chemical reaction that occurs inside a reactor. In total, 34 variables are measured over time. As a result, we have a set of time-series representing how the reaction progresses over each simulation. In some of these simulations, one of 28 possible faults can be introduced. The goal is, based on the readings, to determine which kind of fault, or its absence, has occurred. As a result, there are 29 classes. Each simulation is also associated with a production mode, which drastically changes the statistical properties of the data generation process, thus introducing a distribution shift. As a result, each domain corresponds to one of these modes of operation. We refer readers to montesuma2024benchmarking and reinartz2021extended for more details.
Algorithm $\mathcal{X}\times\mathcal{Y}$ Mode 1 Mode 2 Mode 3 Mode 4 Mode 5 Mode 6 Avg. $\uparrow$
CNN - 80.82 $\pm$ 0.96 63.69 $\pm$ 1.71 87.47 $\pm$ 0.99 79.96 $\pm$ 1.07 74.44 $\pm$ 1.52 84.53 $\pm$ 1.12 78.48
WJDOT turrisi2022multi - 89.06 $\pm$ 1.34 75.60 $\pm$ 1.84 89.99 $\pm$ 0.86 89.38 $\pm$ 0.77 85.32 $\pm$ 1.29 87.43 $\pm$ 1.23 86.13
DaDiL-E montesuma2023multi - 90.45 $\pm$ 1.02 77.08 $\pm$ 1.21 86.79 $\pm$ 2.14 89.01 $\pm$ 1.35 84.04 $\pm$ 3.16 87.85 $\pm$ 1.06 85.87
DaDiL-R montesuma2023multi - 91.97 $\pm$ 1.22 77.15 $\pm$ 1.32 85.41 $\pm$ 1.69 89.39 $\pm$ 1.03 84.49 $\pm$ 1.95 88.44 $\pm$ 1.29 86.14
GMM-DaDiL montesuma2024lighter - 91.72 $\pm$ 1.41 76.41 $\pm$ 1.89 89.68 $\pm$ 1.49 89.18 $\pm$ 1.17 86.05 $\pm$ 1.46 88.02 $\pm$ 1.12 86.85
Discrete cuturi2014fast 90.38 $\pm$ 1.29 71.17 $\pm$ 1.26 85.27 $\pm$ 1.09 87.29 $\pm$ 0.99 83.83 $\pm$ 1.74 84.94 $\pm$ 1.87 83.81
NormFlow visentin2025computing 87.06 $\pm$ 0.57 66.25 $\pm$ 1.11 88.58 $\pm$ 1.00 89.35 $\pm$ 0.34 80.99 $\pm$ 2.21 85.08 $\pm$ 1.45 82.89
CW2B korotin2021continuous 91.72 $\pm$ 0.79 73.71 $\pm$ 1.42 88.37 $\pm$ 1.31 89.32 $\pm$ 1.14 84.66 $\pm$ 1.75 87.19 $\pm$ 1.24 85.83
NOT kolesov2024estimating 90.31 $\pm$ 1.08 72.37 $\pm$ 0.94 87.99 $\pm$ 1.31 89.28 $\pm$ 0.81 84.35 $\pm$ 1.68 85.29 $\pm$ 2.02 84.93
U-NOT gazdieva2024robust 90.82 $\pm$ 0.65 72.68 $\pm$ 1.58 88.65 $\pm$ 0.95 89.00 $\pm$ 1.28 85.22 $\pm$ 1.07 86.19 $\pm$ 1.70 85.43
WGF (ours) 91.51 $\pm$ 1.32 73.77 $\pm$ 1.24 87.44 $\pm$ 1.16 89.38 $\pm$ 0.62 84.87 $\pm$ 1.75 86.08 $\pm$ 2.23 85.51
Discrete montesuma2023multi 92.38 $\pm$ 0.66 73.74 $\pm$ 1.07 88.89 $\pm$ 0.85 89.38 $\pm$ 1.26 85.53 $\pm$ 1.35 86.60 $\pm$ 1.63 86.09
GMM montesuma2024lighter 92.23 $\pm$ 0.70 71.81 $\pm$ 1.78 84.72 $\pm$ 1.92 89.28 $\pm$ 1.55 87.51 $\pm$ 1.73 82.49 $\pm$ 1.81 84.67
WGF (ours) 92.34 $\pm$ 1.05 76.10 $\pm$ 1.56 89.78 $\pm$ 0.90 89.32 $\pm$ 1.03 85.92 $\pm$ 1.35 87.74 $\pm$ 1.52 86.87
Table 5: Classification accuracy per domain on the Tennessee Eastman Process benchmark.
We report our results in Table 5. For this benchmark, the gap between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) barycenter methods is closer. For instance, CW2B korotin2021continuous achieves 85.83 for the average target domain performance versus 86.09 for the labeled discrete barycenter. Furthermore, it surpasses the GMM method of montesuma2024lighter. In comparison, our methods establish a new state-of-the-art for barycenter-based domain adaptation in this benchmark, having similar performance to GMM-DaDiL montesuma2024lighter.
Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 Avg. $\uparrow$
CBraMod - 52.22 40.45 64.23 40.41 47.74 48.34 53.64 50.52 55.03 50.30
DaDiL-R montesuma2023multi - 58.68 40.27 66.14 41.45 59.54 51.38 53.12 57.46 55.21 53.69
GMM-DaDiL montesuma2024lighter - 60.59 42.36 70.31 48.12 61.11 50.34 62.15 58.85 60.00 57.10
Discrete cuturi2014fast 60.93 43.40 69.96 50.00 61.28 50.69 61.63 61.45 57.11 57.38
NormFlow visentin2025computing 57.11 43.22 67.53 47.71 60.24 51.04 56.25 60.76 55.20 55.45
CW2B korotin2021continuous 60.00 43.57 70.13 48.95 60.93 51.21 61.28 61.63 57.46 57.25
NOT kolesov2024estimating 61.28 43.22 70.31 47.91 60.59 51.21 61.46 61.80 57.29 57.23
U-NOT gazdieva2024robust 60.59 44.09 69.44 47.70 61.11 50.34 61.97 61.11 58.15 57.17
WGF (ours) 60.93 43.75 69.96 49.16 60.76 51.38 61.97 63.54 57.64 57.68
Discrete montesuma2023multi 60.76 42.53 70.66 48.54 61.28 50.69 61.63 61.28 59.54 57.43
GMM montesuma2024lighter 60.59 42.53 69.79 47.08 61.45 50.34 62.32 59.54 59.72 57.04
WGF (ours) 60.93 46.18 68.92 50.21 62.32 50.69 61.63 61.11 57.11 57.68
Table 6: Classification accuracy per domain on the BCI-CIV-2a benchmark.
BCI-CIV-2a brunner2008bci Contains electroencephalogram (EEG) data from nine subjects. In brief, an EEG measures the electrical activity of the brain over time. As such, like the TEP benchmark, this benchmark has a collection of time series data. The BCI-CIV-2a data is concerned with motor imagery tasks, that is, the imagination of movement of the left hand, right hand, both feet, and tongue. These correspond to class 1, 2, 3, and 4. Therefore, in this benchmark we have four classes. Our goal here is to perform cross-subject adaptation, namely, we use data from a given set of source subjects, and try to predict on a target subject. Therefore, each domain is a subject, so there are nine domains in total. There are, on average, $560$ samples per subject, with a total of $5088$ samples.
We show our results in Table 6. For cross-subject studies, we highlight that, besides improving individual domain performance, it is desirable to not degrade over the source-only performance. This is indeed the case for almost all methods. For instance, DaDiL montesuma2023multi degrades performance of subjects 2 and 7 by a small margin. For this benchmark, the difference between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) is not as marked as before. We hypothesize that this benchmark falls into the covariate shift hypothesis, namely, $Q_{S}(X)\neq Q_{T}(X)$, rather than the more general joint shift $Q_{S}(X,Y)\neq Q_{T}(X,Y)$.
Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
CBraMod - 78.97 80.35 70.46 85.59 80.69 72.85 77.00 73.29 83.41 68.95 76.63
WJDOT turrisi2022multi - 80.41 78.25 65.69 82.85 81.74 77.74 76.33 68.75 86.95 70.81 76.95
DaDiL-E montesuma2023multi - 76.12 77.90 65.93 83.21 81.28 68.45 75.66 72.61 79.75 65.93 75.14
DaDiL-R montesuma2023multi - 78.06 78.13 66.97 84.64 81.51 72.02 76.33 72.84 80.85 67.55 75.89
GMM-DaDiL montesuma2024lighter - 76.63 78.95 66.51 82.73 77.44 73.81 77.00 71.70 83.05 66.86 75.47
Discrete cuturi2014fast 78.77 80.58 69.76 85.83 79.06 73.45 78.33 72.50 84.14 68.48 77.09
NormFlow visentin2025computing 81.12 82.90 75.58 84.28 80.34 78.69 75.88 70.34 80.48 70.81 78.05
CW2B korotin2021continuous 79.08 78.25 69.07 86.54 78.48 70.35 76.88 70.90 82.80 66.04 75.84
NOT kolesov2024estimating 78.57 79.88 67.32 85.59 78.72 73.21 77.77 72.50 83.41 67.21 76.44
U-NOT gazdieva2024robust 80.20 83.48 73.72 87.50 76.51 72.62 76.66 70.22 78.17 71.51 77.06
WGF (ours) 79.69 81.86 70.58 86.66 78.14 77.74 78.77 72.16 85.97 70.00 78.16
Discrete montesuma2023multi 78.87 80.69 72.09 86.07 78.60 78.45 78.00 72.95 85.48 70.81 78.20
GMM montesuma2024lighter 75.20 75.46 69.88 80.71 78.60 71.90 76.33 68.41 78.17 68.14 74.28
WGF (ours) 80.10 80.81 76.62 85.95 76.86 80.35 76.55 77.72 84.87 73.95 80.02
Table 7: Classification accuracy per domain on the ISRUC benchmark.
ISRUC khalighi2013automatic is a benchmark for EEG data for sleep staging. In brief, sleep staging is the classification of physiological data, such as EEG, recorded during sleep into different sleep stages. In this context, the ISRUC benchmark is the most large scale benchmark used in this paper. This study comprises 100 subjects, with a total of 89240 samples (approximately 900 samples per domain). The sleep staging problem in this benchmark has five classes: awake, non-rapid eye movement sleep (subdivided into N1, N2, N3), and rapid eye movement sleep.
Due the elevated number of subjects in this benchmark, we adopt a slightly different evaluation protocol. Instead of using the leave-of-domain-out strategy, we fix the first $90$ subjects as training subjects and evaluate the performance of adapting to the remaining $10$ subjects. This means that we compute Wasserstein barycenters of $K=90$ input measures. Overall, our WGF method has clear advantage over all other methods, with unsupervised WGF (78.16% average target accuracy) being the best over unlabeled barycenters, and labeled barycenters as well (80.02% average test accuracy).
F.3 Complexity Analysis and Additional Running Time
We start with analyzing the complexity of evaluating the functionals used in our work. We recall that,
$\mathbb{V}(P)=\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)}),\quad\mathbb{U}(P)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)}),$
where, assuming that the complexity of evaluating $V$ and $U$ is $\mathcal{O}(1)$, the complexity of evaluating $\mathbb{V}$ is $\mathcal{O}(n)$ and for $\mathbb{U}$ is $\mathcal{O}(n^{2})$. The complexity of $\mathbb{B}$ is usually the bottleneck, as it involves computing $K$ OT problems. For exact OT, this means we need to solve $K$ OT problems of size $m\times n$, where $m$ is the batch size and $n$ is the barycenter support size. These OT problems need to be solved sequentially, as parallelizing $K$ independent linear programs is non-trivial. In practice, we use the network simplex solver of bonneel2011displacement, wrapped in Python by flamary2021pot. This algorithm has worst-case complexity of $\mathcal{O}(n^{3})$ but in practice, runs in $\mathcal{O}(n^{2})$. Assuming $n\gg m$, the complexity becomes,
$\text{Exact barycenter functional complexity}=\mathcal{O}(K\times m\times n),$
which scales linearly with the barycenter support size $n$. Though this might appear contradictory with the usual $\mathcal{O}(n^{3})$ complexity of OT, we highlight that we are keeping $m$ constant as we scale the barycenter support size $n$. The same analysis can be done for entropic OT,
$\text{Entropic barycenter functional complexity}=\mathcal{O}(K\times L\times m\times n),$
where, instead of a factor $m^{2}$, we have $L\times m$, where $L>0$ is the number of Sinkhorn iterations. Both of these approaches scale linearly with the number of samples in the barycenter support. However, there are 2 important advantages of using entropic OT. First, we can use GPU acceleration for computing the OT variables. Second, as we shown in Section 3.1, we can vectorize the $K$ OT problems. While this does not alleviate the $K$ term in the complexity, it does lead to further speedups. These factors explain the difference, in running time, between exact and entropic OT.
(a) Scaling with $n$ ($\epsilon=0$)
(b) Scaling with $n$ ($\epsilon>0$)
(c) Scaling with $\epsilon$ ($m=256$)
Figure 8: Running time analysis in the Swiss roll example. In (a) and (b), we show that our WGF approach improves the scaling of the discrete solver of cuturi2014fast. In (c), we isolate the speedups of GPU parallelism and Sinkhorn vectorization.
Next, we perform a running time analysis of the tested methods on the ISRUC benchmark, which contains the most number of samples and domains. For all neural net methods, we fix a batch size of $256$ samples. Overall, our proposed methods run significantly faster than previous methods. For instance, in comparison with the discrete solver of cuturi2014fast, our algorithm is approximately seven times faster per iteration. It is noteworthy that some neural net methods are relatively faster per iteration (e.g. NormFlow). However, we empirically verified that these methods sometimes need a large number of iterations to converge to a good barycenter (e.g. 5000 iterations).
Method Complexity per iteration Parameter Complexity Running time per iteration (s) # Iterations Running Time (s)
Discrete cuturi2014fast, montesuma2023multi $Kn^{3}\log n$ $n$ 4.28 50 214.37
NormFlow visentin2025computing N/A $|f|$ 0.17 5000 878.32
CW2B korotin2021continuous N/A $K(|f|+|g|)$ 2.18 1000 2181.00
NOT kolesov2024estimating N/A $K(|f|+|g|)$ 41.63 25 1040.75
U-NOT gazdieva2024robust N/A $K(|f|+|g|)$ 13.29 100 1329.77
WGF (ours, $\epsilon=0$) $Km^{2}n$ $n$ 0.05 200 10.43
WGF (ours, $\epsilon=10^{-3}$) KLmnKLmn $n$ 0.03 200 7.00
Table 8: Complexity and running time (in seconds) of different barycenter calculation strategies on the ISRUC benchmark. $n_{b}$ denotes the batch size, $n$ number of samples, $K$ number of input measures, $d$ number of dimensions. For NormFlow, $f$ denotes the neural net implementing the normalizing flow. For CW2B, NOT and U-NOT, $f$ and $g$ denote the neural networks approximating the Kantorovich potentials. $|f|$ and $|g|$ denote the number of their parameters.
F.4 Effect of Different Functionals
(a) $(10^{-3},10^{-3})$
(b) $(10^{-4},10^{-3})$
(c) $(10^{-5},10^{-3})$
(d) $(10^{-3},0.0)$
(e) $(10^{-4},0.0)$
(f) $(10^{-5},0.0)$
Figure 9: Wasserstein barycenter as a function of $(\eta_{G},\epsilon)$, i.e., diffusion coefficient $\xi$, and entropy regularization $\epsilon$.
On the effect of $\mathbb{G}$. The internal energy functional [chizat2025doubly] controls the smoothness of $\mathbb{G}$, and acts, in the case of the entropy functional $G(s)=s\log s$, as a diffusion term. In Figure 9, we show the effect of varying $\xi$ (c.f. equation 16) in the final barycenter. Overall, an aggressive diffusion coefficient leads to fuzzy barycenters (c.f. Figure 9 (a)), while a coefficient too small leads to low variance measures (c.f. Figure 9 (c)). This is, in fact, an artifact of the entropic OT. While a large diffusion coefficient does lead to fuzzy barycenters with exact OT (c.f. Figure 9 (d)), using a small coefficient does not degrade the quality of the computed barycenter.
(a) Unsupervised
(b) Unsupervised
(c) Supervised
(d) Supervised
Figure 10: t-SNE embeddings of unsupervised (a, b) and supervised (c, d) barycenters. While it is possible to group some samples according their pseudo-labels based on $\mathbb{U}$, the task is in general difficult. For the supervised case, the interaction energy is able to make classes more separable (d, in comparison to c).
On the effect of $\mathbb{U}$. We now visualize the effect of adding the repulsion interaction energy functional in the Office-Home benchmark. For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 22 and we fix $\text{margin}=1$, which encourages class clusters to live in orthogonal sub-spaces of the latent space. Overall, adding $\mathbb{U}$ leads to well separated classes.
Figure 11: Distribution of label entropy values as a function of potential energy scaling $\eta$.
In Figure 10, we show a comparison between unsupervised and supervised regularized barycenters. In (a), unsupervised, unregularized barycenters do not have any label structure, since there is no label supervision. In (b), we show that by adding the interaction energy functional, we can cluster samples with the same assigned label – even without input measure label supervision. In (c) and (d), we show the supervised case, in which labels in the input measures are available. In (c), classes are well clustered, but some of them overlap in the embedding space. Adding the interaction energy functional $\mathbb{U}$ effectively pushes these apart, as shown in (d). Overall, we conclude that enforcing well separated class clusters is not straightforward without the inductive bias coming from the labels.
On the effect of $\mathbb{V}$. In Figure 11, we show the distribution of label entropy values (c.f. equation 22) as a function of the potential scaling $\eta$,
$\mathbb{V}(P)=\frac{\eta}{n}\sum_{i=1}^{n}V(y_{i}^{(P)}).$
We see that increasing the weight of this functional leads to increasingly sharp labels.
F.5 Hyperparameter Setting
In this section, we detail the hyperparameters of our methods. For both flows, we have mainly three parameters,
  • number of samples in the barycenter support, $n$.
  • learning rate $\alpha$ for the gradient flow,
  • number of iterations $n_{\text{iter}}$,
additionally, for the empirical flow, we have the batch size $m$ for the input measures. We show the complete list of hyperparameters in Table 9. Additionally to these parameters, we also have the choice of regularizing functionals, $\mathbb{V}$ and $\mathbb{U}$.
Benchmark $n$ $m$ $\alpha$ $n_{\text{iter}}$ $\epsilon$ $\mathbb{V}$ $\eta_{V}$ $\mathbb{U}$ $\eta_{U}$ margin $\eta_{G}$
Supervised $(\mathcal{X}\times\mathcal{Y})$
Office31 620 620 1.0 100 $10^{-3}$ Entropy $10^{2}$ Repulsion $^{3} 1.01.0$ $10^{-2}$
[…]
TEP 870 145 1.0 50 $0.0$ Entropy $10^{2}$ Repulsion $10^{3}$ 1.0 0.0
BCI-CIV-2a 40 80 1.0 50 $10^{-2}$ Entropy $10^{2}$ Repulsion $10^{5}$ 1.0 0.0
ISRUC 200 100 0.1 200 0.0 - - - - - -
Unsupervised $(\mathcal{X})$
[…]
TEP 580 145 1.0 50 0.0 - - - - - -
BCI-CIV-2a 80 40 1.0 50 $10^{-2}$ - - - - - -
ISRUC 250 50 0.1 200 $10^{-3}$ - - - - - -
Table 9: Hyperparameters used in each benchmark. $n$ denotes number of samples, $m$ denotes the batch size, $\alpha$ denotes learning rate, $n_{\text{iter}}$ denotes number of iterations.
We offer some additional reasoning on how to choose these parameters. The number of samples controls the complexity of the calculated barycenter. We express these values in terms of the number of classes. For instance, $n=1550$ in the Office 31 benchmark translates to $n=31\times 50$ samples, that is, $50$ samples per class. The same reasoning applies to the batch size $m$. Furthermore, our method is fairly robust to the choice of learning rate. For instance, as we discussed in section C.1, the fixed-point iterations in the discrete barycenter are equivalent to setting $\alpha=n/2$, where $n$ is the number of samples in the Wasserstein barycenter. This generally strikes a balance on how fast the algorithm converges.
F.6 Ablation
Table 10 ablates the effect of the functionals $\mathbb{G},\mathbb{V},\mathbb{U}$ in the Office-Home benchmark. First, our WGF consistently outperforms the discrete baseline (77.17 vs. 77.53 and 77.24 vs. 78.42 in the unlabeled and labeled cases). Second, $\mathbb{V}$ acts as a stabilizer with high entropic regularization ($\epsilon=10^{-2}$). While the entropic regularizer blurs the barycentric measure, $\mathbb{V}$ has a sharpness effect. Third, our WGF is more sensitive to the entropic regularization, $\epsilon$, particularly when used as a loss function (our case) rather than for OT plan calculations, as in cuturi2014fast. This latter point highlights the importance of using a sharpness functional. We explore the effects of $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ in further detail in Appendix F.4.
Method $\epsilon$ $\mathbb{B}$ $\mathbb{B}{+}\mathbb{U}$ $\mathbb{B}{+}\mathbb{V}$ $\mathbb{B}{+}\mathbb{V}{+}\mathbb{U}$ $\mathbb{B}{+}\mathbb{G}$ $\mathbb{B}{+}\mathbb{G}{+}\mathbb{V}$ $\mathbb{B}{+}\mathbb{G}{+}\mathbb{U}$ $\mathbb{B}{+}\mathbb{G}{+}\mathbb{V}{+}\mathbb{U}$
Unlabeled setting ($\mathcal{X}$ only)
Discrete 0.0 77.17
Discrete 0.01 76.61
Discrete 0.001 76.56
WGF (ours) 0.0 77.41 76.83 77.41 77.13 76.88 76.88 77.53 77.53
WGF (ours) 0.01 75.54 74.42 75.26 75.26 74.55 74.55 75.00 75.00
WGF (ours) 0.001 77.15 76.88 76.94 76.94 77.28 77.28 77.26 77.26
Labeled setting ($\mathcal{X}\times\mathcal{Y}$)
Discrete 0.0 77.24
Discrete 0.01 76.73
Discrete 0.001 76.64
WGF (ours) 0.0 77.97 78.12 78.07 78.25 78.11 78.11 78.19 78.19
WGF (ours) 0.01 77.59 77.87 77.83 77.83 77.83 77.88 77.89 77.94
WGF (ours) 0.001 78.16 78.16 78.15 78.16 77.99 78.07 78.42 78.37
Table 10: Ablation on the contribution of each energy functional on domain adaptation performance.